### v1

# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

Abstract
Most existing necessary conditions for optimal control based on adjoining methods require both state information and costate information, yet the lack of costates for a given feasible trajectory in practice impedes the determination of optimality. This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law that represents the input control and the feasibility in a compact form. Given a feasible trajectory, the disturbed trajectory under the constraints of augmented switching law is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information. A first order necessary condition is proposed that the Jacobian matrix of the augmented switching law is not full row rank, which also results in an approach to optimizing a given feasible trajectory further. The proposed necessary condition is applied to the chain-of-integrators systems with full box constraints, contributing to some conclusions challenging to reason by traditional costate-based necessary conditions.
Index Terms:
Optimal control, linear systems, variational methods, switched systems, necessary condition.
I Introduction
Time-optimal control for controllable linear systems achieves universal applications in aerospace [1], manufacturing [2, 3], robotic control [4, 5], and autonomous driving [6]. Numerous works have been conducted on the behaviors of the optimal controls, resulting in some well-known necessary or sufficient conditions [7]. However, while referring to existing conclusions of optimal control for problems in general forms, it remains difficult to fully solve a high order problem in practice, especially when state inequality constraints are introduced into the system [8]. Specifically, when given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on state information since most existing necessary conditions require costate information.
In the domain of optimal control theory, numerous necessary conditions are developed based on the variational method [9] and Pontryagin’s maximum principle (PMP) [10]. The former one provides a necessary condition on the extreme of a functional in an open feasible set, while PMP has the potential to deal with inequality constraints on states or controls [11]. Among them, the state-control inequality constraints pose a challenge in solution since they involve the connection of different arcs. Chang [12], Dreyfus [13], and other scholars developed the direct adjoining method for the connection of unconstrained arcs and constrained arcs, i.e., the junction conditions of costates. Jacobson et al. [14] pointed out that constrained arcs are not allowed when the state constraint is of odd order $p>1$, and instead, the unconstrained arc is tangent to the constrained boundary at a single point. Makowski and Neustadt [15] generalized the PMP-based necessary condition to the mixed constraints of state and control. Bryson et al. [16] developed the necessary condition based on the indirect adjoining method, further supplemented by Kreindler [17]. Hartl et al. [7] systematically reviewed existing PMP-based necessary conditions and summarized their general forms. The necessary conditions are usually applied to rule out some forms of non-optimal controls in practice. In some simple cases, the optimal control can be fully solved based on some necessary conditions [18, 19], while in other cases, the necessary conditions are applied to guide the design of suboptimal trajectories [8]. However, the above necessary conditions require information on the states and the costates simultaneously, while the costates are not directly observable in practice. Given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on the pure information on states, let alone the further optimization based on the given trajectory.
In this paper, a necessary condition is called to be state-centric if it only requires information on states and controls, without dependence on costate information. There have existed some state-centric necessary so far, most of which are based on Bellman’s principle of optimality [20]. Bellman’s principle of optimality states that a sub-arc of an optimal trajectory is also optimal for the induced sub-problem, where the costates are not introduced. The Hamilton-Jacobi-Bellman (HJB) equation [21] represents a state-centric necessary in the form of the partial differential equation, which is widely used in optimal control [22] including reinforcement learning [23]. Evans and James [24] investigated the HJB equation for time-optimal control. Wolenski and Zhuang [25] showed that the minimal time function is the unique proximal solution to the HJB equation. However, most numerical methods based on the HJB equation require high computational cost [26], where the continuous time domain is discretized into a grid. A trade-off exists between the computational cost and the computational accuracy in practice. It remains challenging to develop a state-centric necessary condition for time-optimal control in a compact form with limited computational cost.
This paper sets out to establish a theoretical framework for time-optimal control of controllable linear systems and develop a novel state-centric necessary condition. The optimal control problem is formulated and investigated based on Hamiltonian in Section II. Section III analyzes the structure of arcs and the switching law of the optimal trajectory, where the connections between adjacent arcs are fully discussed. The keypoints for feasibility are introduced in Section IV, where the augmented switching law is proposed to represent the control and the feasibility in a compact form. Based on the uniqueness of the optimal control and the feasibility of the disturbed trajectory, Section V develops a novel state-centric necessary condition for time-optimal control. The proposed necessary condition is further applied to the chain-of-integrators systems with full box constraints in Section VI as examples. The contributions of this paper are as follows:
  1. 1.
    This paper establishes an innovative theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law which fully represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form. The established framework focuses on the control of each unconstrained arc and constrained arc, the constraints induced by the connections of adjoining arcs, the feasibility of the disturbed trajectory near the constrained boundary, and the optimality of a given feasible trajectory. In the proposed theory, the control can be fully represented by the augmented switching law and the motion time of each arc, where the constraints directly induced by the augmented switching law can guarantee the feasibility of the trajectory near the constrained boundary in a sufficient and necessary sense. The developed framework provides a novel variational approach as well as a local optimization method with feasibility assurance, i.e., disturbing the motion time of each arc under fixed augmented switching law, resulting in disturbed trajectories with feasibility guarantees.
  2. 2.
    This paper proposes a novel state-centric necessary condition for optimal control of controllable linear systems, which only requires the state and control information without dependence on costate information. Given a Bang-Singular-Bang feasible trajectory which is claimed to be optimal, all disturbed feasible trajectories should result in a terminal time which is strictly longer than that of the original trajectory. In this way, a first order necessary condition is proposed that the Jacobian matrix of the augmented switching law should not be full row rank. The higher order necessary conditions are also expected to be defined based on the strict local optimality of the trajectory. Based on the proposed necessary condition, a novel optimization approach with a given Bang-Singular-Bang feasible trajectory as initial value is proposed, where the optimized trajectory always has the same augmented switching law as the initial trajectory in a Bang-Singular-Bang form and the feasibility is always guaranteed during iteration.
  3. 3.
    The proposed state-centric necessary condition is applied to high-order chain-of-integrators systems with full box constraints as examples, contributing to some conclusions that are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed necessary condition. Note that time-optimal control for chain-of-integrators systems with full box constraints represents a challenging and open problem in optimal control theory and kinematics, yet to be resolved. The number of arcs is proved to be less than the order of the system plus the sum of the order of arcs’ constraints under some conditions, while the traditional methods have to analyze the costates case-by-case. The proposed state-centric necessary condition successfully proves the recursive equations of motion time if the chattering phenomenon occurs and is induced by the 2nd order state constraints, which is challenging based on costate analysis due to the high degree-of-freedom of costates in high order problems. As a result, it is hopeful to determine the existence of chattering induced by the 2nd order state constraints in chain-of-integrators systems in future works.
II Problem Formulation and Assumptions
This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is defined as follows:
$\min_{u}\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t={t_{\mathrm{f}}},$ (1a)
${\mathrm{s.t.}}\quad$ $\dot{{\boldsymbol{x}}}\left(t\right)={\boldsymbol{A}}x\left(t\right)+{\boldsymbol{b}}u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (1b)
${\boldsymbol{C}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}},\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (1c)
${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{t_{\mathrm{f}}},$ (1d)
$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (1e)
[…]
Assume that the system (1b) is controllable, i.e.,
${\mathrm{rank}}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]=n.$ (2)
Remark.
Generally, an equality constraint ${\boldsymbol{F}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ can be introduced into problem (1a). However, it is not considered in this paper since the equality constraint can be equivalently eliminated by a linear transformation.
In order to solve (1a), the Hamiltonian is constructed as
${\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)$ (3)
$=$ $\lambda_{0}+{\boldsymbol{\lambda}}^{\top}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)+{\boldsymbol{\eta}}^{\top}\left({\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}\right).$
Among them, $\lambda_{0}\geq 0$ is a constant. ${\boldsymbol{\lambda}}=\left(\lambda_{k}\right)_{k=1}^{n}$ is the costate vector. $\lambda_{0}$ and ${\boldsymbol{\lambda}}$ satisfy $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\not=0$. The Euler-Lagrange equations [27] holds that
$\dot{\boldsymbol{\lambda}}=-\frac{\partial{\mathcal{H}}}{\partial{\boldsymbol{x}}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}-{\boldsymbol{C}}^{\top}{\boldsymbol{\eta}}.$ (4)
In (3), ${\boldsymbol{\eta}}=\left(\eta_{p}\right)_{p=1}^{P}$ is the multiplier vector induced by the state inequality constraint (1c), satisfying
$\eta_{p}\geq 0,\,\eta_{k}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)=0,\,\forall 1\leq p\leq P.$ (5)
Therefore, $\forall t\in\left[0,t_{\mathrm{f}}\right]$, $\eta_{p}\left(t\right)>0$ only if ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}=0$.
Pontryagin’s maximum principle (PMP) [7] states that the optimal control $u\left(t\right)$ minimizes the Hamiltonian ${\mathcal{H}}$, i.e.,
$u\left(t\right)\in\mathop{\arg\min}\limits_{\left|U\right|\leq{u_{\mathrm{m}}}}{\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),U,\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right).$ (6)
By (3), ${\mathcal{H}}$ is affine in $u$; hence,
$u\left(t\right)=\begin{dcases}{u_{\mathrm{m}}},&{\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\left(t\right)<0,\\ *,&{\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\left(t\right)=0,\\ -{u_{\mathrm{m}}},&{\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\left(t\right)>0.\end{dcases}$ (7)
$u\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ is undetermined during ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, which is called a singular arc. The notation $\equiv$” means that the equality holds for a continuous period. (7) is the well-known Bang-Singular-Bang law [8].
Note that the objective function $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
$\forall t\in\left[0,t_{\mathrm{f}}\right],\,{\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)\equiv 0.$ (8)
For convenience of discussion, the following assumptions are introduced:
Assumption 1.
Assume that problem (1a) is feasible, and the optimal control exists. Furthermore, assume that the chattering phenomenon does not occur in the optimal profile.
Remark.
In Assumption 1, the chattering phenomenon [28] means that the optimal control switches infinitely many times in a finite period. At a one-sided neighborhood of a chattering limit time point, an infinite number of constrained arcs are joined at the boundary of inequality state constraints, where chattering limit time points are usually isolated [28]. If a chattering phenomenon occurs, then, by Bellman’s principle of optimality [20], a sub-arc of the optimal trajectory is also optimal in the sub-problem. Hence, a sub-arc that contains no chattering limit time points can be considered in this paper.
In the following of this paper, Assumption 1 is considered to hold throughout unless otherwise specified.
III Arc Representation of the Optimal Trajectory
An arc of a trajectory is defined as a continuous segment of the optimal trajectory conditioned by the same input control. Two kinds of arcs exist, i.e., unconstrained arcs and constrained arcs, which are discussed in Section III-A and Section III-B, respectively. Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C. Finally, if the chattering phenomenon does not occur, an optimal trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
III-A Unconstrained Arcs in Problem (1a)
In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds strictly almost everywhere, and the control does not switch. The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
Proposition 1.
In an unconstrained arc, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ almost everywhere, and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
Proof.
Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period. By (5), ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ in an unconstrained arc. Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$. Considering the $k$-th order derivative of ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, it has
$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}\equiv 0.$ (9)
By (2), ${\boldsymbol{\lambda}}\equiv 0$. Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$. Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$.∎
Remark.
By $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$, either ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period or ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 for finite times, while the former case is ruled out in an unconstrained arc by Proposition 1. The control switches between $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$ when ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0, where two unconstrained arcs are connected.
III-B Constrained Arcs in Problem (1a)
Compared to unconstrained arcs, constrained arcs exhibit more complex behavior. In a constrained arc, $\exists 1\leq p\leq P$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$. The following proposition characterizes an unconstrained arc.
Proposition 2.
If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in a constrained arc, then $\exists 1\leq r_{p}\leq n$, s.t. $\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$. During the constrained arc,
$u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}},$ (10)
and
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\equiv 0.\end{dcases}$ (11)
Conversely, if the initial value ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{0}\right)=0,\end{dcases}$ (12)
then under the driving of (10), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds.
Proof.
Since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$, (2) implies the existence of the above $r_{p}$. Considering the $k$-th order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$. Hence, (10) holds.
Assume that (10) holds and ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (12). Note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}=0$; hence, the initial condition implies that (11) holds. ∎
In the above proposition, $r_{p}$ is called the order of the equality state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0$. Similarly, the analytical expression of a constrained arc can be obtained by Lemma 1.
Cases exist where multiple inequality constraints are active in a constrained arc. For this case, the active constraints should be consistent. Specifically, denote that $\mathcal{P}=\left\{p\in\left[P\right]:{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0\right\}$, where $\left[P\right]\triangleq\left\{1,2,\dots,P\right\}$. Then, $\forall p_{1},p_{2}\in\mathcal{P}$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}-1}{\boldsymbol{b}}}=\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}-1}{\boldsymbol{b}}}$ holds in the arc due to (10), and $\forall p\in\mathcal{P}$, (11) holds. The following example illustrates the above analysis:
Example 1.
Consider an $n$-th order chain-of-integrators system, i.e.,
$\begin{dcases}\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right],\,\forall 1<k\leq n,\\ \dot{x}_{1}\left(t\right)=u\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right].\end{dcases}$ (13)
Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$. Assume that both $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ hold in a constrained arc. Then, $u\equiv-x_{2}=0$; hence, $x_{3}\equiv 1$, $x_{2}\equiv 0$, and $x_{1}\equiv 0$. Therefore, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent. Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are not consistent.
III-C System Behavior
Following the terminology introduced in our previous work [8], the system behavior is defined as follows:
Definition 1.
A system behavior in problem (1a) is an arc with constant system dynamics, denoted as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$. During the arc, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$. $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{g}}}\equiv{\boldsymbol{0}}$ provides fully equality state constraints, where $\widehat{{\boldsymbol{F}}}$ has full row rank. $\mathcal{P}$ is the set of active inequality state constraints.
In Definition 1, an unconstrained arc can be denoted as ${\mathcal{S}}=\left({\boldsymbol{A}},u_{0}{\boldsymbol{b}},\sim,\sim,\varnothing\right)$, where $\sim$ means that no inequality constraints are introduced, and $u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ is a constant.
For a constrained arc, $\mathcal{P}\not=\varnothing$. (10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{p}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}_{p}$, where
$\widehat{{\boldsymbol{A}}}_{p}={\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\widehat{{\boldsymbol{b}}}_{p}={\boldsymbol{0}}.$ (14)
Although it allows that $\exists p_{1},p_{2}\in\mathcal{P}$, s.t. $\widehat{{\boldsymbol{A}}}_{p_{1}}\not=\widehat{{\boldsymbol{A}}}_{p_{2}}$, it is guaranteed by the consistency of active constraints that $\widehat{{\boldsymbol{A}}}_{p_{1}}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{A}}}_{p_{2}}{\boldsymbol{x}}$ during the arc. For convenience of discussion, let $\widehat{{\boldsymbol{A}}}$ be $\widehat{{\boldsymbol{A}}}_{p}$ with the minimum $p\in\mathcal{P}$, i.e., $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$. Evidently, $\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}}\right)$ is well-defined in Definition 1.
The active inequality state constraints state that $\forall p\in\mathcal{P}$, (11) holds. The above equalities induce $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ where $\widehat{{\boldsymbol{F}}}$ has full row rank. For well-posedness, $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ are determined by the equality constraints with minimum indices.
Example 2.
Consider a chain-of-integrators system (13) with full box state constraints, i.e., $\forall 1\leq k\leq n$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$-th order, where $\forall 1\leq j<k$, $x_{j}\equiv 0$, and $u\equiv 0$. Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$. Among them, $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i,j=1}^{n}$, where $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$, and otherwise, $\hat{a}_{ij}=0$.
In a system behavior ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ holds. Therefore, the analytical expression of ${\boldsymbol{x}}\left(t\right)$ in ${\mathcal{S}}$ can be given by the following lemma.
Lemma 1.
The initial value problem (IVP) where $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}\left(t\right)$, $t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ has a unique solution:
$\forall t\in{\mathbb{R}},\,{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau.$ (15)
Proof.
$\frac{\mathrm{d}}{\mathrm{d}t}\left({\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}\left(\dot{\boldsymbol{x}}\left(t\right)-{\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{b}}\left(t\right)$ implies that ${\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)-{\mathrm{e}}^{-{\boldsymbol{A}}t_{0}}{\boldsymbol{x}}_{0}=\int_{t_{0}}^{t}{\mathrm{e}}^{-{\boldsymbol{A}}\tau}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau$. ∎
Lemma 1 implies that for an unconstrained arc where $u\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$,
${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right)+u_{0}\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (16)
For a constrained arc,
${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{\widehat{{\boldsymbol{A}}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right).$ (17)
For a given system behavior, ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ has an explicit elementary expression since every scalar element of ${\boldsymbol{e}}^{\widehat{{\boldsymbol{A}}}t}$ is composed of exponentials and polynomials w.r.t. $t$.
III-D Switching Law
Based on the analysis in Section III-C, if the optimal trajectory consists of a finite number of arcs, then the optimal trajectory can be represented by a sequence of system behaviors. The switching law is defined as follows:
Definition 2.
Assume that the optimal trajectory in problem (1a) consists of a finite number of arcs represented as ${\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{N}$ successively. The switching law of the optimal trajectory is defined as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
Remark.
The switching law focuses on the arcs that compose the optimal trajectory, while the motion time of each arc is not included in a switching law. Furthermore, Definition 2 does not characterize the chattering phenomenon since chattering induces an infinite number of arcs.
The connection between arcs is analyzed in Proposition 3. Denote the two adjacent arcs as ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$, where ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. $\forall i=1,2$, ${\mathcal{S}}_{i}=\left(\widehat{{\boldsymbol{A}}}_{i},\widehat{{\boldsymbol{b}}}_{i},\widehat{{\boldsymbol{F}}}_{i},\widehat{{\boldsymbol{g}}}_{i},\mathcal{P}_{i}\right)$. Among them, ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ is defined as $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}$, $\widehat{{\boldsymbol{b}}}_{1}=\widehat{{\boldsymbol{b}}}_{2}$, and $\mathcal{P}_{1}=\mathcal{P}_{2}$ since $\widehat{{\boldsymbol{F}}}_{i}$ and $\widehat{{\boldsymbol{g}}}_{i}$ can be induced by $\mathcal{P}_{i}$. Before introducing Proposition 3, a lemma is provided as preparation.
Lemma 2.
Consider the IVP $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}$, $t\in{\mathbb{R}}$, where ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, ${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$. If ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$, s.t. $\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$.
Proof.
According to Lemma 1, $\forall t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\left(\tau\right)\mathrm{d}\tau\right){\boldsymbol{b}}$. Considering the Jordan normal form of ${\boldsymbol{A}}$, it has $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{i=1}^{l}p_{i}\left(t-t_{0}\right){\boldsymbol{e}}^{\lambda_{i}\left(t-t_{0}\right)}$, where $\left\{\lambda_{i}\right\}_{i=1}^{l}$ are the eigenvalues of ${\boldsymbol{A}}$, and $p_{i}$ is a polynomial of degree not greater than $n$. Evidently, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}$ is an analytic function; hence, $\forall t\in{\mathbb{R}}$, it has
${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}.$ (18)
The case where ${\boldsymbol{A}}={\boldsymbol{0}}$ is trivial. Assume that ${\boldsymbol{A}}\not={\boldsymbol{0}}$. Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t. $p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$. Therefore, $\forall r\geq n$, $\exists p_{n}^{*}$ s.t. $p_{n}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{n}^{*}\left({\boldsymbol{A}}\right)$.
Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$. Since $\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, it holds that $\forall r\geq n+1$,
$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)$ (19)
$=$ ${\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$
By (18), $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right)$. ∎
Proposition 3.
Assume that $\forall i=1,2$, ${\mathcal{S}}_{i}$ occurs for $t\in\left[t_{i-1},t_{i}\right]$ in the above notation. ${\boldsymbol{x}}$ enters ${\mathcal{S}}_{2}$ from ${\mathcal{S}}_{1}$ at $t_{1}$.
  1. 1.
    If ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ are unconstrained arcs, then ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 at $t_{1}$. $\widehat{{\boldsymbol{b}}}_{1}=u_{1}{\boldsymbol{b}}$ and $\widehat{{\boldsymbol{b}}}_{2}=-u_{1}{\boldsymbol{b}}$, where $u_{1}\in\left\{\pm{u_{\mathrm{m}}}\right\}$.
  2. 2.
    If ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ are constrained arcs, then $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$. Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (20)
    $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (21)
    Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ for $\left(t_{1},t_{2}\right)$. If ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (20) for $p\in\mathcal{P}_{1}$, (21) for $p\in\mathcal{P}_{2}$, and (12) for $p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, then $\exists\varepsilon>0$, ${\mathcal{S}}_{1}$ occurs for $\left(t_{1}-\varepsilon,t_{1}\right)$ and ${\mathcal{S}}_{2}$ occurs for $\left(t_{1},t_{1}+\varepsilon\right)$.
  3. 3.
    If ${\mathcal{S}}_{1}$ is a constrained arc and ${\mathcal{S}}_{2}$ is an unconstrained arc, then $\forall p,q\in\mathcal{P}_{1}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{2}$, $\forall p\in\mathcal{P}_{1}$, it has
    $u\equiv u_{2}=-{\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right){u_{\mathrm{m}}}.$ (22)
    Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)=0.\end{dcases}$ (23)
    Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$ for $\left(t_{1},t_{2}\right)$ where $u_{2}$ is given in (22). If ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (23) and (12) for $p\in\mathcal{P}_{1}$, then $\exists\varepsilon>0$, ${\mathcal{S}}_{1}$ occurs for $\left(t_{1}-\varepsilon,t_{1}\right)$ and ${\mathcal{S}}_{2}$ occurs for $\left(t_{1},t_{1}+\varepsilon\right)$.
  4. 4.
    If ${\mathcal{S}}_{1}$ is an unconstrained arc and ${\mathcal{S}}_{2}$ is a constrained arc, then $\forall p,q\in\mathcal{P}_{2}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{1}$, $\forall p\in\mathcal{P}_{2}$, it has
    $u\equiv u_{1}=\left(-1\right)^{r_{p}-1}{\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right){u_{\mathrm{m}}}.$ (24)
    Furthermore, $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)=0.\end{dcases}$ (25)
    Conversely, assume that $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{1}{\boldsymbol{b}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{1},t_{2}\right)$ where $u_{1}$ is given in (24). If ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (25) and (12) for $p\in\mathcal{P}_{2}$, then $\exists\varepsilon>0$, ${\mathcal{S}}_{1}$ occurs for $\left(t_{1}-\varepsilon,t_{1}\right)$ and ${\mathcal{S}}_{2}$ occurs for $\left(t_{1},t_{1}+\varepsilon\right)$.
Proof.
Consider the connection between two unconstrained arcs. By Proposition 1, assume that $\forall i=1,2$, $u\equiv u_{i}$ on ${\mathcal{S}}_{i}$, where $u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Then, $\widehat{{\boldsymbol{b}}}_{i}=u_{i}{\boldsymbol{b}}$. Since $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}={\boldsymbol{A}}$ and $\mathcal{P}_{1}=\mathcal{P}_{2}=\varnothing$, it holds that $\widehat{{\boldsymbol{b}}}_{1}\not=\widehat{{\boldsymbol{b}}}_{2}$; hence, $u_{1}=-u_{2}$. Therefore, Proposition 3-1 holds.
Consider the connection between two constrained arcs. Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not=\varnothing$. Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration. According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}}}{{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}$ holds. Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$. By Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds. Therefore, $\mathcal{P}_{1}=\mathcal{P}_{2}$, which contradicts ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. Hence, $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. $\exists\varepsilon>0$, s.t. $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{d}}_{p}<0$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ and $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)$. Lemma 2 implies that $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$. Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, it has $\hat{r}_{p}\geq r_{p}$. In other words, (20) holds. For the same reason, (21) holds.
Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ for $\left(t_{1},t_{2}\right)$. Assume that ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (20) for $p\in\mathcal{P}_{1}$, (21) for $p\in\mathcal{P}_{2}$, and (12) for $p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$. Then, $\forall p\in\mathcal{P}_{1}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0$ and
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (26)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. For the same reason, $\exists\varepsilon^{\prime}\in\left(0,\varepsilon\right)$, $\forall p\in\mathcal{P}_{2}$, $t\in\left(t_{1}-\varepsilon^{\prime},t_{1}\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. Hence, Proposition 3-2 holds.
Consider the connection between an unconstrained arc and a constrained arc. Assume that ${\boldsymbol{x}}$ enters an unconstrained arc ${\mathcal{S}}_{2}$ from a constrained arc ${\mathcal{S}}_{1}$. By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{-}}=0$; hence,
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0,\\ {\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u\left(t_{1}^{-}\right){\boldsymbol{b}}\right)=0.\end{dcases}$ (27)
Assume that $u\equiv u_{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ on ${\mathcal{S}}_{2}$. Then, $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0$ since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$. Since $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$, it has $\left.\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)\leq 0$. Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\left(u_{2}-u\left(t_{1}^{-}\right)\right)\leq 0$. Since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$ and $\left|u\left(t_{1}^{-}\right)\right|\leq\left|u_{2}\right|={u_{\mathrm{m}}}$, (22) holds.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{2}$ where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}<0$ for $\left(t_{1},t_{1}+\varepsilon\right)$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$; hence, $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)$. Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, Lemma 2 implies that $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$ for $\left(t_{1},t_{2}\right)$. Assume that ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (23) and (12) for $p\in\mathcal{P}_{1}$. Then, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0$ and
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (28)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $t\in\left[t_{0},t_{1}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$. So Proposition 3-3 holds. Similarly, Proposition 3-4 holds. ∎
Remark.
Proposition 3 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection. Note that the equality constraints at the connection time, i.e., (12), (21), (23), and (25), are not independent since (12) implies the latter three equations for $1\leq r<r_{p}$. When examining feasibility near the connection time, it is sufficient to consider all inequality constraints and the maximal linearly independent set of the above equality constraints. Furthermore, Proposition 3 temporarily does not consider constraints on $u$ and ${\boldsymbol{x}}$, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ which will be discussed in Section IV-A.
IV Keypoints for the Feasibility of the Optimal Trajectory
The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, but it fails to fully present information on the feasibility during an arc. To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of the “keypoints” in an arc. Section IV-A analyzes some constraints of the end of an arc besides Proposition 3. The tangent marker is proposed in Section IV-B to describe the feasibility of the optimal trajectory tangent to the constrained boundary. Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
IV-A Additional Constraints at the End of An Arc
Proposition 3 fully discusses the feasibility of the connection of two adjacent arcs w.r.t. constraints induced by $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. However, the constraints on $u$ and ${\boldsymbol{x}}$, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ still lacks discussion. The following proposition provides a sufficient and necessary condition for the feasibility at the end of an arc.
Proposition 4.
Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
  1. 1.
    $\forall p\not\in\mathcal{P}$, $\exists\varepsilon>0$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\leq 0$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds:
    1. (a)
      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}<0$.
    2. (b)
      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$. $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\forall 1\leq r<\hat{r}_{p}$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
    Similar conclusions hold for ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}^{-}\right)+d_{p}\leq 0$.
  2. 2.
    Assume that ${\mathcal{S}}$ is a constrained arc. Then, $\exists\varepsilon>0$, s.t. $u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds, where $p\in\mathcal{P}$:
    1. (a)
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<{u_{\mathrm{m}}}$.
    2. (b)
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\forall 1\leq r<\hat{r}_{p}$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
    3. (c)
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$, and $\forall 1\leq r\leq n$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
    Similar conclusions hold for $\left|u\left(t_{0}^{+}\right)\right|,\left|u\left(t_{1}^{-}\right)\right|\leq{u_{\mathrm{m}}}$.
Proof.
Proposition 4-1 holds for the same reason as that of Proposition 3.
Consider Proposition 4-2. Proposition 2 implies that $u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}$. $\exists\varepsilon>0$, s.t. $u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if (1) $\exists\varepsilon>0$, s.t. $u<{u_{\mathrm{m}}}$ for $\left(t_{0},t_{0}+\varepsilon\right)$, or (2) $\exists\varepsilon>0$, s.t. $u\left(t\right)\equiv{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$. By Proposition 2, (1) is equivalent to $\exists 0\leq\hat{r}_{p}\leq n$, s.t. $\forall 0\leq r<\hat{r}_{p}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$. (2) is equivalent to $\forall 0\leq r\leq n$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Hence, Proposition 4-2 holds. ∎
Remark.
Proposition 4 encompasses discussions on the cases covered by Proposition 3, while Proposition 3 provides more details for $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
Based on Proposition 4, the additional end-constraint is defined as follows:
Definition 3.
Assume that ${\mathcal{S}}$ is a given arc for $t\in\left[t_{0},t_{1}\right]$. At an end of ${\mathcal{S}}$, i.e., $t_{i}$, $i\in\left\{0,1\right\}$, ${\mathcal{S}}$ and its adjacent arc, if it exists, correspond to the active constraints $\mathcal{P}$ and $\mathcal{P}^{\prime}$. $\mathcal{P}$ and $\mathcal{P}^{\prime}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{i}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank. The additional end-constraint of ${\boldsymbol{x}}\left(t_{i}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$. Among them, $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by Proposition 4 for the active constraints of ${\mathcal{S}}$ and its adjacent arc at $t_{i}$, except for those in $\mathcal{P}\cup\mathcal{P}^{\prime}$. $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
Remark.
Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 4 is directly determined. In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}\cup\mathcal{P}^{\prime}$. To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}\cup\mathcal{P}^{\prime}$ are made linearly independent through elimination. The elimination process is guaranteed by the feasibility of the given trajectory.
IV-B Tangent Markers in An Arc
Proposition 3 and Proposition 4 analyzed the feasibility of an arc near the end points. An arc might be tangent to some constrained boundary, where light perturbations may affect the feasibility of the arc. The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to the constrained boundary.
Proposition 5.
Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is an arc for $t\in\left[t_{0},t_{2}\right]$. ${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists 1\leq\hat{r}\leq n$ is even, s.t.
$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall 1\leq r<\hat{r},\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
Proof.
${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0$. $\exists\varepsilon>0$, s.t. $B_{\varepsilon}\left(t_{1}\right)\subset\left(t_{0},t_{2}\right)$ and $\forall t\in{\left.B_{\varepsilon}\left(t_{1}\right)\middle\backslash\left\{t_{1}\right\}\right.}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d<0$. Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)={\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}\right)$.
Assume that $\forall 1\leq r\leq n$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$. By Lemma 2, $\forall t\in B_{\varepsilon}\left(t_{1}\right)$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d\equiv 0$, which contradicts the tangent condition. Hence, $\exists 1\leq\hat{r}\leq n$, s.t. $\forall 1\leq r<\hat{r}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
If $\hat{r}$ is odd, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ strictly crosses 00 at $t_{1}$, which contradicts the tangent condition. Hence, $\hat{r}$ is even. If $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}>0$, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ achieves a strict local minimum at $t_{1}$, which contradicts the maximum condition. Hence, $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}<0$. Therefore, (29) holds.
Conversely, assume that (29) holds. Considering the Taylor of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ at $t_{1}$, it holds that ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d<0$ at a deleted neighborhood of $t_{1}$. Hence, ${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ at $t_{1}$. ∎
No matter the arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is constrained or unconstrained, ${\mathcal{S}}$ can be tangent to the boundary of the state constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$. It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$. For this case, ${\mathcal{S}}$ should be constrained. Arbitrarily consider $p\in\mathcal{P}$. According to Proposition 2, ±$\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$. In other words, the constraints on $u$ are equivalent to the constraints on ${\boldsymbol{x}}$. Hence, Proposition 5 can apply for both state constraints and control constraints. For convenience of discussion, $\forall 1\leq p\leq P$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The $\left(P+1\right)$-th constraint refers to $u\leq{u_{\mathrm{m}}}$, and the $\left(P+2\right)$-th constraint refers to $-u\leq{u_{\mathrm{m}}}$.
Another noteworthy fact is that an arc ${\mathcal{S}}$ can be tangent to multiple constraints $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$ at a time point $t_{1}$. Then, the necessary and sufficient condition for feasibility is $\forall p\in\mathcal{P}^{\prime}$, (29) holds. The tangent marker is defined as follows:
Definition 4.
Assume that $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$, an arc ${\mathcal{S}}$ is tangent to the $p$-th constraint at $t_{1}$. Denote the tangent marker as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$. $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by (29), where $\widehat{{\boldsymbol{C}}}$ has full row rank.
Remark.
In Definition 4, the equality constraints induced by (29) can be linearly dependent. However, for a given feasible trajectory, the equalities should have a feasible solution; hence, equality constraints with full row rank, i.e., $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, can be constructed through elimination. Therefore, the tangent marker is well-defined.
IV-C Augmented Switching Law
Section III, Section IV-A, and Section IV-B analyze the feasibility of the optimal trajectory at the end points of arcs and the tangent points to the constrained boundaries. To characterize the optimal trajectory, this section proposes the augmented switching law, which represents the input control and the feasibility of the trajectory in a compact form.
Definition 5.
Assume that the optimal trajectory of problem (1a) satisfies Assumption 1; hence, it contains a finite number of arcs. The augmented switching law is defined as ${\mathcal{L}}={\mathcal{E}}_{0}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{\left(1\right)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{\left(N\right)}\right\}_{k=1}^{M_{N}}\right){\mathcal{E}}_{N}$. $\forall i\in\left[N\right]$, ${\mathcal{S}}_{i}$ occurs on $\left[t_{0}^{\left(i\right)},t_{0}^{\left(i+1\right)}\right]$. $\forall k\in\left[M_{i}\right]$, the tangent marker ${\mathcal{T}}_{k}^{\left(i\right)}$ occurs at $t_{k}^{\left(i\right)}$ during ${\mathcal{S}}_{i}$, where $t_{0}^{\left(i\right)}<t_{1}^{\left(i\right)}<t_{2}^{\left(i\right)}\dots<t_{M_{i}}^{\left(i\right)}<t_{0}^{\left(i+1\right)}$. The additional end-constraint ${\mathcal{E}}_{i}$ is induced by Proposition 4 at $t_{0}^{\left(i\right)}$. Among them, ${\mathcal{E}}_{N}$ additionally contains the constraints ${\boldsymbol{x}}\left(t_{0}^{\left(N+1\right)}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$, where $t_{0}^{\left(N+1\right)}={t_{\mathrm{f}}}$. $\forall i\in\left[N+1\right]$, $k\in\left[M_{i}\right]\cup\left\{0\right\}$, the points $t_{k}^{\left(i\right)}$ are called keypoints for the feasibility of the trajectory.
Remark.
If a feasible trajectory with a finite number of arcs follows the Bang-Singular-Bang control law (7), then Proposition 3, Proposition 4, and Proposition 5 hold for the trajectory. For this case, the augmented switching law can also represent the feasible trajectory.
Remark.
If a non-chattering Bang-Singular-Bang feasible trajectory is disturbed by a sufficiently small perturbation, then the disturbed trajectory can only exceed the constrained boundary at the keypoints since the distance between the original trajectory and the constrained boundary is 0 at the keypoints, while it is strictly positive at other points. Therefore, the augmented switching law can not only characterize the input control but also describe the feasibility of the trajectory.
Remark.
The switching law in Definition 2 corresponding to the augmented switching law in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$. It can be observed that the augmented switching law provides more detailed information on the feasibility of the trajectory at keypoints than the switching law. Notably, neither the switching law nor the augmented switching law provides the motion time of each arc.
To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an augmented switching law. If no additional end-constraint is induced at the end of an arc, i.e., ${\mathcal{E}}=\left(\sim,\sim,\sim,\sim\right)$, then ${\mathcal{E}}$ is omitted.
For convenience of understanding, the chain-of-integrators system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [8, 18]. The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively. The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively. Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall 1\leq j<k$, $x_{j}\equiv 0$. The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of an even order $2l<k$ in Proposition 5 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively. The additional end-constraint ${\mathcal{E}}=\left(\left(s_{k},r_{k}\right)\right)_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 4, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively. Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc. Based on the above simplified notations, an example is provided to illustrate the augmented switching law.
Example 3.
Consider a position-to-position problem for 4th order chain-of-integrators systems with full box constraints shown in Fig. 1. A feasible suboptimal trajectory with a Bang-Singular-Bang form, planned by the manifold-intercepted method (MIM) [8], is shown in Fig. 1(a). It can be observed that the augmented switching law is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
Figure 1: Two examples of augmented switching laws. For a 4th order chain-of-integrators system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$. (a) A non-chattering feasible trajectory planned by our previous work [8]. (b) A chattering optimal trajectory planned by our previous work [18]. In (b2), $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
Example 4.
Our previous work [18] pointed out that the chattering phenomenon occurs in the optimal trajectory for the problem in Example 3. The chattering optimal trajectory is shown in Fig. 1(b1). Fig. 1(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied. It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$. According to Bellman’s principle of optimality [20], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is small enough. The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
V A State-Centric Necessary Condition for Optimal Control of Problem (1a)
The theoretical framework for the augmented switching law of time-optimal control for linear systems is established in Section IV. Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a). Section V-A discusses the uniqueness of the time-optimal control. Then, given a feasible Bang-Singular-Bang trajectory, Section V-B analyzes the feasibility of the disturbed Bang-Singular-Bang trajectory. Based on the conclusions in Section V-A and Section V-B, Section V-C provides a necessary condition for the optimal trajectory represented by the augmented switching law, which constrains the difference between the switching times and the sum of orders of keypoints in a way.
V-A Uniqueness of Time-Optimal Control
The uniqueness of the time-optimal control for problem (1a) is proved in Theorem 1 as preliminaries.
Theorem 1.
Under Assumption 1, the time-optimal control of problem (1a) is unique in an almost everywhere sense. In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal controls of problem (1a), then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
Proof.
If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then evidently it has $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere since $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{0}$.
Assume that $\exists t^{\prime}\in\left(0,t_{\mathrm{f}}^{*}\right)$, s.t. ${\boldsymbol{x}}_{1}^{*}\left(t^{\prime}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t^{\prime}\right)$. Let $t_{0}=\arg\min\left\{t\in\left(0,t_{\mathrm{f}}^{*}\right):{\boldsymbol{x}}_{1}^{*}\left(t\right)\not={\boldsymbol{x}}_{2}^{*}\left(t\right)\right\}$. Evidently, $0\leq t_{0}<t_{\mathrm{f}}^{*}$. Assume that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$. Then, $t_{0}>0$. Due to the continuity of ${\boldsymbol{x}}$, $\forall 0\leq t<t_{0}$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$, which contradicts the assumption. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ holds. Furthermore, $\forall t\in\left[0,t_{0}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ for $t\in\left[0,t_{0}\right]$ almost everywhere.
According to Bellman’s principle of optimality [20], ${\boldsymbol{x}}={\boldsymbol{x}}_{1}^{*}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}_{2}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ are still the time-optimal trajectories between ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{\mathrm{f}}^{*}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{\mathrm{f}}^{*}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. Let $u_{3}^{*}\left(t\right)=\frac{1}{2}\left(u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)\right)$ and ${\boldsymbol{x}}_{3}^{*}\left(t\right)=\frac{1}{2}\left({\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)\right)$. Evidently, $u_{3}^{*}\left(t\right)$ is feasible for problem (1a) since the feasible set of $\left(u,{\boldsymbol{x}}\right)$ is convex; hence, $u_{3}^{*}\left(t\right)$ is optimal.
Denote that $\forall k=1,2$, the switching law of ${\boldsymbol{x}}={\boldsymbol{x}}_{i}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ is ${\mathcal{L}}_{i}={\mathcal{S}}_{1}^{\left(i\right)}{\mathcal{S}}_{2}^{\left(i\right)}\dots\mathcal{S}_{N_{i}}^{\left(i\right)}$, where $\forall j\in\left[N_{i}\right]$, ${\mathcal{S}}_{j}^{\left(i\right)}=\left(\widehat{{\boldsymbol{A}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{b}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{F}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{g}}}_{j}^{\left(i\right)},\mathcal{P}_{j}^{\left(i\right)}\right)$ lasts for $t\in\left(t_{j-1}^{\left(i\right)},t_{j}^{\left(i\right)}\right)$. Among them, $t_{0}^{\left(i\right)}=t_{0}$ and $t_{N_{i}}^{\left(i\right)}=t_{\mathrm{f}}^{*}$. Evidently, if ${\mathcal{S}}_{1}^{\left(1\right)}={\mathcal{S}}_{1}^{\left(2\right)}$, then $\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Hence, ${\mathcal{S}}_{1}^{\left(1\right)}\not={\mathcal{S}}_{1}^{\left(2\right)}$.
Assume that both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are unconstrained arcs. Without loss of generality, assume that $u_{1}^{*}\left(t\right)\equiv{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(1\right)}$; hence, $u_{2}^{*}\left(t\right)\equiv-{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(2\right)}$. Then, $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$ and ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Therefore, ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the fact that $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\equiv 0$. Therefore, one of ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ is constrained. Without loss of generality, assume that ${\mathcal{S}}_{1}^{\left(1\right)}$ is constrained.
Assume that ${\mathcal{S}}_{1}^{\left(2\right)}$ is an unconstrained arc. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Since ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}}$, it has ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Note that $u_{1}^{*}\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ and $u_{2}^{*}\left(t\right)\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{1}^{*}\left(t\right)\equiv u_{0}$. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ for $\left[t_{0},t_{0}+\varepsilon\right]$, which contradicts the definition of $t_{0}$.
Therefore, both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are constrained arcs. Assume that $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}\not=\varnothing$. Arbitrarily consider $p\in\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}$. $\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$. ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Therefore, $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}=\varnothing$. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{\left(i\right)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$. Hence, $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{3}^{*}\left(t\right)+d_{p}={\boldsymbol{c}}_{p}^{\top}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+u{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+d_{p}<0$. According to Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$. Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere. ∎
Remark.
Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an almost everywhere sense when there exists at most one chattering limit point. The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
According to Theorem 1, if the time-optimal trajectory is disturbed by a sufficiently small perturbation, then the resulting disturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time. The above observation will be applied in Section V-B.
V-B Feasibility of the Disturbed Bang-Singular-Bang Trajectory
[…]
The feasibility of the disturbed trajectory is proved by two steps. Firstly, the consistent error between the disturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory. In this way, the disturbed trajectory can be proved feasible at the points far from keypoints. Secondly, the feasibility of the disturbed trajectory near the keypoints is proved based on Proposition 3, Proposition 4, and Proposition 5. As a result, the feasibility of the disturbed trajectory is proved. The two steps correspond to Proposition 6 and Theorem 2, respectively.
Proposition 6.
${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$. $\exists C>0$, s.t. $\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$,
$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon,$ (30a)
$\max_{i\in\left[M\right]}\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|\leq C\varepsilon.$ (30b)
Among them, $\forall i\in\left[M\right]$, $t\in\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$. ${\boldsymbol{x}}^{\prime}\left(t_{0}\right)={\boldsymbol{x}}\left(t_{0}\right)$ and $t_{0}^{\prime}=t_{0}$. ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$ and ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$. $0<\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$.
Proof.
Since $\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$ and $\left|t_{i}-t_{i}^{\prime}\right|\leq\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, it has $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$. $\forall i\in\left[M\right]$, let $\hat{t}_{2i-1}=\min\left\{t_{i},t_{i}^{\prime}\right\}$, $\hat{t}_{2i}=\max\left\{t_{i},t_{i}^{\prime}\right\}$, and $\hat{t}_{0}=0$. Then $\left\{\hat{t}_{i}\right\}_{i=0}^{2M}$ increases monotonically.
Note that $\forall{\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\mathrm{e}}^{{\boldsymbol{A}}t}$ is locally Lipschitz continuous w.r.t. $t$. Therefore, $\exists C_{1}>0$, s.t. $\forall t\in\left[0,t_{M}-t_{0}\right]$, $\max_{i\in\left[M\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}t}-{\boldsymbol{I}}\right\|\leq C_{1}t$. Let $C_{2}=\max_{i\in\left[M\right]}\left\|{\boldsymbol{b}}_{i}\right\|$ and $C_{3}=\max_{i\in\left[M\right]}\sup_{t\in\left[0,t_{M}-t_{0}\right]}\left\|{\boldsymbol{e}}^{{\boldsymbol{A}}_{i}t}\right\|\geq 1$.
By Lemma 1, $\forall i\in\left[M-1\right]\cup\left\{0\right\}$, $t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$,
$\begin{dcases}{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau,\\ {\boldsymbol{x}}^{\prime}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.\end{dcases}$ (31)
Therefore,
$\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (32)
$=$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\left({\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right)\right\|$
$\leq$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$
$\leq$ $C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|.$
$\forall i\in\left[M\right]$, let
$\begin{dcases}j=i+1,\,j^{\prime}=i,&\text{if }t_{i}\leq t_{i}^{\prime},\\ j=i,\,j^{\prime}=i+1,&\text{if }t_{i}>t_{i}^{\prime}.\end{dcases}$ (33)
Then, $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{j}{\boldsymbol{x}}+{\boldsymbol{b}}_{j}$ and $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{j^{\prime}}{\boldsymbol{x}}+{\boldsymbol{b}}_{j^{\prime}}$. Lemma 1 implies that $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$,
$\begin{dcases}{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)+\int_{\hat{t}_{2i-1}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\tau\right)}{\boldsymbol{b}}_{j}\mathrm{d}\tau,\\ {\boldsymbol{x}}^{\prime}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)+\int_{\hat{t}_{2i-1}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\tau\right)}{\boldsymbol{b}}_{j^{\prime}}\mathrm{d}\tau.\end{dcases}$ (34)
Therefore,
$\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (35)
$\leq$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\varepsilon\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j^{\prime}}\right\|$
$\leq$ $2\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon+C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|.$
Since ${\boldsymbol{x}}\left(\hat{t}_{0}\right)={\boldsymbol{x}}^{\prime}\left(\hat{t}_{0}\right)$, (32) and (35) implies that
$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\leq 2\left(C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$ (36)
Therefore, (30a) holds. Similarly, $\forall i\in\left[M\right]$,
$\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|$ (37)
$\leq$ $\left\|{\boldsymbol{x}}\left(t_{i}^{\prime}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|+\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}\left(t_{i}^{\prime}\right)\right\|$
$\leq$ $\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+C_{1}\left\|{\boldsymbol{x}}\right\|\varepsilon+C_{2}C_{3}\varepsilon$
$\leq$ $\left(2C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$
Therefore, (30b) holds. ∎
Theorem 2.
${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are Bang-Singular-Bang trajectories of problem (1a) without chattering. Assume that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are represented by the same augmented switching law ${\mathcal{L}}={\mathcal{L}}^{\prime}$. ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is feasible for problem (1a). $\exists\varepsilon>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
Proof.
Fix ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$. Firstly, let
$\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}.$ (38)
According to Proposition 6, $\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then (30a) holds.
Secondly, consider the inequality constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\leq 0$. Denote that $\left\{i^{\left(p\right)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t. $i^{\left(p\right)}_{0}=0$, $i^{\left(p\right)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds. If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ since ${\mathcal{L}}={\mathcal{L}}^{\prime}$.
Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$. Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$, and then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{\left(p\right)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{\left(p\right)}}\right)$ is continuous in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$. If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i^{\left(p\right)}_{j-1}}\right)+d_{p}<0$, then let $r_{j}^{\left(p\right)}=0$; otherwise, according to Proposition 3, Proposition 4, and Proposition 5, $\exists 1\leq r_{j}^{\left(p\right)}\leq n$, s.t. $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$. Let $t_{1,j}^{\left(p\right)}>t_{i^{\left(p\right)}_{j-1}}$ is small enough, s.t. $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)\leq\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+1$ for $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$. $t_{2,j}^{\left(p\right)}<t_{i^{\left(p\right)}_{j}}$ is constructed similarly. Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}\right)$. In both case, $\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ due to continuity. Let
$\varepsilon<-\frac{\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}}{2C\left\|{\boldsymbol{c}}_{p}\right\|}.$ (39)
Then, $\forall t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$,
${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}<0.$ (40)
In other words, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible in $\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$.
The feasibility of ${\boldsymbol{x}}^{\prime}$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)\cup\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ needs further discussion. Let
$\begin{dcases}\varepsilon<t_{1,j}^{\left(p\right)}-t_{i^{\left(p\right)}_{j-1}},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}+1\right)\right)^{-1},\\ \varepsilon<\left(C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}+1}+1\right)\right)^{-1},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+8\right)^{-1}.\end{dcases}$ (41)
Denote $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}$. Then, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right]$,
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)$
$=$ $\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
$\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}-\varepsilon,t_{i^{\left(p\right)}_{j-1}}\right]$, it has
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (43)
$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)+1}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$
$\leq$ $\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2.$
Therefore, if $t_{i^{\left(p\right)}_{j-1}}^{\prime}\geq t_{i^{\left(p\right)}_{j-1}}$, then $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$. If $t_{i^{\left(p\right)}_{j-1}}^{\prime}<t_{i^{\left(p\right)}_{j-1}}$, then (30b) implies that $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\leq\frac{3}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$. $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j-1}}\right]$,
$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (44)
$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)+\left(\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2\right)\varepsilon$
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$.
Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)$. Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$. Similarly, let $\varepsilon$ is small enough, and it has ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ in $\left[t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right]$.
In summary, when $\varepsilon$ is small enough, $\forall p\in\left[P\right]$, $t\in\left[t_{0}^{\prime},t_{M}^{\prime}\right]$, it has ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ holds.
Thirdly, the inequality $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is small enough. ∎
Theorem 2 points out that if the disturbed trajectory is represented by the same augmented switching law to the original trajectory and the disturbance is light enough, then the disturbed trajectory is feasible. In other words, Theorem 2 provides a novel variational approach under fixed augmented switching laws with a feasibility guarantee near the constrained boundary.
V-C Necessary Condition for the Optimal Trajectory
Based on Theorem 2, this section provides an approach to constructing a feasible disturbed trajectory with the same augmented switching law. According to Theorem 1, this section shows that the existence of such a disturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
The Jacobian matrix of equality constraints induced by the augmented switching law can be calculated by Proposition 7.
Proposition 7.
${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$, and ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ is fixed. Consider the mapping ${\boldsymbol{x}}_{i}:{\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}\mapsto{\boldsymbol{x}}\left(t_{i}\right)$. Then, $\forall i,j\in\left[M\right]$,
$\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}=\begin{dcases}\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\\ \qquad j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ 0,\,j>i,\end{dcases}$ (45)
where $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}$ refers to
${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}.$ (46)
Proof.
Evidently, $\forall j>i$, ${\boldsymbol{x}}_{i}$ is independent with $t_{j}$; hence, $\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}={\boldsymbol{0}}$. According to Lemma 1,
${\boldsymbol{x}}_{i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\boldsymbol{x}}_{i-1}+\int_{t_{i-1}}^{t_{i}}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.$ (47)
Therefore, $\forall i\in\left[M\right]$,
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}.$ (48)
Applying (48), $\forall i>1$, it has
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i-1}}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}\left(\frac{\partial{\boldsymbol{x}}_{i-1}}{\partial t_{i-1}}-{\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i-1}-{\boldsymbol{b}}_{i}\right)$ (49)
$=$ ${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}\left(\left({\boldsymbol{A}}_{i-1}-{\boldsymbol{A}}_{i}\right){\boldsymbol{x}}_{i-1}+{\boldsymbol{b}}_{i-1}-{\boldsymbol{b}}_{i}\right).$
$\forall i>1$, $j<i-1$, it has
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ (50)
$=$ $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right).$
Therefore, (45) holds. ∎
Utilizing the notations in Section V-B, the augmented switching law provides a system of equalities on every keypoints. Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank. For a constrained arc with active constraints $\mathcal{P}$, the equality constraints induced by $\mathcal{P}$ are added in only one end instead of both ends of the arcs since one end satisfies $\mathcal{P}$ implies that the whole arc satisfies $\mathcal{P}$. Note that the additional end-constraints at both ends are eliminated by equality constraints induced by $\mathcal{P}$. Similar processes are applied if $u\equiv{u_{\mathrm{m}}}$ or $u\equiv-{u_{\mathrm{m}}}$ during a constrained arc. ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$. Then, Theorem 3 provides a necessary condition for the optimal trajectory.
Before providing Theorem 3, the implicit function theorem [29] should be introduced.
Lemma 3 (The Implicit Function Theorem [29]).
Assume $S\subset{\mathbb{R}}^{p+q}$ is open and non-empty. ${\boldsymbol{F}}:S\to{\mathbb{R}}^{q}$ is $C^{1}$ continuous. $\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\in S$, satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)={\boldsymbol{0}}$ and $\det\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\not=0$. Then, $\exists\delta_{1},\delta_{2}>0$, s.t. $\forall{\boldsymbol{\xi}}\in B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)$, $\exists!{\boldsymbol{\eta}}\in B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$ satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}},{\boldsymbol{\eta}}\right)={\boldsymbol{0}}$. The above relation induces a mapping ${\boldsymbol{f}}:B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)\to B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$. Then, ${\boldsymbol{f}}$ is $C^{1}$ continuous, where $\frac{\mathrm{d}{\boldsymbol{f}}}{\mathrm{d}{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}}\right)=-\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)^{-1}\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)$.
Theorem 3 (State-Centric Necessary Condition).
Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1. Denote the equality constraints induced by the augmented switching law as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$. Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
Proof.
Assume that ${\boldsymbol{H}}:{\mathbb{R}}^{M}\to{\mathbb{R}}^{M^{\prime}}$. $M^{\prime}$ is the number of linearly independent equality constraints on keypoints, and $M$ is the number of keypoints. Assume that $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ has full row rank, i.e.,
${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)=M^{\prime}\leq M-1.$ (51)
Assume that the $i$-th row of $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ is linearly independent where $i\in\mathcal{I}\subset\left[M-1\right]$ and #$\#\mathcal{I}=M^{\prime}$. Let $\varepsilon>0$ be small enough to satisfy the condition of Theorem 2 and Lemma 3. The implicit function ${\boldsymbol{f}}:\left(t_{i}\right)_{i\in{\left.\left[M\right]\middle\backslash\mathcal{I}\right.}}\mapsto\left(t_{i}\right)_{i\in\mathcal{I}}$ is induced by ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)={\boldsymbol{0}}$. Based on ${\boldsymbol{f}}$, $\exists{\boldsymbol{t}}^{\prime}$ satisfies ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$, $t_{M}^{\prime}<t_{M}^{*}$, and $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}^{*}\right\|<\varepsilon$. According to Theorem 2, the disturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible since ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$ implies that ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{*}$ have the same augmented switching law. However, ${\boldsymbol{x}}^{\prime}$ achieves a shorter terminal time $t_{M}^{\prime}<t_{M}^{*}$, which contradicts the optimality of ${\boldsymbol{x}}^{*}$. ∎
Remark.
[…]
Remark.
In fact, if a given non-chattering Bang-Singular-Bang trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical local optimization of trajectories based on Theorem 3. For example, let $t_{M}$ be the optimization objective and consider the equality and inequality constraints of the augmented switching law as the constraints of the optimization problem. In this way, it is hopeful to obtain a locally optimal solution or a stationary point of problem (1a) through iteratively solving an $M$-dimensional nonlinear problem. Note that when applying shooting methods [30], problem (1a) is highly nonlinear and non-convex, and the long motion time may result in unacceptable computation cost. In contrast, the local optimization method on ${\boldsymbol{t}}$ based on Theorem 1a will be more efficient due to the significantly lower dimensions and the suboptimal feasible initial value. The above method will be developed in future works.
Theorem 3 provides a necessary condition for the optimal control of problem (1a) under Assumption 1. Different from the existing costate-based necessary conditions [14], Theorem 3 requires no information on the costates, which significantly simplifies the computation of the necessary conditions. Compared to existing state-centric necessary conditions [22], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
Some further discussions on the Jacobian matrix are necessary. Consider a keypoint $t_{i}$ with a constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ active at $t_{i}$. Assume that the constraint is of $\hat{r}$-th order w.r.t the arc $\left(t_{i-1},t_{i}\right)$ where $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$. In other words, if $1\leq\hat{r}<n$, then $\left(-1\right)^{\hat{r}}\frac{\mathrm{d}^{\hat{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{\hat{r}}}\left(t_{i}^{-}\right)>0$ and $\forall 1\leq r<\hat{r}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$. If $\hat{r}=n$, then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$. Among them. $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)$. According to Proposition 7, $\forall r\in{\mathbb{N}}$,
$\frac{\partial}{\partial t_{i}}\left.\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\right|_{t_{i}^{-}}=\frac{\partial}{\partial t_{i}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)$ (52)
$=$ $\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)=\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}.$
In other words, the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t. $t_{i}$ is $\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}{\boldsymbol{e}}_{\hat{r}}$. Specifically, if $\hat{r}=n$, then the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t. $t_{i}$ is ${\boldsymbol{0}}$.
VI Applications in Chain-of-Integrators Systems and Simulation Experiments
This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a). For convenience, the chain-of-integrators system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
Time-optimal control for chain-of-integrators systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved. The problem can be summarized as follows:
$\min\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (53a)
${\mathrm{s.t.}}\quad$ $\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),\,\forall 1<k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53b)
$\dot{x}_{1}\left(t\right)=u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (53c)
${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}},$ (53d)
$\left|x_{k}\left(t\right)\right|\leq x_{{\mathrm{m}}k},\,\forall 1\leq k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53e)
$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (53f)
The Jacobian matrix, except the last column induced by the augmented switching law, denoted as ${\boldsymbol{J}}$, can be calculated as follows:
Corollary 1.
Assume that $\forall i\in\left[N\right]$, $u\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}},0\right\}$ in $\left(t_{i-1},t_{i}\right)$, and denote ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$. Then,
$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}-\Delta u_{j+1}{\boldsymbol{\phi}}\left(t_{i}-t_{j}\right),&j<i\\ \left[\begin{array}[]{c}u_{i}\\ {\boldsymbol{x}}_{i,1:\left(n-1\right)}\end{array}\right],&j=i\\ {\boldsymbol{0}},&j>i.\end{dcases}$ (54)
Among them, $\Delta u_{i}=u_{i}-u_{i-1}$ and $u_{0}=0$. ${\boldsymbol{\phi}}\left(t\right)\triangleq\left(\phi_{k}\left(t\right)\right)_{k=1}^{n}$ where $\phi_{k}\left(t\right)=\frac{t^{k-1}}{\left(k-1\right)!}$.
Proof.
Proposition 7 directly implies Corollary 1. ∎
This section provides some applications of Theorem 3. Section VI-A proves a trivial conclusion from a state-centric perspective. Section VI-B and Section VI-C prove two corollaries that are challenging to prove by traditional costate-based necessary conditions.
VI-A The Case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$
The case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$ has a well-known conclusion that the optimal control switches no more than $\left(n-1\right)$ times [19]. This section proves the above conclusion based on the proposed Theorem 3.
Corollary 2.
In problem (53a), assume that $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$. Then, the optimal control switches no more than $\left(n-1\right)$ times.
Proof.
Chattering does not occur since no state inequality constraints exist. Assume that the optimal control switches $\left(N-1\right)$ times. In other words, $\forall i\in\left[N\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $u\left(t\right)\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. $u_{i}\not=u_{i-1}$, implies that $\Delta u_{i}\not=0$. Then,
${\boldsymbol{J}}=\left[-\Delta u_{i+1}{\boldsymbol{\phi}}\left(t_{N}-t_{i}\right)\right]_{i=1}^{N-1}\in{\mathbb{R}}^{n\times\left(N-1\right)}.$ (55)
The Vandermonde form of ${\boldsymbol{J}}$ implies that ${\mathrm{rank}}{\boldsymbol{J}}=\min\left\{n,N-1\right\}$. Theorem 3 implies that ${\mathrm{rank}}{\boldsymbol{J}}<n$; hence, $N-1<n$. ∎
Corollary 2 implies that the switching law is in the form of $\overline{0}\underline{0}\overline{0}\underline{0}\dots$ or $\underline{0}\overline{0}\underline{0}\overline{0}\dots$ with no more than $n$ arcs. The case in Section VI-A is trivial since no state constraints exist and the costate vector follows a fixed equation that $\forall k\in\left[n-1\right]$, $\dot{\lambda}_{i}=-\lambda_{i+1}$ and $\lambda_{n}\equiv{\mathrm{const}}$.
However, when the state constraints are induced, the behavior of the optimal control can be complex. On the one hand, the multiplier ${\boldsymbol{\eta}}$ in (4) can be non-zero during a constrained arc. On the other hand, the junction of ${\boldsymbol{\lambda}}$ can occur when a state constraint switches between active and inactive [7]. For this case, the costate analysis can be complex, while the proposed state-centric necessary condition provides a simple and effective approach for analyzing the optimal control.
VI-B The Case where Only System Behaviors Occur
The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to the constrained boundary, is widely applied in existing works on trajectory planning [31, 32, 8].
Denote the augmented switching law of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. $\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$. The non-existence of tangent markers and additional end-constraints in ${\mathcal{L}}$ means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility. An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
Corollary 3.
Assume that the augmented switching law consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
  1. 1.
    $\forall 1\leq j<i$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=j+1}^{i}\left|{\mathcal{S}}_{k}\right|<i-j$.
  2. 2.
    $\forall i<j\leq N$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{j-1}\left|{\mathcal{S}}_{k}\right|<j-i$.
  3. 3.
    $\forall i\in\left[N\right]$, if $\left|{\mathcal{S}}_{i}\right|>0$ and $\forall i<j\leq N$, $\left|{\mathcal{S}}_{j}\right|<\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{N}\left|{\mathcal{S}}_{k}\right|\leq N-i$.
Then,
$N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|\leq n.$ (56)
Proof.
For a constrained arc ${\mathcal{S}}_{i}$, add the equality constraint at $t_{i}$, i.e., ${\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|}$. Note that if ${\mathcal{S}}_{i}$ is constrained, i.e., $\left|{\mathcal{S}}_{i}\right|\not=0$, then $u\equiv 0$ in ${\mathcal{S}}_{i}$. Therefore, the constraints induced by ${\mathcal{L}}$ are
$\begin{dcases}{\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|},\,\forall i\in\left[N\right],\,\left|{\mathcal{S}}_{i}\right|\not=0,\\ {\boldsymbol{x}}_{N}={{\boldsymbol{x}}_{\mathrm{f}}}.\end{dcases}$ (57)
By Condition 3, $\left|{\mathcal{S}}_{N}\right|=0$. Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$. According to Corollary 1, ${\boldsymbol{J}}$ contains the following rows:
${\boldsymbol{J}}=\left(-\Delta u_{j+1}{\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]},$ (58)
where $\delta_{A}$ is the indicator function of condition $A$, i.e.,
$\delta_{A}=\begin{dcases}1,&\text{condition }A\text{ holds},\\ 0,&\text{condition }A\text{ does not hold}.\end{dcases}$ (59)
Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that ${\boldsymbol{A}}$ has full row rank if and only if ${\boldsymbol{B}}$ has full row rank. Since $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it has
${\boldsymbol{J}}\leftrightarrow\left({\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}.$ (60)
By iteratively taking differences between adjacent rows of (60), dividing by the time difference, and eliminating, it can be proved that ${\boldsymbol{J}}$ has full rank based on Conditions 1, 2, and 3. The proof is similar to our previous work [8] and is omitted due to space limitation. According to Theorem 3, the number of rows should be greater than the number of columns, i.e.,
$\sum_{i\in\mathcal{I}}\left|{\mathcal{S}}_{i}\right|=\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|+n>N-1.$ (61)
Therefore, (56) holds. ∎
Remark.
Further research suggests that Conditions 1, 2, and 3 in Corollary 3 might hold almost everywhere. In other words, $\left({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{\mathrm{f}},{\boldsymbol{x}}_{\mathrm{m}},{u_{\mathrm{m}}}\right)$ whose resulting Bang-Singular-Bang trajectory does not satisfy Conditions 1, 2, and 3 forms a zero-measure set.
Remark.
In our previous work [8], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the augmented switching law ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$ locally forms a submanifold of $r$-dimensions. Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$. However, the local optimality is not discussed in [8]. Corollary 3 proves that the augmented switching law of dimensions $r>n$ fails to achieve optimality.
${\boldsymbol{J}}=\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
$\left[\begin{array}[]{cc}\left(\delta_{j=i}-\delta_{j=i-1}\right)_{i\in\left[N\right],j\in\left[N\right]}&\\ &{\boldsymbol{I}}_{n-1}\end{array}\right]\left[\begin{array}[]{cccc}{\boldsymbol{I}}_{N-1}&&&\\ &0&1&\\ &1&0&\\ &&&{\boldsymbol{I}}_{n-2}\end{array}\right]\left[\begin{array}[]{cc}{\boldsymbol{I}}_{N+1}&\\ &\left(\phi_{i+2-j}\left(t_{\infty}-t_{3N}\right)\right)_{i\in\left[n-2\right],j\in\left[n\right]}\end{array}\right]$ (63)
It is challenging to reason Corollary 3 by existing costate-based necessary conditions [14] due to the complex behavior of the costates. In contrast, existing costate-based necessary conditions can try to reason the costate for a given trajectory case-by-case.
For example, in a 4th order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$. $\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can serve as a candidate for the optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$. Of course, both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ cannot be optimal in a 3rd order problem. Though Corollary 3 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing the optimal control of problem (53a).
VI-C The Chattering Phenomenon induced by $\left|x_{2}\right|\leq x_{{\mathrm{m}}2}$
Our previous work [18] points out that if the chattering phenomenon occurs in problem (53a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon. An infinite number of unconstrained arcs are joined by the constrained boundary, i.e., the unconstrained arcs are tangent to the boundary for infinite times. In contrast, constrained arcs do not occur. This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition. [18] provides the following conclusion.
Lemma 4.
[18] Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (1a) in a left neighborhood of $t_{\infty}$, where $t_{\infty}$ is the chattering limit point. Then, $\exists\left\{t_{3k}\right\}_{k=0}^{\infty}\subset\left[t_{0},t_{\infty}\right]$ increasing strictly monotonically and converging to $t_{\infty}$, s.t. $x_{2}$ is tangent to $x_{{\mathrm{m}}2}$ at $t_{3k}$. $\forall k\in{\mathbb{N}}^{*}$, $u$ switches at most 2 times during $\left(t_{3k-3},t_{3k}\right)$.
Theorem 3 and Lemma 4 imply the following corollary.
Corollary 4.
Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (1a) in a left neighborhood of $t_{\infty}$. Let $\tau_{k}=t_{\infty}-t_{3k}$, $\forall k\in{\mathbb{N}}$. Denote $f_{m}\left(a,b,c\right)$ as
$\left(b+3a\right)^{m}-3\left(3b+a\right)^{m}+3\left(c+3b\right)^{m}-\left(3c+b\right)^{m}.$ (64)
Then, $\forall N\in{\mathbb{N}}^{*}$, ${\boldsymbol{J}}^{\prime}$ does not have full row rank, where
${\boldsymbol{J}}^{\prime}=\left(f_{i+1}\left(\tau_{j-1},\tau_{j},\tau_{j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[N\right]}.$ (65)
Furthermore, $\forall k\geq n-2$, $\tau_{k}$ can be calculated by the following recursive equation:
$\det\left(f_{i+1}\left(\tau_{k-j-1},\tau_{k-j},\tau_{k-j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[n-2\right]}=0.$ (66)
Proof.
Since $\forall k\in{\mathbb{N}}^{*}$, ${\boldsymbol{x}}_{1:2}\left(t_{3k-3}\right)={\boldsymbol{x}}_{1:2}\left(t_{3k}\right)=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, it can be proved that the control $u$ in $\left(t_{3k-3},t_{3k}\right)$ is
$\begin{dcases}u\left(t\right)\equiv u_{3k-2}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-3},t_{3k-2}\right),\\ u\left(t\right)\equiv u_{3k-1}={u_{\mathrm{m}}},&t\in\left(t_{3k-2},t_{3k-1}\right),\\ u\left(t\right)\equiv u_{3k}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-1},t_{3k}\right).\end{dcases}$ (67)
Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$. Hence, $\Delta u_{3k-1}=2{u_{\mathrm{m}}}$, $\Delta u_{3k}=-2{u_{\mathrm{m}}}$, and $\Delta u_{3k+1}=0$.
$\forall N\in{\mathbb{N}}^{*}$, the augmented switching law between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$. According to Corollary 1, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (62). $\forall k\in\left[N-1\right]$, the $3k$-th columns and the $\left(2k-1\right)$-th rows can be eliminated, and the coefficients of each column, i.e., ±$\pm 2{u_{\mathrm{m}}}$, can be normalized, resulting in ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}_{1}$. Upon left multiplication by (63) and right multiplication by $\left(\delta_{j=i}-\delta_{j=i+1}\right)_{i\in\left[N\right],j\in\left[N\right]}^{3}$, it has ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$. Considering the rank of columns $\left(k-n+3\right)$ through $k$ of ${\boldsymbol{J}}^{\prime}$, (66) holds. ∎
Remark.
According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ when applying (66). When given initial values, (66) is a polynomial equation w.r.t. $\tau_{k}$. It is evident that $n=3$ does not hold since it has $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$. Through more refined derivation, it can be rigorously proven that $n=4$ does not hold. The proof is lengthy and omitted due to space limitation, whose approach is similar to our previous work [18]. In other words, the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ can only occur in a 5th order problem or higher. One can try to determine whether the chattering phenomenon can occur in an $n$-th order problem for $n\geq 5$ based on the proposed Corollary 3.
Remark.
It is challenging to obtain the recursive equation (66) for problem (53a) of high order based on the traditional costate-based necessary condition [14] since the behaviors of costates are significantly complex for high order problems.
Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 4 provides an example to apply the proposed necessary condition to the case the chattering phenomenon occurs, where a sub-arc without chattering of the optimal trajectory is investigated.
VI-D Simulation Experiment
Figure 2: Results of the simulation experiment. (a) The plot of the local minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$. (b) The original trajectory and the optimized trajectory. Among them, ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, ${u_{\mathrm{m}}}=1$.
A simulation experiment is conducted to verify the proposed state-centric necessary condition. As shown in Fig. 2(b), our previous work [8] provided a 4th order suboptimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, $u=u\left(t\right)$, $t\in\left[t,t_{\mathrm{f}}\right]$. The augmented switching law of the original trajectory is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$. According to Corollary 3, $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$ implies that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is not optimal. Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$. Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns is full row rank. Then, disturb $t_{4}$ into $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (53a). As shown in Fig. 2(a), $t_{\mathrm{f}}^{\prime}$ achieves local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$ shown in Fig. 2(b). It can be observed in Fig. 2(b4) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory. In other words, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint; hence, ${\boldsymbol{t}}^{\prime}$ cannot be further optimized under fixing ${\mathcal{L}}$. Among them,
$\begin{dcases}{\boldsymbol{t}}=\left(1.7500,2.1562,3.1562,4.5342,5.5342,\right.\\ \quad\left.6.0342,6.5221,7.0100,7.2226,8.9165,9.8604\right),\\ {\boldsymbol{t}}^{\prime}=\left(1.7500,2.1562,3.1562,4.5550,5.5550,\right.\\ \quad\left.6.2023,6.6621,7.1218,7.2166,8.9104,9.8543\right).\end{dcases}$ (68)
Therefore, ${\boldsymbol{x}}^{\prime}$ achieves a $0.6\%$ reduction in the terminal time compared to ${\boldsymbol{x}}$. The proposed state-centric necessary condition is verified through the simulation experiment. More efficient optimization algorithms based on the proposed state-centric necessary condition can be developed in future works.
VII Conclusions
This paper has set out to establish an innovative theoretical framework for the augmented switching law for time-optimal control of controllable linear systems and its applications. The proposed augmented switching law represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form. Specifically, the equality and inequality constraints induced by the augmented switching law represent a sufficient and necessary condition for the feasibility of the Bang-Singular-Bang trajectory. Based on the augmented switching law, a given feasible trajectory can be disturbed with the fixing augmented switching law and feasibility guarantees, resulting in a state-centric necessary condition for optimal control since any disturbed feasible trajectory should have a strictly longer terminal time than the optimal terminal time due to the uniqueness of optimal control. The proposed necessary condition states that the Jacobian matrix induced by the augmented switching law must not be full row rank. High order conditions can be derived based on the local optimality of the optimal trajectory. Different from the traditional necessary condition based on adjoining methods, the developed necessary condition requires only state and control information without dependence on costate information.
The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrators systems with full box constraints. Under the condition that every arc is strictly feasible except for the constraints active during the whole arc, the proposed state-centric necessary condition implies that the number of arcs should be no greater than the sum of the order of constrained arcs plus the order of the problem. Considering the chattering phenomenon induced by the 2nd order state constraints, this paper proves a recursive equation for the junction time before the chattering limit point. The above two conclusions are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed state-centric necessary conditions.
Acknowledgment
This work was supported by the National Key Research and Development Program of China under Grant 2023YFB4302003.

### v2

@@
-A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws
+A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)
 Abstract
-Most existing necessary conditions for optimal control based on adjoining methods require both state information and costate information, yet the lack of costates for a given feasible trajectory in practice impedes the determination of optimality.
-This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law that represents the input control and the feasibility in a compact form.
-Given a feasible trajectory, the disturbed trajectory under the constraints of augmented switching law is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information.
-A first order necessary condition is proposed that the Jacobian matrix of the augmented switching law is not full row rank, which also results in an approach to optimizing a given feasible trajectory further.
-The proposed necessary condition is applied to the chain-of-integrators systems with full box constraints, contributing to some conclusions challenging to reason by traditional costate-based necessary conditions.
+Most existing necessary conditions for optimal control based on adjoining methods require both state and costate information, yet the unobservability of costates for a given feasible trajectory impedes the determination of optimality in practice.
+This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems with a single input, proposing the augmented switching law (ASL) that represents the input control and the feasibility in a compact form.
+Given a feasible trajectory, the perturbed trajectory under the constraints of ASL is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information.
+A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures.
+The proposed necessary condition is applied to high-order chain-of-integrator systems with full box constraints, contributing to some theoretical results challenging to reason by costate-based conditions.
 Index Terms:
@@
 Numerous works have been conducted on the behaviors of the optimal controls, resulting in some well-known necessary or sufficient conditions [7].
-However, while referring to existing conclusions of optimal control for problems in general forms, it remains difficult to fully solve a high order problem in practice, especially when state inequality constraints are introduced into the system [8].
-Specifically, when given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on state information since most existing necessary conditions require costate information.
+However, it remains difficult to fully solve a high-order problem in practice, especially when state inequality constraints are introduced [8, 9].
+Specifically, when given a feasible trajectory, it is challenging to determine the trajectory’s optimality based on state information since most existing necessary conditions require costate information, especially when the trajectory is not planned by costate-based methods.
 In the domain of optimal control theory, numerous necessary conditions are developed based on the variational method [10] and Pontryagin’s maximum principle (PMP) [11].
@@
 Jacobson et al.
-[14] pointed out that constrained arcs are not allowed when the state constraint is of odd order $p>1$, and instead, the unconstrained arc is tangent to the constrained boundary at a single point.
+[9] pointed out that constrained arcs are not allowed when the state constraint is of odd order $p>1$, and instead, the unconstrained arc is tangent to the constraint’s boundary at a single point.
 Makowski and Neustadt [12] generalized the PMP-based necessary condition to the mixed constraints of state and control.
@@
 [15] developed the necessary condition based on the indirect adjoining method, further supplemented by Kreindler [16].
-Hartl et al.
-[7] systematically reviewed existing PMP-based necessary conditions and summarized their general forms.
 The necessary conditions are usually applied to rule out some forms of non-optimal controls in practice.
-In some simple cases, the optimal control can be fully solved based on some necessary conditions [18, 19], while in other cases, the necessary conditions are applied to guide the design of suboptimal trajectories [8].
-However, the above necessary conditions require information on the states and the costates simultaneously, while the costates are not directly observable in practice.
-Given a planned feasible trajectory, it is challenging to determine whether the trajectory is optimal based on the pure information on states, let alone the further optimization based on the given trajectory.
-In this paper, a necessary condition is called to be state-centric if it only requires information on states and controls, without dependence on costate information.
-There have existed some state-centric necessary so far, most of which are based on Bellman’s principle of optimality [20].
-Bellman’s principle of optimality states that a sub-arc of an optimal trajectory is also optimal for the induced sub-problem, where the costates are not introduced.
-The Hamilton-Jacobi-Bellman (HJB) equation [21] represents a state-centric necessary in the form of the partial differential equation, which is widely used in optimal control [22] including reinforcement learning [23].
+However, the above necessary conditions require information on both states and costates, while costates are usually unobservable in practice, resulting in a challenge in determining the optimality of a given feasible trajectory by costate-based conditions.
+Based on PMP, one can get rid of dependence on costates by deriving the optimal trajectory’s structure.
+Switching surfaces can be constructed in the state space for time-optimal control of linear systems.
+Walther et al.
+[17] computated switching surfaces for some systems through Gröebner basis.
+Patil et al.
+[18] represented switching surfaces using state equations for single-input diagonal linear systems without state constraints.
+However, for systems with state constraints, it remains challenging to completely construct switching surfaces due to the lack of optimal switching laws.
+For example, He et al.
+[19] provided explicit equations for switching surfaces of 3rd-order chain-of-integrator systems with full state constraints, limited to zero terminal states.
+Wang et al.
+[20] extended [19] to non-zero terminal conditions, failing to achieve time-optimality for higher-order systems.
+Numerical methods [21] face the curse of dimensionality.
+Hence, there is a lack of efficient and general optimality conditions without dependence on costates for high-order systems with state constraints.
+Another type of necessary condition leverages Bellman’s principle of optimality [22], asserting that a sub-arc of an optimal trajectory is itself optimal for the induced sub-problem, without explicitly involving costates.
+The Hamilton-Jacobi-Bellman (HJB) equation [23] represents a typical necessary condition in the form of partial differential equations, which is widely used in optimal control [24, 25].
 Evans and James [26] investigated the HJB equation for time-optimal control.
 Wolenski and Zhuang [27] showed that the minimal time function is the unique proximal solution to the HJB equation.
-However, most numerical methods based on the HJB equation require high computational cost [26], where the continuous time domain is discretized into a grid.
-A trade-off exists between the computational cost and the computational accuracy in practice.
-It remains challenging to develop a state-centric necessary condition for time-optimal control in a compact form with limited computational cost.
-This paper sets out to establish a theoretical framework for time-optimal control of controllable linear systems and develop a novel state-centric necessary condition.
-The optimal control problem is formulated and investigated based on Hamiltonian in Section II.
-Section III analyzes the structure of arcs and the switching law of the optimal trajectory, where the connections between adjacent arcs are fully discussed.
-The keypoints for feasibility are introduced in Section IV, where the augmented switching law is proposed to represent the control and the feasibility in a compact form.
-Based on the uniqueness of the optimal control and the feasibility of the disturbed trajectory, Section V develops a novel state-centric necessary condition for time-optimal control.
-The proposed necessary condition is further applied to the chain-of-integrators systems with full box constraints in Section VI as examples.
+However, most numerical HJB-based methods require high computational cost [28] due to the discretized grids.
+From a theoretical perspective, HJB-based methods also fail to preserve arc structures, which hinders the analytical reasoning.
+Compared with the PMP-based and HJB-based necessary conditions, this paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information.
+Sections III and IV derive trajectories’ structures on arcs and keypoints for feasibility, respectively.
+Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively.
 The contributions of this paper are as follows:
   1. 1.
-    This paper establishes an innovative theoretical framework for time-optimal control of controllable linear systems, proposing the augmented switching law which fully represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form.
-The established framework focuses on the control of each unconstrained arc and constrained arc, the constraints induced by the connections of adjoining arcs, the feasibility of the disturbed trajectory near the constrained boundary, and the optimality of a given feasible trajectory.
-In the proposed theory, the control can be fully represented by the augmented switching law and the motion time of each arc, where the constraints directly induced by the augmented switching law can guarantee the feasibility of the trajectory near the constrained boundary in a sufficient and necessary sense.
-The developed framework provides a novel variational approach as well as a local optimization method with feasibility assurance, i.e., disturbing the motion time of each arc under fixed augmented switching law, resulting in disturbed trajectories with feasibility guarantees.
+    This paper establishes a novel theoretical framework for time-optimal control of controllable single-input linear systems, proposing the augmented switching law (ASL) which represents structures and feasibility of bang-bang and singular (BBS) trajectories in a compact form.
+Compared to existing representations of switching surfaces, the ASL can generally deal with high-order systems with state constraints.
+The developed framework provides a novel variational approach with feasibility assurance, i.e., perturbing the motion time of each arc with a fixed ASL and resulting in feasible perturbed BBS trajectories.
   2. 2.
-    This paper proposes a novel state-centric necessary condition for optimal control of controllable linear systems, which only requires the state and control information without dependence on costate information.
-Given a Bang-Singular-Bang feasible trajectory which is claimed to be optimal, all disturbed feasible trajectories should result in a terminal time which is strictly longer than that of the original trajectory.
-In this way, a first order necessary condition is proposed that the Jacobian matrix of the augmented switching law should not be full row rank.
-The higher order necessary conditions are also expected to be defined based on the strict local optimality of the trajectory.
-Based on the proposed necessary condition, a novel optimization approach with a given Bang-Singular-Bang feasible trajectory as initial value is proposed, where the optimized trajectory always has the same augmented switching law as the initial trajectory in a Bang-Singular-Bang form and the feasibility is always guaranteed during iteration.
+    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not full row rank.
+The condition is derived from the fact that if a BBS trajectory is optimal, then all perturbed feasible trajectories have a strictly longer terminal time than the original one.
+Compared to existing costate-based conditions, the proposed state-centric necessary condition only requires state information without dependence on costates.
+Note that costates are usually unobservable in practice, especially when the trajectory is not planned by costate-based optimization.
   3. 3.
-    The proposed state-centric necessary condition is applied to high-order chain-of-integrators systems with full box constraints as examples, contributing to some conclusions that are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed necessary condition.
-Note that time-optimal control for chain-of-integrators systems with full box constraints represents a challenging and open problem in optimal control theory and kinematics, yet to be resolved.
-The number of arcs is proved to be less than the order of the system plus the sum of the order of arcs’ constraints under some conditions, while the traditional methods have to analyze the costates case-by-case.
-The proposed state-centric necessary condition successfully proves the recursive equations of motion time if the chattering phenomenon occurs and is induced by the 2nd order state constraints, which is challenging based on costate analysis due to the high degree-of-freedom of costates in high order problems.
-As a result, it is hopeful to determine the existence of chattering induced by the 2nd order state constraints in chain-of-integrators systems in future works.
-II Problem Formulation and Assumptions
-This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is defined as follows:
+    This paper proposes a potential optimization idea based on the developed necessary condition.
+Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not full row rank.
+During iteration, the feasibility is guaranteed based on the developed ASL framework.
+Compared to discretized methods, the proposed approach can preserve the BBS arc structure and has the potential to significantly reduce control oscillations and open-loop integration errors of terminal states.
+  4. 4.
+    The proposed state-centric necessary condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain.
+An upper bound on the number of arcs is determined in a general sense, while case-by-case analysis is necessary in traditional methods.
+A recursive equation of junction time is reasoned in chattering induced by 2nd-order constraints, which is challenging to prove by costate analysis.
+As a result, one could determine the existence of chattering induced by 2nd-order constraints in chain-of-integrator systems in future works.
+The above results demonstrate the theoretical effectiveness of the proposed necessary condition.
+II Problem Formulation
+Some notations should be introduced.
+For an integer $n\in{\mathbb{N}}$, the set $\left[n\right]$ refers to $\left\{1,2,\dots,n\right\}$.
+$\left(x_{i}\right)_{i\in\mathcal{I}}\in{\mathbb{R}}^{\left|\mathcal{I}\right|}$ is a column vector and $\left(x_{ij}\right)_{i\in\mathcal{I},j\in\mathcal{J}}\in{\mathbb{R}}^{\left|\mathcal{I}\right|\times\left|\mathcal{J}\right|}$ is a matrix, where $\mathcal{I}$ and $\mathcal{J}$ are index sets.
+Specifically, $\left(x_{i}\right)_{i=1}^{n}\in{\mathbb{R}}^{n}$ refers to $\left(x_{i}\right)_{i\in\left[n\right]}$.
+This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is as follows:
 $\min_{u}\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t={t_{\mathrm{f}}},$ (1a)
@@
 ${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{t_{\mathrm{f}}},$ (1d)
-$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (1e)
-[…]
+$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (1e)
+where ${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$, ${\boldsymbol{C}}=\left({\boldsymbol{c}}_{p}\right)_{p=1}^{P}\in{\mathbb{R}}^{P\times n}$, and ${\boldsymbol{d}}=\left(d_{p}\right)_{p=1}^{P}\in{\mathbb{R}}^{P}$, where $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$.
+In problem (1a), the terminal time ${t_{\mathrm{f}}}$ is free.
+${\boldsymbol{x}}=\left(x_{k}\right)_{k=1}^{n}\in{\mathbb{R}}^{n}$ is the state vector.
+The state constraint (1c) requires that $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$.
+The input control $u$ is subject to a box constraint (1e).
 Assume that the system (1b) is controllable, i.e.,
@@
 Remark.
-Generally, an equality constraint ${\boldsymbol{F}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ can be introduced into problem (1a).
+Generally, an equality constraint ${\boldsymbol{F}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ can be incorporated into problem (1a).
 However, it is not considered in this paper since the equality constraint can be equivalently eliminated by a linear transformation.
@@
 ${\boldsymbol{\lambda}}=\left(\lambda_{k}\right)_{k=1}^{n}$ is the costate vector.
-$\lambda_{0}$ and ${\boldsymbol{\lambda}}$ satisfy $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\not=0$.
+$\forall t\in\left[0,{t_{\mathrm{f}}}\right]$, $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\in{\mathbb{R}}^{n+1}$ is not ${\boldsymbol{0}}$ [8].
 The Euler-Lagrange equations [8] holds that
@@
 In (3), ${\boldsymbol{\eta}}=\left(\eta_{p}\right)_{p=1}^{P}$ is the multiplier vector induced by the state inequality constraint (1c), satisfying
-$\eta_{p}\geq 0,\,\eta_{k}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)=0,\,\forall 1\leq p\leq P.$ (5)
+$\eta_{p}\geq 0,\,\eta_{p}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)=0,\,\forall p\in\left[P\right].$ (5)
 Therefore, $\forall t\in\left[0,t_{\mathrm{f}}\right]$, $\eta_{p}\left(t\right)>0$ only if ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}=0$.
@@
 The notation $\equiv$” means that the equality holds for a continuous period.
-(7) is the well-known Bang-Singular-Bang law [8].
+(7) is the well-known bang-bang and singular (BBS) control law [11, 29].
+In other words, the optimal control is always ±$\pm{u_{\mathrm{m}}}$, except during singular arcs.
 Note that the objective function $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
 $\forall t\in\left[0,t_{\mathrm{f}}\right],\,{\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)\equiv 0.$ (8)
-For convenience of discussion, the following assumptions are introduced:
 Assumption 1.
@@
 In the following of this paper, Assumption 1 is considered to hold throughout unless otherwise specified.
+Given a feasible trajectory $\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right)\right)$, one can determine the optimality through PMP-based necessary conditions like (4)–(8).
+However, in practice, only $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$ is observable, while $\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right)\right)$ is difficult to fully solve based on provided $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$.
+Consequently, the focus of this paper is to determine the optimality of a feasible trajectory in problem (1a) using the provided $\left({\boldsymbol{x}}\left(t\right),u\left(t\right)\right)$.
 III Arc Representation of the Optimal Trajectory
-An arc of a trajectory is defined as a continuous segment of the optimal trajectory conditioned by the same input control.
-Two kinds of arcs exist, i.e., unconstrained arcs and constrained arcs, which are discussed in Section III-A and Section III-B, respectively.
+An arc of a trajectory is defined as a continuous segment conditioned by a fixed dynamic equation, i.e., $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ where $\widehat{{\boldsymbol{A}}}$ and $\widehat{{\boldsymbol{b}}}$ are constant.
+Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A and Section III-B, respectively.
 Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C.
-Finally, if the chattering phenomenon does not occur, an optimal trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
+Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
 III-A Unconstrained Arcs in Problem (1a)
-In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds strictly almost everywhere, and the control does not switch.
+In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds almost everywhere (a.e.), and the control does not switch.
 The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
 Proposition 1.
-In an unconstrained arc, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ almost everywhere, and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
+In an unconstrained arc, it holds a.e.
+that ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
 Proof.
 Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period.
-By (5), ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ in an unconstrained arc.
+By (5), ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ implies that ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ holds a.e.
+in an unconstrained arc.
 Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$.
-Considering the $k$-th order derivative of ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, it has
-$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}\equiv 0.$ (9)
-By (2), ${\boldsymbol{\lambda}}\equiv 0$.
+Therefore,
+$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}=\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\left({\boldsymbol{\lambda}}^{\top}{\boldsymbol{b}}\right)\equiv 0,$ (9)
+i.e., ${\boldsymbol{\lambda}}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\equiv{\boldsymbol{0}}$.
+However, by (2), ${\boldsymbol{\lambda}}\equiv 0$ holds a.e.
 Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$.
-Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$.∎
+Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$ holds a.e.
+∎
 Remark.
@@
 Compared to unconstrained arcs, constrained arcs exhibit more complex behavior.
-In a constrained arc, $\exists 1\leq p\leq P$, s.t.
+In a constrained arc, $\exists p\in\left[P\right]$, s.t.
 ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$.
-The following proposition characterizes an unconstrained arc.
+The following proposition characterizes a constrained arc.
 Proposition 2.
-If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in a constrained arc, then $\exists 1\leq r_{p}\leq n$, s.t.
-$\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$.
-During the constrained arc,
-$u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}},$ (10)
-and
-$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\equiv 0.\end{dcases}$ (11)
-Conversely, if the initial value ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies
-$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{0}\right)=0,\end{dcases}$ (12)
-then under the driving of (10), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds.
+If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in a constrained arc for $t\in\left[t_{1},t_{2}\right]$, then $\exists r_{p}\in\left[n\right]$, s.t.
+$\forall r\in\left[r_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$.
+During the constrained arc, it holds that
+$u=\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}},\,\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}},\,\widehat{{\boldsymbol{a}}}_{p}\triangleq-\frac{\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\right)^{\top}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},$ (10)
+and $\forall t\in\left[t_{1},t_{2}\right]$,
+$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0,\\ \forall r\in\left[r_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}=0.\end{dcases}$ (11)
+$r_{p}$ is called the order of the active constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$.
+Conversely, if $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (11), then under the driving of (10), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in $\left[t_{1},t_{2}\right]$.
 Proof.
-Since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$, (2) implies the existence of the above $r_{p}$.
-Considering the $k$-th order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$.
+(2) implies ${\boldsymbol{c}}_{p}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\not={\boldsymbol{0}}$ since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$; hence, $r_{p}$ exists and serves as the minimum non-zero column index of a non-vzero matrix.
+For $r\in\left[r_{p}-1\right]$, considering the $r$th-order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$.
 Hence, (10) holds.
-Assume that (10) holds and ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (12).
-Note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}=0$; hence, the initial condition implies that (11) holds.
+Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (11).
+Under the driving of (10), note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}=0$; hence, (11) holds in $\left[t_{1},t_{2}\right]$.
 ∎
-In the above proposition, $r_{p}$ is called the order of the equality state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0$.
-Similarly, the analytical expression of a constrained arc can be obtained by Lemma 1.
-Cases exist where multiple inequality constraints are active in a constrained arc.
-For this case, the active constraints should be consistent.
-Specifically, denote that $\mathcal{P}=\left\{p\in\left[P\right]:{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0\right\}$, where $\left[P\right]\triangleq\left\{1,2,\dots,P\right\}$.
-Then, $\forall p_{1},p_{2}\in\mathcal{P}$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}-1}{\boldsymbol{b}}}=\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}-1}{\boldsymbol{b}}}$ holds in the arc due to (10), and $\forall p\in\mathcal{P}$, (11) holds.
-The following example illustrates the above analysis:
+It is allowed that multiple inequality constraints are active in a constrained arc, as proved in Proposition 3.
+Lemma 1.
+${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$ and ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$.
+The initial value problem where $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}$, ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ has a unique solution:
+$\forall t\in{\mathbb{R}},\,{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (12)
+Furthermore, if ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$ satisfies $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$.
+Proof.
+The proof of (12) is omitted since it is elementary [31].
+Assume that $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
+Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t.
+$p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$.
+Therefore, $\forall r\geq n$, $\exists p_{r}^{*}$ s.t.
+$p_{r}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{r}^{*}\left({\boldsymbol{A}}\right)$.
+Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$.
+Since $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, Note that $\forall r\geq n+1$,
+$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$ (13)
+It can be proved that ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)$ is analytic. By (13), $\forall t\in{\mathbb{R}}$,
+${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right).$ (14)
+In other words, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$. ∎
+Proposition 3.
+[…]
+$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0,\\ \forall q\in\mathcal{P},\,r\in\left[n\right],\,{\boldsymbol{c}}_{q}^{\top}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right)^{r}{\boldsymbol{x}}=0.\end{dcases}$ (15)
+Active constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ for $p\in\mathcal{P}$ are called consistent if linear equations (15) have a feasible solution ${\boldsymbol{x}}$.
+Conversely, assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (15).
+Then, under the driving of (10) with an arbitrary $p\in\mathcal{P}$, it holds that $\forall q\in\mathcal{P}$, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}+d_{q}\equiv 0$ in $\left[t_{1},t_{2}\right]$.
+Proof.
+Assume that $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in the constrianed arc.
+By Proposition 2, (10) holds $\forall p\in\mathcal{P}$.
+$\forall q\in\mathcal{P}$, since ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv{\mathrm{const}}$, (15) can be derived through considering the $r$th-order derivative of ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}$ for $r\in\left[n\right]$.
+Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (15).
+$\exists p\in\mathcal{P}$, (15) holds in $\left[t_{1},t_{2}\right]$.
+Then, $\forall q\in\mathcal{P}$, $r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds.
+By Lemma 1, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv-d_{q}$ holds.
+∎
 Example 1.
-Consider an $n$-th order chain-of-integrators system, i.e.,
-$\begin{dcases}\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right],\,\forall 1<k\leq n,\\ \dot{x}_{1}\left(t\right)=u\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right].\end{dcases}$ (13)
+Consider a chain-of-integrator system of order $n$, i.e.,
+[…]
 Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$.
-Assume that both $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ hold in a constrained arc.
-Then, $u\equiv-x_{2}=0$; hence, $x_{3}\equiv 1$, $x_{2}\equiv 0$, and $x_{1}\equiv 0$.
-Therefore, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent.
-Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are not consistent.
+$x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold.
+In this case, $x_{2}\equiv 0$ implies $x_{1}\equiv 0$ and $u\equiv 0$; hence, $x_{1}+x_{3}\leq 1$ implies $x_{3}\equiv 1$.
+Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are inconsistent, i.e., in a constrained arc, if $x_{1}+x_{3}\equiv 1$ holds, then $x_{2}\equiv 1$ cannot hold.
 III-C System Behavior
-Following the terminology introduced in our previous work [8], the system behavior is defined as follows:
 Definition 1.
-A system behavior in problem (1a) is an arc with constant system dynamics, denoted as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$.
-During the arc, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$.
-$\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{g}}}\equiv{\boldsymbol{0}}$ provides fully equality state constraints, where $\widehat{{\boldsymbol{F}}}$ has full row rank.
-$\mathcal{P}$ is the set of active inequality state constraints.
-In Definition 1, an unconstrained arc can be denoted as ${\mathcal{S}}=\left({\boldsymbol{A}},u_{0}{\boldsymbol{b}},\sim,\sim,\varnothing\right)$, where $\sim$ means that no inequality constraints are introduced, and $u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ is a constant.
-For a constrained arc, $\mathcal{P}\not=\varnothing$.
-(10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{p}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}_{p}$, where
-$\widehat{{\boldsymbol{A}}}_{p}={\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\widehat{{\boldsymbol{b}}}_{p}={\boldsymbol{0}}.$ (14)
-Although it allows that $\exists p_{1},p_{2}\in\mathcal{P}$, s.t.
-$\widehat{{\boldsymbol{A}}}_{p_{1}}\not=\widehat{{\boldsymbol{A}}}_{p_{2}}$, it is guaranteed by the consistency of active constraints that $\widehat{{\boldsymbol{A}}}_{p_{1}}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{A}}}_{p_{2}}{\boldsymbol{x}}$ during the arc.
-For convenience of discussion, let $\widehat{{\boldsymbol{A}}}$ be $\widehat{{\boldsymbol{A}}}_{p}$ with the minimum $p\in\mathcal{P}$, i.e., $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$.
-Evidently, $\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}}\right)$ is well-defined in Definition 1.
-The active inequality state constraints state that $\forall p\in\mathcal{P}$, (11) holds.
-The above equalities induce $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ where $\widehat{{\boldsymbol{F}}}$ has full row rank.
-For well-posedness, $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ are determined by the equality constraints with minimum indices.
+A system behavior in problem (1a) is an arc with constant system dynamics, denoted as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$.
+During the arc, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ holds.
+Induced by Proposition 3, ${\boldsymbol{F}}{\boldsymbol{x}}+{\boldsymbol{g}}\equiv{\boldsymbol{0}}$ provides full equality constraints, where ${\boldsymbol{F}}$ has full row rank.
+$\mathcal{P}$ is the set of consistent active state constraints.
+In Definition 1, an unconstrained arc can be denoted as ${\mathcal{S}}=\left({\boldsymbol{A}},u_{0}{\boldsymbol{b}},\sim,\sim,\varnothing\right)$, where $\sim$ means that no equality constraints are introduced.
+$u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ is a constant.
+By Lemma 1,
+${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right)+u_{0}\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (17)
+For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not=\varnothing$.
+(10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}$, where $\widehat{{\boldsymbol{a}}}_{p}$ is defined in (10).
+Although it allows that $\exists p,q\in\mathcal{P}$, s.t.
+$\widehat{{\boldsymbol{a}}}_{p}\not=\widehat{{\boldsymbol{a}}}_{q}$, it is guaranteed by the consistency of $\mathcal{P}$ that $\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{a}}}_{q}^{\top}{\boldsymbol{x}}$ during the arc.
+For convenience, let $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$.
+Then, ${\mathcal{S}}$ in Definition 1 is well-defined.
+By Lemma 1, during ${\mathcal{S}}$,
+${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{\widehat{{\boldsymbol{A}}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right).$ (18)
 Example 2.
-Consider a chain-of-integrators system (13) with full box state constraints, i.e., $\forall 1\leq k\leq n$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
+Consider a chain-of-integrator system (16) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
 The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
-A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$-th order, where $\forall 1\leq j<k$, $x_{j}\equiv 0$, and $u\equiv 0$.
-Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$.
-Among them, $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i,j=1}^{n}$, where $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$, and otherwise, $\hat{a}_{ij}=0$.
-In a system behavior ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ holds.
-Therefore, the analytical expression of ${\boldsymbol{x}}\left(t\right)$ in ${\mathcal{S}}$ can be given by the following lemma.
-Lemma 1.
-The initial value problem (IVP) where $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}\left(t\right)$, $t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ has a unique solution:
-$\forall t\in{\mathbb{R}},\,{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau.$ (15)
-Proof.
-$\frac{\mathrm{d}}{\mathrm{d}t}\left({\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}\left(\dot{\boldsymbol{x}}\left(t\right)-{\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{b}}\left(t\right)$ implies that ${\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)-{\mathrm{e}}^{-{\boldsymbol{A}}t_{0}}{\boldsymbol{x}}_{0}=\int_{t_{0}}^{t}{\mathrm{e}}^{-{\boldsymbol{A}}\tau}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau$.
-∎
-Lemma 1 implies that for an unconstrained arc where $u\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$,
-${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right)+u_{0}\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (16)
-For a constrained arc,
-${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{\widehat{{\boldsymbol{A}}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right).$ (17)
-For a given system behavior, ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ has an explicit elementary expression since every scalar element of ${\boldsymbol{e}}^{\widehat{{\boldsymbol{A}}}t}$ is composed of exponentials and polynomials w.r.t.
-$t$.
+A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$th-order, where $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$, and $u\equiv 0$.
+Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$ where $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i\in\left[n\right],j\in\left[n\right]}$.
+$\hat{a}_{ij}=1$ if $k<i=j+1\leq n$; otherwise, $\hat{a}_{ij}=0$.
+Figure 1: A feasible BBS trajectory under the constraint ${\boldsymbol{x}}\geq{\boldsymbol{0}}$.
+The trajectory consists of 3 arcs, i.e., ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{2}$, and ${\mathcal{S}}_{3}$.
+${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ is connected at ${\boldsymbol{0}}$, denoted as $\mathcal{E}_{1}$.
+The trajectory is tangent to $\left\{x_{3}=0\right\}$ at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, and $\mathcal{T}_{2}^{(2)}$.
 III-D Switching Law
@@
 Definition 2.
-Assume that the optimal trajectory in problem (1a) consists of a finite number of arcs represented as ${\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{N}$ successively.
-The switching law of the optimal trajectory is defined as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
-Remark.
+Assume that a BBS trajectory in problem (1a) consists of a finite number of arcs ${\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{N}$ successively.
+The switching law is denoted as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
 The switching law focuses on the arcs that compose the optimal trajectory, while the motion time of each arc is not included in a switching law.
-Furthermore, Definition 2 does not characterize the chattering phenomenon since chattering induces an infinite number of arcs.
-The connection between arcs is analyzed in Proposition 3.
-Denote the two adjacent arcs as ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$, where ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$.
-$\forall i=1,2$, ${\mathcal{S}}_{i}=\left(\widehat{{\boldsymbol{A}}}_{i},\widehat{{\boldsymbol{b}}}_{i},\widehat{{\boldsymbol{F}}}_{i},\widehat{{\boldsymbol{g}}}_{i},\mathcal{P}_{i}\right)$.
-Among them, ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ is defined as $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}$, $\widehat{{\boldsymbol{b}}}_{1}=\widehat{{\boldsymbol{b}}}_{2}$, and $\mathcal{P}_{1}=\mathcal{P}_{2}$ since $\widehat{{\boldsymbol{F}}}_{i}$ and $\widehat{{\boldsymbol{g}}}_{i}$ can be induced by $\mathcal{P}_{i}$.
-Before introducing Proposition 3, a lemma is provided as preparation.
-Lemma 2.
-Consider the IVP $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}$, $t\in{\mathbb{R}}$, where ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, ${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$.
-If ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$, s.t.
-$\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$.
-Proof.
-According to Lemma 1, $\forall t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\left(\tau\right)\mathrm{d}\tau\right){\boldsymbol{b}}$.
-Considering the Jordan normal form of ${\boldsymbol{A}}$, it has $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{i=1}^{l}p_{i}\left(t-t_{0}\right){\boldsymbol{e}}^{\lambda_{i}\left(t-t_{0}\right)}$, where $\left\{\lambda_{i}\right\}_{i=1}^{l}$ are the eigenvalues of ${\boldsymbol{A}}$, and $p_{i}$ is a polynomial of degree not greater than $n$.
-Evidently, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}$ is an analytic function; hence, $\forall t\in{\mathbb{R}}$, it has
-${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}.$ (18)
-The case where ${\boldsymbol{A}}={\boldsymbol{0}}$ is trivial.
-Assume that ${\boldsymbol{A}}\not={\boldsymbol{0}}$.
-Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t.
-$p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$.
-Therefore, $\forall r\geq n$, $\exists p_{n}^{*}$ s.t.
-$p_{n}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{n}^{*}\left({\boldsymbol{A}}\right)$.
-Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$.
-Since $\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, it holds that $\forall r\geq n+1$,
-$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)$ (19)
-$=$ ${\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$
-By (18), $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right)$. ∎
+For the example in Fig.
+1, the switching law is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}$.
+${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{3}$ are unconstrained arcs where ${\boldsymbol{x}}>{\boldsymbol{0}}$ holds a.e.
+${\mathcal{S}}_{2}$ is a constrained arc where $x_{2}\equiv 0$.
+The following proposition discusses connections of two adjacent arcs, including unconstrained and constrained arcs.
 Proposition 4.
-Assume that $\forall i=1,2$, ${\mathcal{S}}_{i}$ occurs for $t\in\left[t_{i-1},t_{i}\right]$ in the above notation. ${\boldsymbol{x}}$ enters ${\mathcal{S}}_{2}$ from ${\mathcal{S}}_{1}$ at $t_{1}$.
+Assume that ${\mathcal{S}}_{i}=\left(\widehat{{\boldsymbol{A}}}_{i},\widehat{{\boldsymbol{b}}}_{i},{\boldsymbol{F}}_{i},{\boldsymbol{g}}_{i},\mathcal{P}_{i}\right)$ occurs in $\left[t_{i-1},t_{i}\right]$, $i=1,2$, where $\left[\widehat{{\boldsymbol{A}}}_{1},\widehat{{\boldsymbol{b}}}_{1}\right]\not=\left[\widehat{{\boldsymbol{A}}}_{2},\widehat{{\boldsymbol{b}}}_{2}\right]$.
+The following conclusions hold at the connection of ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
   1. 1.
@@
     If ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ are constrained arcs, then $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$.
-Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
-    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (20)
-    $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
-    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (21)
+Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
+    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (19)
+    $\forall p\in\mathcal{P}_{2}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
+    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (20)
     Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ for $\left(t_{1},t_{2}\right)$.
@@
     If ${\mathcal{S}}_{1}$ is a constrained arc and ${\mathcal{S}}_{2}$ is an unconstrained arc, then $\forall p,q\in\mathcal{P}_{1}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$.
-On ${\mathcal{S}}_{2}$, $\forall p\in\mathcal{P}_{1}$, it has
+On ${\mathcal{S}}_{2}$, $\forall p\in\mathcal{P}_{1}$, it holds that
     $u\equiv u_{2}=-{\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right){u_{\mathrm{m}}}.$ (21)
-    Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
-    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)=0.\end{dcases}$ (23)
+    Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
+    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)=0.\end{dcases}$ (22)
     Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$ for $\left(t_{1},t_{2}\right)$ where $u_{2}$ is given in (21).
@@
     If ${\mathcal{S}}_{1}$ is an unconstrained arc and ${\mathcal{S}}_{2}$ is a constrained arc, then $\forall p,q\in\mathcal{P}_{2}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$.
-On ${\mathcal{S}}_{1}$, $\forall p\in\mathcal{P}_{2}$, it has
+On ${\mathcal{S}}_{1}$, $\forall p\in\mathcal{P}_{2}$, it holds that
     $u\equiv u_{1}=\left(-1\right)^{r_{p}-1}{\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right){u_{\mathrm{m}}}.$ (23)
-    Furthermore, $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
-    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)=0.\end{dcases}$ (25)
+    Furthermore, $\forall p\in\mathcal{P}_{2}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
+    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)=0.\end{dcases}$ (24)
     Conversely, assume that $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{1}{\boldsymbol{b}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{1},t_{2}\right)$ where $u_{1}$ is given in (23).
@@
 Since $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}={\boldsymbol{A}}$ and $\mathcal{P}_{1}=\mathcal{P}_{2}=\varnothing$, it holds that $\widehat{{\boldsymbol{b}}}_{1}\not=\widehat{{\boldsymbol{b}}}_{2}$; hence, $u_{1}=-u_{2}$.
-Therefore, Proposition 3-1 holds.
+Therefore, Proposition 4.1 holds.
 Consider the connection between two constrained arcs.
@@
 Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration.
-According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}}}{{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}$ holds.
-Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$.
+According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p^{*}}^{\top}\right){\boldsymbol{x}}$ holds.
+Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall r\in\left[r_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$.
 By Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds.
@@
 Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ and $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)$.
-Lemma 2 implies that $\exists 1\leq\hat{r}_{p}\leq n$, s.t.
-$\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
-Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, it has $\hat{r}_{p}\geq r_{p}$.
+Lemma 1 implies that $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
+$\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall r\in\left[\hat{r}_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
+Since $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, it holds that $\hat{r}_{p}\geq r_{p}$.
 In other words, (19) holds.
@@
 Then, $\forall p\in\mathcal{P}_{1}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0$ and
-$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (26)
+$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (25)
 Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$.
 For the same reason, $\exists\varepsilon^{\prime}\in\left(0,\varepsilon\right)$, $\forall p\in\mathcal{P}_{2}$, $t\in\left(t_{1}-\varepsilon^{\prime},t_{1}\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$.
-Hence, Proposition 3-2 holds.
+Hence, Proposition 4.2 holds.
 Consider the connection between an unconstrained arc and a constrained arc.
@@
 By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{-}}=0$; hence,
-$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0,\\ {\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u\left(t_{1}^{-}\right){\boldsymbol{b}}\right)=0.\end{dcases}$ (27)
+$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0,\\ \forall r\in\left[r_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0,\\ {\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u\left(t_{1}^{-}\right){\boldsymbol{b}}\right)=0.\end{dcases}$ (26)
 Assume that $u\equiv u_{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ on ${\mathcal{S}}_{2}$.
-Then, $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0$ since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$.
+Then, $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0$ since ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$.
 Since $\exists\varepsilon>0$, s.t.
-$\forall t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$, it has $\left.\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)\leq 0$.
+$\forall t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$, it holds that $\left.\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)\leq 0$.
 Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\left(u_{2}-u\left(t_{1}^{-}\right)\right)\leq 0$.
@@
 Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$; hence, $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)$.
-Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, Lemma 2 implies that $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
-$\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
+Since $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, Lemma 1 implies that $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
+$\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall r\in\left[\hat{r}_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$.
 Conversely, assume that $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{1}{\boldsymbol{x}}$ for $\left(t_{0},t_{1}\right)$ and $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+u_{2}{\boldsymbol{b}}$ for $\left(t_{1},t_{2}\right)$.
@@
 Then, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}=0$ and
-$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (28)
+$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (27)
 Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$.
 By Proposition 2, $\forall p\in\mathcal{P}_{1}$, $t\in\left[t_{0},t_{1}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$.
-So Proposition 3-3 holds.
-Similarly, Proposition 3-4 holds.
+So Proposition 4.3 holds.
+Similarly, Proposition 4.4 holds.
 ∎
@@
 Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection.
-Note that the equality constraints at the connection time, i.e., (12), (21), (23), and (25), are not independent since (12) implies the latter three equations for $1\leq r<r_{p}$.
-When examining feasibility near the connection time, it is sufficient to consider all inequality constraints and the maximal linearly independent set of the above equality constraints.
-Furthermore, Proposition 3 temporarily does not consider constraints on $u$ and ${\boldsymbol{x}}$, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ which will be discussed in Section IV-A.
-IV Keypoints for the Feasibility of the Optimal Trajectory
-The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, but it fails to fully present information on the feasibility during an arc.
-To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of the “keypoints” in an arc.
-Section IV-A analyzes some constraints of the end of an arc besides Proposition 3.
-The tangent marker is proposed in Section IV-B to describe the feasibility of the optimal trajectory tangent to the constrained boundary.
+Note that the equality constraints at the connection time, i.e., (11), (20), (22), and (24), are allowed linearly dependent.
+When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
+IV Keypoints for the Trajectory’s Feasibility
+The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, while the feasibility during an arc remains inadequately investigated.
+To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of keypoints in an arc.
+As formally defined in Definition 5, a keypoint, where the trajectory touches state constraints’ boundaries in isolation, directly influences the feasibility of the perturbed trajectory.
+For the example shown in Fig.
+1, the time points at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, $\mathcal{T}_{2}^{(2)}$, and ends of arcs are keypoints.
+Section IV-A analyzes additional constraints of an arc’s end besides Proposition 4.
+The tangent marker is proposed in Section IV-B to describe the feasibility of a BBS trajectory tangent to constraints’ boundaries.
 Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
@@
 Proposition 5.
-Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
+Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
   1. 1.
@@
       ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$.
-$\exists 1\leq\hat{r}_{p}\leq n$, s.t.
-$\forall 1\leq r<\hat{r}_{p}$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
-    Similar conclusions hold for ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}^{-}\right)+d_{p}\leq 0$.
+Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
+$\forall r\in\left[\hat{r}_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
+    Similar conclusions hold for $\left[t_{1}-\varepsilon,t_{1}\right]$.
   2. 2.
@@
       $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
-$\exists 1\leq\hat{r}_{p}\leq n$, s.t.
-$\forall 1\leq r<\hat{r}_{p}$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
+Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
+$\forall r\in\left[\hat{r}_{p}-1\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
     3. (c)
-      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$, and $\forall 1\leq r\leq n$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
-    Similar conclusions hold for $\left|u\left(t_{0}^{+}\right)\right|,\left|u\left(t_{1}^{-}\right)\right|\leq{u_{\mathrm{m}}}$.
+      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
+Furthermore, $\forall r\in\left[n\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
+    Similar conclusions hold for $u\geq-{u_{\mathrm{m}}}$ and $\left[t_{1}-\varepsilon,t_{1}\right]$.
 Proof.
-Proposition 4-1 holds for the same reason as that of Proposition 3.
-Consider Proposition 4-2.
+Proposition 5.1 holds for the same reason as that of Proposition 4.
+Consider Proposition 5.2.
 Proposition 2 implies that $u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}$.
@@
 $u\left(t\right)\equiv{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$.
-By Proposition 2, (1) is equivalent to $\exists 0\leq\hat{r}_{p}\leq n$, s.t.
-$\forall 0\leq r<\hat{r}_{p}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$.
-(2) is equivalent to $\forall 0\leq r\leq n$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
-Hence, Proposition 4-2 holds.
+By Proposition 2, (1) is equivalent to $\exists\hat{r}_{p}\in\left[0,n\right]\cap{\mathbb{N}}$, s.t.
+$\forall r\in\left[0,\hat{r}_{p}-1\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$.
+(2) is equivalent to $\forall r\in\left[0,n\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
+Hence, Proposition 5.2 holds.
 ∎
-Remark.
-Proposition 4 encompasses discussions on the cases covered by Proposition 3, while Proposition 3 provides more details for $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
-Based on Proposition 4, the additional end-constraint is defined as follows:
+${\mathcal{L}}=\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(1)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(2)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{2}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{(N)}\right\}_{k=1}^{M_{N}}\right)$ (28)
+Consider two connected arcs ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
+Proposition 4 guarantees the feasibility of constraints active along ${\mathcal{S}}_{1}$ or ${\mathcal{S}}_{2}$.
+If some constraints are active at the ends of arcs but are inactive in ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ a.e., then additional constraints at the connection, as defined in Definition 3, should be introduced for feasibility based on Proposition 5.
 Definition 3.
-Assume that ${\mathcal{S}}$ is a given arc for $t\in\left[t_{0},t_{1}\right]$.
-At an end of ${\mathcal{S}}$, i.e., $t_{i}$, $i\in\left\{0,1\right\}$, ${\mathcal{S}}$ and its adjacent arc, if it exists, correspond to the active constraints $\mathcal{P}$ and $\mathcal{P}^{\prime}$.
-$\mathcal{P}$ and $\mathcal{P}^{\prime}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{i}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank.
-The additional end-constraint of ${\boldsymbol{x}}\left(t_{i}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$.
-Among them, $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{i}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by Proposition 4 for the active constraints of ${\mathcal{S}}$ and its adjacent arc at $t_{i}$, except for those in $\mathcal{P}\cup\mathcal{P}^{\prime}$.
-$\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
+Consider two arcs ${\mathcal{S}}_{i}$ with active constraints $\mathcal{P}_{i}$ for $t\in\left[t_{i-1},t_{i}\right]$, $i\in\left\{1,2\right\}$.
+$\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{1}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank.
+By Proposition 5, state constraints except $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ and control constraints induce $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
+Then, the additional end-constraint of ${\boldsymbol{x}}\left(t_{1}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$.
+Furthermore, if $\widehat{{\boldsymbol{C}}}$ and $\widehat{{\boldsymbol{F}}}$ are empty matrices, then ${\mathcal{E}}$ is called empty.
 Remark.
 Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 5 is directly determined.
-In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}\cup\mathcal{P}^{\prime}$.
-To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}\cup\mathcal{P}^{\prime}$ are made linearly independent through elimination.
+In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
+To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ are made linearly independent through elimination.
 The elimination process is guaranteed by the feasibility of the given trajectory.
@@
 Proposition 4 and Proposition 5 analyzed the feasibility of an arc near the end points.
-An arc might be tangent to some constrained boundary, where light perturbations may affect the feasibility of the arc.
-The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to the constrained boundary.
+An arc might be tangent to some constrained boundary, where small perturbations may affect the feasibility of the arc.
+The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to constraints’ boundaries.
 Proposition 6.
-Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is an arc for $t\in\left[t_{0},t_{2}\right]$.
-${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists 1\leq\hat{r}\leq n$ is even, s.t.
-$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall 1\leq r<\hat{r},\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
+Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $\left[t_{0},t_{2}\right]$.
+${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists\hat{r}\in\left[n\right]$ is even, s.t.
+$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall r\in\left[\hat{r}-1\right],\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
 Proof.
@@
 Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)={\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}\right)$.
-Assume that $\forall 1\leq r\leq n$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$.
+Assume that $\forall r\in\left[n\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$.
 By Lemma 1, $\forall t\in B_{\varepsilon}\left(t_{1}\right)$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d\equiv 0$, which contradicts the tangent condition.
-Hence, $\exists 1\leq\hat{r}\leq n$, s.t.
-$\forall 1\leq r<\hat{r}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
+Hence, $\exists\hat{r}\in\left[n\right]$, s.t.
+$\forall r\in\left[\hat{r}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
 If $\hat{r}$ is odd, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ strictly crosses 00 at $t_{1}$, which contradicts the tangent condition.
@@
 ∎
-No matter the arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is constrained or unconstrained, ${\mathcal{S}}$ can be tangent to the boundary of the state constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$.
+No matter an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained or not, ${\mathcal{S}}$ can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$.
 It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$.
@@
 Hence, Proposition 6 can apply for both state constraints and control constraints.
-For convenience of discussion, $\forall 1\leq p\leq P$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$.
-The $\left(P+1\right)$-th constraint refers to $u\leq{u_{\mathrm{m}}}$, and the $\left(P+2\right)$-th constraint refers to $-u\leq{u_{\mathrm{m}}}$.
+For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$.
+The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$, respectively.
 Another noteworthy fact is that an arc ${\mathcal{S}}$ can be tangent to multiple constraints $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$ at a time point $t_{1}$.
@@
 Definition 4.
-Assume that $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$, an arc ${\mathcal{S}}$ is tangent to the $p$-th constraint at $t_{1}$.
-Denote the tangent marker as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$.
-$\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by (29), where $\widehat{{\boldsymbol{C}}}$ has full row rank.
+Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $t\in\left[t_{0},t_{2}\right]$.
+Assume that at $t_{1}\in\left(t_{0},t_{2}\right)$, the arc ${\mathcal{S}}$ is tangent to some constraints’ boundaries whose indices constitute a nonempty set $\mathcal{P}^{\prime}\subset\left[P+2\right]$.
+By Proposition 6, the tangent condition (29) is equivalent to $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
+Then, the tangent marker is defined as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$.
 Remark.
@@
 Definition 5.
-Assume that the optimal trajectory of problem (1a) satisfies Assumption 1; hence, it contains a finite number of arcs.
-The augmented switching law is defined as ${\mathcal{L}}={\mathcal{E}}_{0}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{\left(1\right)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{\left(N\right)}\right\}_{k=1}^{M_{N}}\right){\mathcal{E}}_{N}$.
-$\forall i\in\left[N\right]$, ${\mathcal{S}}_{i}$ occurs on $\left[t_{0}^{\left(i\right)},t_{0}^{\left(i+1\right)}\right]$.
-$\forall k\in\left[M_{i}\right]$, the tangent marker ${\mathcal{T}}_{k}^{\left(i\right)}$ occurs at $t_{k}^{\left(i\right)}$ during ${\mathcal{S}}_{i}$, where $t_{0}^{\left(i\right)}<t_{1}^{\left(i\right)}<t_{2}^{\left(i\right)}\dots<t_{M_{i}}^{\left(i\right)}<t_{0}^{\left(i+1\right)}$.
-The additional end-constraint ${\mathcal{E}}_{i}$ is induced by Proposition 4 at $t_{0}^{\left(i\right)}$.
-Among them, ${\mathcal{E}}_{N}$ additionally contains the constraints ${\boldsymbol{x}}\left(t_{0}^{\left(N+1\right)}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$, where $t_{0}^{\left(N+1\right)}={t_{\mathrm{f}}}$.
-$\forall i\in\left[N+1\right]$, $k\in\left[M_{i}\right]\cup\left\{0\right\}$, the points $t_{k}^{\left(i\right)}$ are called keypoints for the feasibility of the trajectory.
+Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$.
+During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$.
+At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$.
+Then, time points at tangent markers and ends of arcs are called keypoints for the trajectory’s feasibility.
+The augmented switching law (ASL) ${\mathcal{L}}$ is defined in (28).
 Remark.
-If a feasible trajectory with a finite number of arcs follows the Bang-Singular-Bang control law (7), then Proposition 3, Proposition 4, and Proposition 5 hold for the trajectory.
-For this case, the augmented switching law can also represent the feasible trajectory.
+If a feasible trajectory with a finite number of arcs follows the BBS control law (7), then Propositions 4, 5, and 6 hold for the trajectory.
+For this case, the ASL can also represent the feasible trajectory.
 Remark.
-If a non-chattering Bang-Singular-Bang feasible trajectory is disturbed by a sufficiently small perturbation, then the disturbed trajectory can only exceed the constrained boundary at the keypoints since the distance between the original trajectory and the constrained boundary is 0 at the keypoints, while it is strictly positive at other points.
-Therefore, the augmented switching law can not only characterize the input control but also describe the feasibility of the trajectory.
+If a feasible non-chattering BBS trajectory is perturbed by a sufficiently small perturbation, then the perturbed trajectory can only exceed constraints’ boundaries at the keypoints since the distance between the original trajectory and constraints’ boundaries is 0 at the keypoints, while it is strictly positive at other points.
+Therefore, the ASL can not only characterize the input control but also describe the feasibility of the trajectory.
 Remark.
-The switching law in Definition 2 corresponding to the augmented switching law in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
-It can be observed that the augmented switching law provides more detailed information on the feasibility of the trajectory at keypoints than the switching law.
-Notably, neither the switching law nor the augmented switching law provides the motion time of each arc.
-To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an augmented switching law.
-If no additional end-constraint is induced at the end of an arc, i.e., ${\mathcal{E}}=\left(\sim,\sim,\sim,\sim\right)$, then ${\mathcal{E}}$ is omitted.
-For convenience of understanding, the chain-of-integrators system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [8, 18].
+The switching law in Definition 2 corresponding to the ASL in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
+It can be observed that the ASL provides more detailed information on the feasibility of the trajectory at keypoints than the switching law.
+Notably, neither the switching law nor the ASL provides the motion time of each arc.
+To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an ASL.
+If ${\mathcal{E}}$ is empty, then ${\mathcal{E}}$ is omitted.
+The additional end-constraints before ${\mathcal{S}}_{1}$ and after ${\mathcal{S}}_{N}$ are omitted since the initial and terminal states are given, i.e., ${\boldsymbol{x}}_{0}$ and ${\boldsymbol{x}}_{\mathrm{f}}$.
+For the example shown in Fig.
+1, the ASL is $\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{1}^{(1)}\right\}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{2},\left\{{\mathcal{T}}_{2}^{(1)},{\mathcal{T}}_{2}^{(2)}\right\}\right){\mathcal{S}}_{3}$.
+For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 32].
 The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively.
 The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively.
-Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall 1\leq j<k$, $x_{j}\equiv 0$.
-The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of an even order $2l<k$ in Proposition 5 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively.
-The additional end-constraint ${\mathcal{E}}=\left(\left(s_{k},r_{k}\right)\right)_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 4, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively.
+Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$.
+The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively.
+The additional end-constraint =${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively.
 Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc.
-Based on the above simplified notations, an example is provided to illustrate the augmented switching law.
+Based on the above simplified notations, an example is provided to illustrate the ASL.
 Example 3.
-Consider a position-to-position problem for 4th order chain-of-integrators systems with full box constraints shown in Fig.
-1.
-A feasible suboptimal trajectory with a Bang-Singular-Bang form, planned by the manifold-intercepted method (MIM) [8], is shown in Fig.
-1(a).
-It can be observed that the augmented switching law is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
-Figure 1: Two examples of augmented switching laws.
-For a 4th order chain-of-integrators system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$.
-(a) A non-chattering feasible trajectory planned by our previous work [8].
-(b) A chattering optimal trajectory planned by our previous work [18].
-In (b2), $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
+Consider a position-to-position problem for 4th-order chain-of-integrator systems with full box constraints shown in Fig.
+2.
+A feasible suboptimal trajectory with a BBS form, planned by the manifold-intercepted method (MIM) [20], is shown in Fig.
+2(a).
+It can be observed that the ASL is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
+Figure 2: Two examples of ASLs.
+For a 4th-order chain-of-integrator system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$.
+(a) A non-chattering feasible trajectory planned by [20].
+(b) A chattering optimal trajectory planned by [32].
+(b2) is the plot of $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
 Example 4.
@@
 The chattering optimal trajectory is shown in Fig.
-1(b1).
+2(b1).
 Fig.
-1(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied.
+2(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied.
 It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$.
-According to Bellman’s principle of optimality [20], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is small enough.
+According to Bellman’s principle of optimality [22], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is sufficiently small.
 The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
-V A State-Centric Necessary Condition for Optimal Control of Problem (1a)
-The theoretical framework for the augmented switching law of time-optimal control for linear systems is established in Section IV.
+V A State-Centric Necessary Condition
+The theoretical framework for the ASL of time-optimal control for linear systems is established in Section IV.
 Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a).
 Section V-A discusses the uniqueness of the time-optimal control.
-Then, given a feasible Bang-Singular-Bang trajectory, Section V-B analyzes the feasibility of the disturbed Bang-Singular-Bang trajectory.
-Based on the conclusions in Section V-A and Section V-B, Section V-C provides a necessary condition for the optimal trajectory represented by the augmented switching law, which constrains the difference between the switching times and the sum of orders of keypoints in a way.
+Then, given a feasible BBS trajectory, Section V-B analyzes the feasibility of the perturbed BBS trajectory.
+Based on the conclusions in Sections V-A and V-B, Section V-C provides a necessary condition of optimality.
 V-A Uniqueness of Time-Optimal Control
@@
 Theorem 1.
-Under Assumption 1, the time-optimal control of problem (1a) is unique in an almost everywhere sense.
-In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal controls of problem (1a), then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
+Under Assumption 1, the time-optimal control of problem (1a) is unique in an a.e.
+sense.
+In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal, then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e.; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
 Proof.
-If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then evidently it has $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere since $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{0}$.
+If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then it holds a.e.
+that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ since ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{0}$.
 Assume that $\exists t^{\prime}\in\left(0,t_{\mathrm{f}}^{*}\right)$, s.t.
@@
 Then, $t_{0}>0$.
-Due to the continuity of ${\boldsymbol{x}}$, $\forall 0\leq t<t_{0}$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$, which contradicts the assumption.
+Due to the continuity of ${\boldsymbol{x}}$, $\forall t\in\left(0,t_{0}\right)$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$, which contradicts the assumption.
 Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ holds.
-Furthermore, $\forall t\in\left[0,t_{0}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ for $t\in\left[0,t_{0}\right]$ almost everywhere.
+Furthermore, $\forall t\in\left[0,t_{0}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ for $t\in\left[0,t_{0}\right]$ a.e.
 According to Bellman’s principle of optimality [22], ${\boldsymbol{x}}={\boldsymbol{x}}_{1}^{*}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}_{2}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ are still the time-optimal trajectories between ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{\mathrm{f}}^{*}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{\mathrm{f}}^{*}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$.
 Let $u_{3}^{*}\left(t\right)=\frac{1}{2}\left(u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)\right)$ and ${\boldsymbol{x}}_{3}^{*}\left(t\right)=\frac{1}{2}\left({\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)\right)$.
-Evidently, $u_{3}^{*}\left(t\right)$ is feasible for problem (1a) since the feasible set of $\left(u,{\boldsymbol{x}}\right)$ is convex; hence, $u_{3}^{*}\left(t\right)$ is optimal.
-Denote that $\forall k=1,2$, the switching law of ${\boldsymbol{x}}={\boldsymbol{x}}_{i}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ is ${\mathcal{L}}_{i}={\mathcal{S}}_{1}^{\left(i\right)}{\mathcal{S}}_{2}^{\left(i\right)}\dots\mathcal{S}_{N_{i}}^{\left(i\right)}$, where $\forall j\in\left[N_{i}\right]$, ${\mathcal{S}}_{j}^{\left(i\right)}=\left(\widehat{{\boldsymbol{A}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{b}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{F}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{g}}}_{j}^{\left(i\right)},\mathcal{P}_{j}^{\left(i\right)}\right)$ lasts for $t\in\left(t_{j-1}^{\left(i\right)},t_{j}^{\left(i\right)}\right)$.
-Among them, $t_{0}^{\left(i\right)}=t_{0}$ and $t_{N_{i}}^{\left(i\right)}=t_{\mathrm{f}}^{*}$.
-Evidently, if ${\mathcal{S}}_{1}^{\left(1\right)}={\mathcal{S}}_{1}^{\left(2\right)}$, then $\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$.
-Hence, ${\mathcal{S}}_{1}^{\left(1\right)}\not={\mathcal{S}}_{1}^{\left(2\right)}$.
-Assume that both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are unconstrained arcs.
-Without loss of generality, assume that $u_{1}^{*}\left(t\right)\equiv{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(1\right)}$; hence, $u_{2}^{*}\left(t\right)\equiv-{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(2\right)}$.
+Evidently, $u_{3}^{*}\left(t\right)$ is feasible since the feasible set of $\left(u,{\boldsymbol{x}}\right)$ is convex; hence, $u_{3}^{*}\left(t\right)$ is optimal.
+Denote that $\forall k=1,2$, the switching law of ${\boldsymbol{x}}={\boldsymbol{x}}_{i}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ is ${\mathcal{L}}_{i}={\mathcal{S}}_{1}^{(i)}{\mathcal{S}}_{2}^{(i)}\dots\mathcal{S}_{N_{i}}^{(i)}$, where $\forall j\in\left[N_{i}\right]$, ${\mathcal{S}}_{j}^{(i)}=\left(\widehat{{\boldsymbol{A}}}_{j}^{(i)},\widehat{{\boldsymbol{b}}}_{j}^{(i)},{\boldsymbol{F}}_{j}^{(i)},{\boldsymbol{g}}_{j}^{(i)},\mathcal{P}_{j}^{(i)}\right)$ lasts for $t\in\left(t_{j-1}^{(i)},t_{j}^{(i)}\right)$.
+Among them, $t_{0}^{(i)}=t_{0}$ and $t_{N_{i}}^{(i)}=t_{\mathrm{f}}^{*}$.
+Evidently, if ${\mathcal{S}}_{1}^{(1)}={\mathcal{S}}_{1}^{(2)}$, then $\forall t\in\left(t_{0},\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}\right)$, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$.
+So ${\mathcal{S}}_{1}^{(1)}\not={\mathcal{S}}_{1}^{(2)}$.
+Assume that both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are unconstrained arcs.
+Without loss of generality, assume that $u_{1}^{*}\left(t\right)\equiv{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{(1)}$; hence, $u_{2}^{*}\left(t\right)\equiv-{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{(2)}$.
 Then, $\exists\varepsilon>0$, s.t.
@@
 According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the fact that $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\equiv 0$.
-Therefore, one of ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ is constrained.
-Without loss of generality, assume that ${\mathcal{S}}_{1}^{\left(1\right)}$ is constrained.
-Assume that ${\mathcal{S}}_{1}^{\left(2\right)}$ is an unconstrained arc.
-Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t.
+Therefore, one of ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ is constrained.
+Without loss of generality, assume that ${\mathcal{S}}_{1}^{(1)}$ is constrained.
+Assume that ${\mathcal{S}}_{1}^{(2)}$ is an unconstrained arc.
+Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}-t_{0}\right)$, s.t.
 $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$.
-Since ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}}$, it has ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$.
+Since ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}}$, it holds that ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$.
 In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$.
@@
 Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ for $\left[t_{0},t_{0}+\varepsilon\right]$, which contradicts the definition of $t_{0}$.
-Therefore, both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are constrained arcs.
-Assume that $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}\not=\varnothing$.
-Arbitrarily consider $p\in\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}$.
-$\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$.
+Therefore, both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are constrained arcs.
+Assume that $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}\not=\varnothing$.
+Arbitrarily consider $p\in\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}$.
+$\forall t\in\left(t_{0},\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$.
 ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$.
-Therefore, $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}=\varnothing$.
-Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t.
-$\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{\left(i\right)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$.
+Therefore, $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}=\varnothing$.
+Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}-t_{0}\right)$, s.t.
+$\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{(i)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$.
 Hence, $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{3}^{*}\left(t\right)+d_{p}={\boldsymbol{c}}_{p}^{\top}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+u{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+d_{p}<0$.
-According to Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$.
-Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere.
+By Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$.
+Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e.
 ∎
 Remark.
-Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an almost everywhere sense when there exists at most one chattering limit point.
+Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an a.e.
+sense when there exists at most one chattering limit point.
 The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
-According to Theorem 1, if the time-optimal trajectory is disturbed by a sufficiently small perturbation, then the resulting disturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time.
+According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time.
 The above observation will be applied in Section V-B.
-V-B Feasibility of the Disturbed Bang-Singular-Bang Trajectory
+V-B Feasibility of the Perturbed BBS Trajectory
 […]
-The feasibility of the disturbed trajectory is proved by two steps.
-Firstly, the consistent error between the disturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory.
-In this way, the disturbed trajectory can be proved feasible at the points far from keypoints.
-Secondly, the feasibility of the disturbed trajectory near the keypoints is proved based on Proposition 3, Proposition 4, and Proposition 5.
-As a result, the feasibility of the disturbed trajectory is proved.
-The two steps correspond to Proposition 6 and Theorem 2, respectively.
+The feasibility of the perturbed trajectory is proved by two steps.
+Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory.
+In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory.
+Secondly, Theorem 2 proves the perturbed trajectory’s feasibility based on Propositions 4, 5, 6, and 7.
+In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible.
+Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
 Proposition 7.
-${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$.
-$\exists C>0$, s.t.
-$\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$,
-$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon,$ (30a)
-$\max_{i\in\left[M\right]}\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|\leq C\varepsilon.$ (30b)
-Among them, $\forall i\in\left[M\right]$, $t\in\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$.
-${\boldsymbol{x}}^{\prime}\left(t_{0}\right)={\boldsymbol{x}}\left(t_{0}\right)$ and $t_{0}^{\prime}=t_{0}$.
-${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$ and ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$.
-$0<\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$.
+Consider a BBS trajectory ${\boldsymbol{x}}\left(t\right)$ satisfying $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$.
+Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$.
+$\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
+$\forall\varepsilon\in\left(0,\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}\right)$, $\forall{\boldsymbol{x}}^{\prime}\left(t\right)$ is a perturbed trajectory induced by ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ with $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$,
+$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon$ (30)
+where $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$ and $t_{0}^{\prime}=t_{0}$.
 Proof.
-Since $\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$ and $\left|t_{i}-t_{i}^{\prime}\right|\leq\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, it has $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$.
-$\forall i\in\left[M\right]$, let $\hat{t}_{2i-1}=\min\left\{t_{i},t_{i}^{\prime}\right\}$, $\hat{t}_{2i}=\max\left\{t_{i},t_{i}^{\prime}\right\}$, and $\hat{t}_{0}=0$.
+Since $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$, it holds that $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$.
+$\forall i\in\left[M\right]$, let $\hat{t}_{2i-1}=\min\left\{t_{i},t_{i}^{\prime}\right\}$, $\hat{t}_{2i}=\max\left\{t_{i},t_{i}^{\prime}\right\}$, and $\hat{t}_{0}=t_{0}$.
 Then $\left\{\hat{t}_{i}\right\}_{i=0}^{2M}$ increases monotonically.
@@
 $\begin{dcases}{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau,\\ {\boldsymbol{x}}^{\prime}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.\end{dcases}$ (31)
-Therefore,
-$\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (32)
-$=$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\left({\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right)\right\|$
-$\leq$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$
+Therefore, $\forall t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$, it holds that
+$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$ (32)
 $\leq$ $C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|.$
@@
 $\begin{dcases}{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)+\int_{\hat{t}_{2i-1}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\tau\right)}{\boldsymbol{b}}_{j}\mathrm{d}\tau,\\ {\boldsymbol{x}}^{\prime}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)+\int_{\hat{t}_{2i-1}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\tau\right)}{\boldsymbol{b}}_{j^{\prime}}\mathrm{d}\tau.\end{dcases}$ (34)
-Therefore,
-$\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (35)
-$\leq$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)\right\|$
-$+$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|$
-$+$ $\varepsilon\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j^{\prime}}\right\|$
+Therefore, $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$, it holds that
+$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (35)
+$\leq$ $\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)\right\|$
+$+$ $\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|$
+$+$ $\varepsilon\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j^{\prime}}\right\|$
 $\leq$ $2\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon+C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|.$
@@
 $\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\leq 2\left(C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$ (36)
-Therefore, (30a) holds. Similarly, $\forall i\in\left[M\right]$,
-$\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|$ (37)
-$\leq$ $\left\|{\boldsymbol{x}}\left(t_{i}^{\prime}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|+\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}\left(t_{i}^{\prime}\right)\right\|$
-$\leq$ $\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+C_{1}\left\|{\boldsymbol{x}}\right\|\varepsilon+C_{2}C_{3}\varepsilon$
-$\leq$ $\left(2C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$
-Therefore, (30b) holds. ∎
+Therefore, (30) holds.
 Theorem 2.
-${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are Bang-Singular-Bang trajectories of problem (1a) without chattering.
-Assume that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are represented by the same augmented switching law ${\mathcal{L}}={\mathcal{L}}^{\prime}$.
-${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is feasible for problem (1a).
-$\exists\varepsilon>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
-if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
+Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is ${\mathcal{L}}$ in problem (1a).
+Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
+$\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is trajectory also represented by ${\mathcal{L}}^{\prime}={\mathcal{L}}$, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
+${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by ${\boldsymbol{t}}$ and ${\boldsymbol{t}}^{\prime}$ in the same way as that in Proposition 7.
 Proof.
@@
 According to Proposition 7, $\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
-if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then (30a) holds.
+if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then (30) holds.
 Secondly, consider the inequality constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\leq 0$.
-Denote that $\left\{i^{\left(p\right)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t.
-$i^{\left(p\right)}_{0}=0$, $i^{\left(p\right)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds.
-If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ since ${\mathcal{L}}={\mathcal{L}}^{\prime}$.
-Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$.
-Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$, and then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{\left(p\right)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{\left(p\right)}}\right)$ is continuous in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$.
-If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i^{\left(p\right)}_{j-1}}\right)+d_{p}<0$, then let $r_{j}^{\left(p\right)}=0$; otherwise, according to Proposition 3, Proposition 4, and Proposition 5, $\exists 1\leq r_{j}^{\left(p\right)}\leq n$, s.t.
-$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$.
-Let $t_{1,j}^{\left(p\right)}>t_{i^{\left(p\right)}_{j-1}}$ is small enough, s.t.
-$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$ and $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)\leq\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+1$ for $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$.
-$t_{2,j}^{\left(p\right)}<t_{i^{\left(p\right)}_{j}}$ is constructed similarly.
-Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}\right)$.
-In both case, $\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ due to continuity.
-Let
-$\varepsilon<-\frac{\sup_{t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}}{2C\left\|{\boldsymbol{c}}_{p}\right\|}.$ (39)
-Then, $\forall t\in\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$,
-${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}<0.$ (40)
-In other words, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible in $\left[t_{1,j}^{\left(p\right)},t_{2,j}^{\left(p\right)}\right]$.
-The feasibility of ${\boldsymbol{x}}^{\prime}$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)\cup\left(t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ needs further discussion.
-Let
-$\begin{dcases}\varepsilon<t_{1,j}^{\left(p\right)}-t_{i^{\left(p\right)}_{j-1}},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}+1\right)\right)^{-1},\\ \varepsilon<\left(C\left\|{\boldsymbol{c}}_{p}\right\|\left(\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}+1}+1\right)\right)^{-1},\\ \varepsilon<-\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\left(4\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+8\right)^{-1}.\end{dcases}$ (41)
-Denote $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}$. Then, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}},t_{1,j}^{\left(p\right)}\right]$,
-$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
-$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)$
-$=$ $\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
-$\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}-\varepsilon,t_{i^{\left(p\right)}_{j-1}}\right]$, it has
-$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (43)
-$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{\left(p\right)}}\right\|^{r_{j}^{\left(p\right)+1}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$
-$\leq$ $\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2.$
-Therefore, if $t_{i^{\left(p\right)}_{j-1}}^{\prime}\geq t_{i^{\left(p\right)}_{j-1}}$, then $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$.
-If $t_{i^{\left(p\right)}_{j-1}}^{\prime}<t_{i^{\left(p\right)}_{j-1}}$, then (30b) implies that $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\leq\frac{3}{4}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0$.
-$\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j-1}}\right]$,
-$\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t\right)$ (44)
-$\leq$ $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)+\left(\left|\frac{\mathrm{d}^{r_{j}^{\left(p\right)}+1}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}+1}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)\right|+2\right)\varepsilon$
-$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)<0.$
-Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, it has $\frac{\mathrm{d}^{r_{j}^{\left(p\right)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{\left(p\right)}}}\left(t\right)<0$.
-Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, $\forall 1\leq r<r_{j}^{\left(p\right)}$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{\left(p\right)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right)$.
-Therefore, $\forall t\in\left[t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{1,j}^{\left(p\right)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$.
-Similarly, let $\varepsilon$ is small enough, and it has ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ in $\left[t_{2,j}^{\left(p\right)},t_{i^{\left(p\right)}_{j}}^{\prime}\right]$.
-In summary, when $\varepsilon$ is small enough, $\forall p\in\left[P\right]$, $t\in\left[t_{0}^{\prime},t_{M}^{\prime}\right]$, it has ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ holds.
-Thirdly, the inequality $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is small enough. ∎
-Theorem 2 points out that if the disturbed trajectory is represented by the same augmented switching law to the original trajectory and the disturbance is light enough, then the disturbed trajectory is feasible.
-In other words, Theorem 2 provides a novel variational approach under fixed augmented switching laws with a feasibility guarantee near the constrained boundary.
-V-C Necessary Condition for the Optimal Trajectory
-Based on Theorem 2, this section provides an approach to constructing a feasible disturbed trajectory with the same augmented switching law.
-According to Theorem 1, this section shows that the existence of such a disturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
-The Jacobian matrix of equality constraints induced by the augmented switching law can be calculated by Proposition 7.
+Denote $\left\{i^{(p)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t.
+$i^{(p)}_{0}=0$, $i^{(p)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ switches between active and inactive at $t_{i^{(p)}_{j}}$.
+In other words, $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i_{j}^{(p)}}\right)+d_{p}=0$ holds.
+Furthermore, $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds.
+Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, the above property on $\left({\boldsymbol{x}}\left(t\right),{\boldsymbol{t}}\right)$ holds for $\left({\boldsymbol{x}}^{\prime}\left(t\right),{\boldsymbol{t}}^{\prime}\right)$ as well.
+Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig.
+3(a).
+${\mathcal{L}}={\mathcal{L}}^{\prime}$ implies that ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\equiv 0$ holds.
+Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right)$.
+Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig.
+3(b).
+The feasibility of ${\boldsymbol{x}}^{\prime}\left(t\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right]$ is discussed in two parts, i.e., far from and near the keypoints.
+Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$ and $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$.
+Then, $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ is continuous in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$.
+If $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ holds, then let $r_{j}^{(p)}=0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$; otherwise, according to Propositions 4, 5, and 6, $\exists r_{j}^{(p)}\in\left[n\right]$, s.t.
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ and $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$.
+Let $t_{1,j}^{(p)}\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$ is sufficiently small, s.t.
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ holds $\forall t\in\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$.
+$t_{2,j}^{(p)}\leq t_{i^{(p)}_{j}}$ is constructed similarly.
+Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{(p)},t_{i^{(p)}_{j}}\right)$.
+Due to the continuity of $f_{p}$, $\sup_{t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]}f_{p}\left(t\right)<0$ holds.
+For the part far from keypoints, i.e., $\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$, let
+$\varepsilon<-\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\right)^{-1}\sup\nolimits_{t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]}f_{p}\left(t\right).$ (38)
+Then, according to (30), $\forall t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$, it holds that
+$f_{p}^{\prime}\left(t\right)\leq f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}<0.$ (39)
+In other words, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]$.
+Consider the part near the left keypoint, i.e., $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$.
+Assume that $t_{i^{(p)}_{j-1}}^{\prime}<t_{1,j}^{(p)}$; otherwise, $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]=\varnothing$.
+For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$, let $\varepsilon$ be sufficiently small, s.t.
+$f_{p}\left(t\right)\leq\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$.
+By (30) and (38), $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$, $f_{p}^{\prime}\left(t\right)\leq 0$ is feasible since
+$f_{p}^{\prime}\left(t\right)\leq$ $f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (40)
+$\leq$ $\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)-\frac{1}{4}f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0.$
+Figure 3: Feasibility of the perturbed trajectory near the keypoints. (a) Constrained arc. (b) Unconstrained arc.
+For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)=0$ and $t_{i_{j-1}^{(p)}}<t_{1,j}^{(p)}$, let
+$\varepsilon<\hat{\varepsilon}\triangleq-\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}+1\right)^{-1}.$ (41)
+Hence, $\forall t\in\left[t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right]$, it holds that
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
+$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
+If $t_{i^{(p)}_{j-1}}^{\prime}<t_{i^{(p)}_{j-1}}$, let $\varepsilon$ be sufficiently small, s.t.
+$\sup\nolimits_{\delta\in\left[0,\varepsilon\right]}\left\|{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}-\delta\right)-{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)\right\|<\hat{\varepsilon}.$ (43)
+Then, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]\subset\left[t_{i^{(p)}_{j-1}}-\varepsilon,t_{i^{(p)}_{j-1}}\right]$, it holds that
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r_{j}^{(p)}-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ (44)
+$+$ $\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left(\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)\right\|\right)$
+$\leq$ $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
+Therefore, once $\varepsilon$ is sufficiently small, then $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}<0$ holds in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$.
+${\mathcal{L}}^{\prime}={\mathcal{L}}$ implies that $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right)$.
+Therefore, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible.
+Similarly, once $\varepsilon$ is sufficiently small, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible in $\left[t_{2,j}^{(p)},t_{i^{(p)}_{j}}^{\prime}\right]$.
+Note that the above conditions on $\varepsilon$ only depends on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$.
+In summary, once $\varepsilon$ is sufficiently small, then $\forall p\in\left[P\right]$, $t\in\left[t_{0}^{\prime},t_{M}^{\prime}\right]$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ holds.
+Thirdly, the control constraints $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is sufficiently small. ∎
+Theorem 2 points out that if the perturbed trajectory is represented by the same ASL to the original trajectory and the perturbance is small enough, then the perturbed trajectory is feasible.
+In other words, Theorem 2 provides a novel variational approach under the fixed ASL with a feasibility guarantee near constraints’ boundaries.
+V-C State-Centric Necessary Condition of Optimality
+Based on Theorem 2, this section provides an approach to constructing a feasible perturbed trajectory with the same ASL.
+According to Theorem 1, this section shows that the existence of such a perturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
+The Jacobian matrix of equality constraints induced by the ASL can be calculated through Proposition 8.
 Proposition 8.
@@
 Then, $\forall i,j\in\left[M\right]$,
-$\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}=\begin{dcases}\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\\ \qquad j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ 0,\,j>i,\end{dcases}$ (45)
-where $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}$ refers to
-${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}.$ (46)
+$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}{\boldsymbol{P}}_{j,i}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\,j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ {\boldsymbol{0}},\,j>i,\end{dcases}$ (45)
+where ${\boldsymbol{P}}_{j,i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}$.
 Proof.
-Evidently, $\forall j>i$, ${\boldsymbol{x}}_{i}$ is independent with $t_{j}$; hence, $\frac{\partial{\boldsymbol{x}}_{i}}{t_{j}}={\boldsymbol{0}}$. According to Lemma 1,
-${\boldsymbol{x}}_{i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\boldsymbol{x}}_{i-1}+\int_{t_{i-1}}^{t_{i}}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.$ (47)
-Therefore, $\forall i\in\left[M\right]$,
-$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}.$ (48)
-Applying (48), $\forall i>1$, it has
+$\forall j>i$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}={\boldsymbol{0}}$ holds evidently.
+According to Lemma 1, ${\boldsymbol{x}}_{i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\boldsymbol{x}}_{i-1}+\int_{t_{i-1}}^{t_{i}}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau$; hence, $\forall i\in\left[M\right]$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}$ holds.
+Then, $\forall i>1$, it holds that
 $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{i-1}}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}\left(\frac{\partial{\boldsymbol{x}}_{i-1}}{\partial t_{i-1}}-{\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i-1}-{\boldsymbol{b}}_{i}\right)$ (46)
 $=$ ${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}\left(\left({\boldsymbol{A}}_{i-1}-{\boldsymbol{A}}_{i}\right){\boldsymbol{x}}_{i-1}+{\boldsymbol{b}}_{i-1}-{\boldsymbol{b}}_{i}\right).$
-$\forall i>1$, $j<i-1$, it has
-$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ (50)
-$=$ $\prod_{k=i}^{j+1}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right).$
-Therefore, (45) holds. ∎
-Utilizing the notations in Section V-B, the augmented switching law provides a system of equalities on every keypoints.
+$\forall i>1$, $j<i-1$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ holds, which implies (45).
+∎
+Utilizing the notations in Section V-B, the ASL provides a system of equalities on every keypoints.
 Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank.
@@
 ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$.
-Then, Theorem 3 provides a necessary condition for the optimal trajectory.
-Before providing Theorem 3, the implicit function theorem [29] should be introduced.
-Lemma 3 (The Implicit Function Theorem [29]).
-Assume $S\subset{\mathbb{R}}^{p+q}$ is open and non-empty.
-${\boldsymbol{F}}:S\to{\mathbb{R}}^{q}$ is $C^{1}$ continuous.
-$\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\in S$, satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)={\boldsymbol{0}}$ and $\det\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}}_{0},{\boldsymbol{\eta}}_{0}\right)\not=0$.
-Then, $\exists\delta_{1},\delta_{2}>0$, s.t.
-$\forall{\boldsymbol{\xi}}\in B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)$, $\exists!{\boldsymbol{\eta}}\in B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$ satisfying ${\boldsymbol{F}}\left({\boldsymbol{\xi}},{\boldsymbol{\eta}}\right)={\boldsymbol{0}}$.
-The above relation induces a mapping ${\boldsymbol{f}}:B_{\delta_{1}}\left({\boldsymbol{\xi}}_{0}\right)\to B_{\delta_{2}}\left({\boldsymbol{\eta}}_{0}\right)$.
-Then, ${\boldsymbol{f}}$ is $C^{1}$ continuous, where $\frac{\mathrm{d}{\boldsymbol{f}}}{\mathrm{d}{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}}\right)=-\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\eta}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)^{-1}\frac{\partial{\boldsymbol{F}}}{\partial{\boldsymbol{\xi}}}\left({\boldsymbol{\xi}},{\boldsymbol{f}}\left({\boldsymbol{\xi}}\right)\right)$.
+Then, Theorem 3 provides a necessary condition of optimality.
 Theorem 3 (State-Centric Necessary Condition).
 Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1.
-Denote the equality constraints induced by the augmented switching law as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$.
+Denote the equality constraints induced by the ASL as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$.
 Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
@@
 ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)=M^{\prime}\leq M-1.$ (47)
-Assume that the $i$-th row of $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ is linearly independent where $i\in\mathcal{I}\subset\left[M-1\right]$ and #$\#\mathcal{I}=M^{\prime}$.
-Let $\varepsilon>0$ be small enough to satisfy the condition of Theorem 2 and Lemma 3.
+Assume that the $i$-th row of $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ is linearly independent where $i\in\mathcal{I}\subset\left[M-1\right]$ and $\left|\mathcal{I}\right|=M^{\prime}$.
+Let $\varepsilon>0$ be sufficiently small to satisfy the condition of Theorem 2 and the implicit function theorem [31].
 The implicit function ${\boldsymbol{f}}:\left(t_{i}\right)_{i\in{\left.\left[M\right]\middle\backslash\mathcal{I}\right.}}\mapsto\left(t_{i}\right)_{i\in\mathcal{I}}$ is induced by ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)={\boldsymbol{0}}$.
 Based on ${\boldsymbol{f}}$, $\exists{\boldsymbol{t}}^{\prime}$ satisfies ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$, $t_{M}^{\prime}<t_{M}^{*}$, and $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}^{*}\right\|<\varepsilon$.
-According to Theorem 2, the disturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible since ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$ implies that ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{*}$ have the same augmented switching law.
+According to Theorem 2, the perturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible since ${\boldsymbol{H}}\left({\boldsymbol{t}}^{\prime}\right)={\boldsymbol{0}}$ implies that ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{*}$ have the same ASL.
 However, ${\boldsymbol{x}}^{\prime}$ achieves a shorter terminal time $t_{M}^{\prime}<t_{M}^{*}$, which contradicts the optimality of ${\boldsymbol{x}}^{*}$.
@@
 Remark.
+In contrast to existing costate-based necessary conditions [9], Theorem 3 requires no information on costates, which significantly simplifies the computation.
+Compared to existing state-centric necessary conditions [24], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
+Some examples are provided in Sections VI-B and VI-C.
+Remark.
 […]
 Remark.
-In fact, if a given non-chattering Bang-Singular-Bang trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical local optimization of trajectories based on Theorem 3.
-For example, let $t_{M}$ be the optimization objective and consider the equality and inequality constraints of the augmented switching law as the constraints of the optimization problem.
-In this way, it is hopeful to obtain a locally optimal solution or a stationary point of problem (1a) through iteratively solving an $M$-dimensional nonlinear problem.
-Note that when applying shooting methods [30], problem (1a) is highly nonlinear and non-convex, and the long motion time may result in unacceptable computation cost.
-In contrast, the local optimization method on ${\boldsymbol{t}}$ based on Theorem 1a will be more efficient due to the significantly lower dimensions and the suboptimal feasible initial value.
-The above method will be developed in future works.
-Theorem 3 provides a necessary condition for the optimal control of problem (1a) under Assumption 1.
-Different from the existing costate-based necessary conditions [14], Theorem 3 requires no information on the costates, which significantly simplifies the computation of the necessary conditions.
-Compared to existing state-centric necessary conditions [22], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
-Some further discussions on the Jacobian matrix are necessary.
-Consider a keypoint $t_{i}$ with a constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ active at $t_{i}$.
-Assume that the constraint is of $\hat{r}$-th order w.r.t the arc $\left(t_{i-1},t_{i}\right)$ where $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$.
-In other words, if $1\leq\hat{r}<n$, then $\left(-1\right)^{\hat{r}}\frac{\mathrm{d}^{\hat{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{\hat{r}}}\left(t_{i}^{-}\right)>0$ and $\forall 1\leq r<\hat{r}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$.
-If $\hat{r}=n$, then $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)=0$.
-Among them.
-$\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\left(t_{i}^{-}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)$.
-According to Proposition 7, $\forall r\in{\mathbb{N}}$,
-$\frac{\partial}{\partial t_{i}}\left.\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r}}\right|_{t_{i}^{-}}=\frac{\partial}{\partial t_{i}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r-1}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)$ (52)
-$=$ $\left({\boldsymbol{c}}^{\top}{\boldsymbol{A}}_{i}^{r}\left({\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i}\right)\right)=\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}.$
-In other words, the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t.
-$t_{i}$ is $\left.\frac{\mathrm{d}^{r+1}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)}{\mathrm{d}t^{r+1}}\right|_{t_{i}^{-}}{\boldsymbol{e}}_{\hat{r}}$.
-Specifically, if $\hat{r}=n$, then the Jacobian matrix of ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq{\boldsymbol{0}}$ w.r.t.
-$t_{i}$ is ${\boldsymbol{0}}$.
-VI Applications in Chain-of-Integrators Systems and Simulation Experiments
+If a given non-chattering BBS trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical optimization of the trajectory based on Theorem 3.
+Numerical examples are provided in Section VII.
+VI Applications in Chain-of-Integrator Systems
 This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a).
-For convenience, the chain-of-integrators system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
-Time-optimal control for chain-of-integrators systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved.
+For convenience, the chain-of-integrator system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
+Time-optimal control for chain-of-integrator systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved.
 The problem can be summarized as follows:
 $\min\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (48a)
-${\mathrm{s.t.}}\quad$ $\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),\,\forall 1<k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53b)
+[…]
 $\dot{x}_{1}\left(t\right)=u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (48c)
 ${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}},$ (48d)
-$\left|x_{k}\left(t\right)\right|\leq x_{{\mathrm{m}}k},\,\forall 1\leq k\leq n,\,t\in\left[0,t_{\mathrm{f}}\right],$ (53e)
+[…]
 $\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (48f)
-The Jacobian matrix, except the last column induced by the augmented switching law, denoted as ${\boldsymbol{J}}$, can be calculated as follows:
-Corollary 1.
-Assume that $\forall i\in\left[N\right]$, $u\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}},0\right\}$ in $\left(t_{i-1},t_{i}\right)$, and denote ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$.
-Then,
-$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}-\Delta u_{j+1}{\boldsymbol{\phi}}\left(t_{i}-t_{j}\right),&j<i\\ \left[\begin{array}[]{c}u_{i}\\ {\boldsymbol{x}}_{i,1:\left(n-1\right)}\end{array}\right],&j=i\\ {\boldsymbol{0}},&j>i.\end{dcases}$ (54)
-Among them, $\Delta u_{i}=u_{i}-u_{i-1}$ and $u_{0}=0$.
-${\boldsymbol{\phi}}\left(t\right)\triangleq\left(\phi_{k}\left(t\right)\right)_{k=1}^{n}$ where $\phi_{k}\left(t\right)=\frac{t^{k-1}}{\left(k-1\right)!}$.
-Proof.
-Proposition 7 directly implies Corollary 1. ∎
 This section provides some applications of Theorem 3.
@@
 VI-B The Case where Only System Behaviors Occur
-The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to the constrained boundary, is widely applied in existing works on trajectory planning [31, 32, 8].
-Denote the augmented switching law of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$.
+The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to constraints’ boundaries, is widely applied in existing works on trajectory planning [19, 34, 20].
+Denote the ASL of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$.
 $\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$.
 The non-existence of tangent markers and additional end-constraints in ${\mathcal{L}}$ means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility.
-An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
+An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
 Corollary 2.
-Assume that the augmented switching law consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
+Assume that the ASL consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
   1. 1.
@@
     $\forall i\in\left[N\right]$, if $\left|{\mathcal{S}}_{i}\right|>0$ and $\forall i<j\leq N$, $\left|{\mathcal{S}}_{j}\right|<\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{N}\left|{\mathcal{S}}_{k}\right|\leq N-i$.
-Then,
+Then, it holds that
 $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|\leq n.$ (50)
+$\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
 Proof.
@@
 Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$.
-According to Corollary 1, ${\boldsymbol{J}}$ contains the following rows:
+According to Proposition 8, ${\boldsymbol{J}}$ contains the following rows:
 ${\boldsymbol{J}}=\left(-\Delta u_{j+1}{\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]},$ (53)
@@
 Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that ${\boldsymbol{A}}$ has full row rank if and only if ${\boldsymbol{B}}$ has full row rank.
-Since $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it has
+By $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it holds that
 ${\boldsymbol{J}}\leftrightarrow\left({\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}.$ (55)
@@
 Remark.
-Further research suggests that Conditions 1, 2, and 3 in Corollary 3 might hold almost everywhere.
-In other words, $\left({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{\mathrm{f}},{\boldsymbol{x}}_{\mathrm{m}},{u_{\mathrm{m}}}\right)$ whose resulting Bang-Singular-Bang trajectory does not satisfy Conditions 1, 2, and 3 forms a zero-measure set.
-Remark.
-In our previous work [8], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the augmented switching law ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$ locally forms a submanifold of $r$-dimensions.
+In our previous work [20], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the ASL ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$ locally forms a submanifold of dimension $r$.
 Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$.
 However, the local optimality is not discussed in [20].
-Corollary 3 proves that the augmented switching law of dimensions $r>n$ fails to achieve optimality.
-${\boldsymbol{J}}=\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
-$\left[\begin{array}[]{cc}\left(\delta_{j=i}-\delta_{j=i-1}\right)_{i\in\left[N\right],j\in\left[N\right]}&\\ &{\boldsymbol{I}}_{n-1}\end{array}\right]\left[\begin{array}[]{cccc}{\boldsymbol{I}}_{N-1}&&&\\ &0&1&\\ &1&0&\\ &&&{\boldsymbol{I}}_{n-2}\end{array}\right]\left[\begin{array}[]{cc}{\boldsymbol{I}}_{N+1}&\\ &\left(\phi_{i+2-j}\left(t_{\infty}-t_{3N}\right)\right)_{i\in\left[n-2\right],j\in\left[n\right]}\end{array}\right]$ (63)
-It is challenging to reason Corollary 3 by existing costate-based necessary conditions [14] due to the complex behavior of the costates.
-In contrast, existing costate-based necessary conditions can try to reason the costate for a given trajectory case-by-case.
-For example, in a 4th order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$.
-$\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can serve as a candidate for the optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$.
-Of course, both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ cannot be optimal in a 3rd order problem.
-Though Corollary 3 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing the optimal control of problem (53a).
+Corollary 2 proves that the ASL with $r>n$ fails to achieve optimality.
+It is challenging to reason Corollary 2 by existing costate-based necessary conditions [9] due to complex behaviors of the costates.
+In contrast to Corollary 2, PMP-based approaches need to derive the costates for a given trajectory case-by-case.
+For example, in a 4th-order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$.
+$\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can be a candidate for optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$.
+Though Corollary 2 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing optimal control of problem (48a).
+Numerical examples are provided in Section VII.
 VI-C The Chattering Phenomenon induced by $\left|x_{2}\right|\leq x_{{\mathrm{m}}2}$
 Our previous work [32] points out that if the chattering phenomenon occurs in problem (48a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon.
-An infinite number of unconstrained arcs are joined by the constrained boundary, i.e., the unconstrained arcs are tangent to the boundary for infinite times.
+An infinite number of unconstrained arcs are joined by constraints’ boundaries, i.e., the unconstrained arcs are tangent to the boundary for infinite times.
 In contrast, constrained arcs do not occur.
 This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition.
-[18] provides the following conclusion.
+[32] provides the following lemma.
 Lemma 2.
@@
 $\forall k\in{\mathbb{N}}^{*}$, $u$ switches at most 2 times during $\left(t_{3k-3},t_{3k}\right)$.
-Theorem 3 and Lemma 4 imply the following corollary.
 Corollary 3.
@@
 $\begin{dcases}u\left(t\right)\equiv u_{3k-2}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-3},t_{3k-2}\right),\\ u\left(t\right)\equiv u_{3k-1}={u_{\mathrm{m}}},&t\in\left(t_{3k-2},t_{3k-1}\right),\\ u\left(t\right)\equiv u_{3k}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-1},t_{3k}\right).\end{dcases}$ (60)
-Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$. Hence, $\Delta u_{3k-1}=2{u_{\mathrm{m}}}$, $\Delta u_{3k}=-2{u_{\mathrm{m}}}$, and $\Delta u_{3k+1}=0$.
-$\forall N\in{\mathbb{N}}^{*}$, the augmented switching law between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$.
-According to Corollary 1, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (62).
-$\forall k\in\left[N-1\right]$, the $3k$-th columns and the $\left(2k-1\right)$-th rows can be eliminated, and the coefficients of each column, i.e., ±$\pm 2{u_{\mathrm{m}}}$, can be normalized, resulting in ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}_{1}$.
-Upon left multiplication by (63) and right multiplication by $\left(\delta_{j=i}-\delta_{j=i+1}\right)_{i\in\left[N\right],j\in\left[N\right]}^{3}$, it has ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$.
+Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$.
+$\forall N\in{\mathbb{N}}^{*}$, the ASL between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$.
+By Proposition 8, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (51).
+Through basic row and column transformations, it holds that ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$.
 Considering the rank of columns $\left(k-n+3\right)$ through $k$ of ${\boldsymbol{J}}^{\prime}$, (59) holds.
-∎
 Remark.
-According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ when applying (66).
-When given initial values, (66) is a polynomial equation w.r.t.
-$\tau_{k}$.
-It is evident that $n=3$ does not hold since it has $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$.
-Through more refined derivation, it can be rigorously proven that $n=4$ does not hold.
-The proof is lengthy and omitted due to space limitation, whose approach is similar to our previous work [18].
-In other words, the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ can only occur in a 5th order problem or higher.
-One can try to determine whether the chattering phenomenon can occur in an $n$-th order problem for $n\geq 5$ based on the proposed Corollary 3.
-Remark.
-It is challenging to obtain the recursive equation (66) for problem (53a) of high order based on the traditional costate-based necessary condition [14] since the behaviors of costates are significantly complex for high order problems.
-Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 4 provides an example to apply the proposed necessary condition to the case the chattering phenomenon occurs, where a sub-arc without chattering of the optimal trajectory is investigated.
-VI-D Simulation Experiment
-Figure 2: Results of the simulation experiment.
-(a) The plot of the local minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$.
-(b) The original trajectory and the optimized trajectory.
-Among them, ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, ${u_{\mathrm{m}}}=1$.
-A simulation experiment is conducted to verify the proposed state-centric necessary condition.
+According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ under the recursive equation (59).
+It is evident that chattering does not occur when $n=3$ since $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$ converges to $\infty$.
+Through more refined derivation, it can be rigorously proved that $n=4$ does not hold.
+In future works, one can try to determine the existence of chattering in problem (48a) of order $n\geq 5$ under constraints on $x_{2}$, where Corollary 3 and Theorem 3 would serve as useful mathematical tools.
+However, it is challenging to obtain the recursive equation (59) for problem (48a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering.
+A comparison between state-centric and PMP-based approaches is provided in Appendix A.
+Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 3 provides an example to apply the proposed necessary condition to a chattering trajectory, where a sub-arc without chattering of the optimal trajectory is investigated.
+VII Numerical Experiments
+Numerical experiments for high-order chain-of-integrator systems (48a) are conducted to provide a potential way to apply the proposed state-centric necessary condition to optimizing trajectories.
+Since this paper focuses on the theoretical necessary condition instead of optimization algorithms, metrics like computational time are not considered.
+Note that this paper does not provide a manual optimization algorithm.
+VII-A Representating a Feasible BBS Trajectory
+Consider problem (1a) whose system matrix is non-diagonalizable with multiple Jordan blocks as follows:
+${\boldsymbol{A}}=\left[\begin{array}[]{cccc}-1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right],\,{\boldsymbol{b}}=\left[\begin{array}[]{c}0\\ 0\\ 2\\ -1\end{array}\right].$ (61)
+The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$.
+As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint.
+As shown in Figs.
+4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper.
+The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$.
+Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$.
+${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$.
+${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$.
+${\mathcal{T}}_{5}^{(1)}$ is a tangent markers where $x_{1}$ is tangent to $-0.7$.
+${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$.
+In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$.
+In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$.
+One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
+A numerical optimal control toolbox, i.e., Yop [35], is applied to solving problem (61).
+The number of intervals is set to 200.
+The resulting trajectory without a specified initial solution ${\boldsymbol{x}}_{\text{Yop1}}$ is significantly far from optimal, as shown in Fig.
+4.
+If setting ${\boldsymbol{x}}_{\text{ASL}}$ as the initial solution, then a near-optimal solution ${\boldsymbol{x}}_{\text{Yop2}}$ is obtained.
+The terminal times of ${\boldsymbol{x}}_{\text{ASL}}$, ${\boldsymbol{x}}_{\text{Yop1}}$, and ${\boldsymbol{x}}_{\text{Yop2}}$ are 5.541, 6.896, and 5.624, respectively.
+Furthermore, costates can be reconstructed from states ${\boldsymbol{x}}_{\text{ASL}}$ through meticulous derivations, as shown in Fig.
+4(d).
+The information on costates also verifies the proposed state-centric necessary condition in a way, noting that PMP (7) holds in this example.
+In contrast, the costates of ${\boldsymbol{x}}_{\text{Yop1}}$ and ${\boldsymbol{x}}_{\text{Yop2}}$ do not exist since Yop does not solve a problem based on costates.
+Figure 4: A feasible BBS trajectory with system behaviors, tangent markers, and additional end-constraints.
+In (d), $\hat{\lambda}\triangleq\left(2\lambda_{3}-\lambda_{4}\right)\xi\left(t\right)$ where $\xi\left(t\right)>0$ serves as a scaling factor to enhance the visibility of ${\mathrm{sgn}}\left(2\lambda_{3}-\lambda_{4}\right)$.
+Figure 5: Optimizing a 4th-order near-optimal trajectory.
+(a) Plot of the locally minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$.
+(b) The original and the optimized trajectories.
+VII-B Determining the Optimality of a 4th-Order Trajectory
+A near-optimal trajectory planning method is proposed in our previous work [20].
+Consider a 4th-order trajectory for high-order chain-of-integrator systems (48a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig.
+5(b).
+The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$.
+Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal.
+Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$.
+Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank.
+Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (48a).
 As shown in Fig.
-2(b), our previous work [8] provided a 4th order suboptimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, $u=u\left(t\right)$, $t\in\left[t,t_{\mathrm{f}}\right]$.
-The augmented switching law of the original trajectory is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$.
-According to Corollary 3, $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$ implies that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is not optimal.
-Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$.
-Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns is full row rank.
-Then, disturb $t_{4}$ into $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (53a).
+5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$.
+It can be observed in Fig.
+5(b2) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory.
+Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint.
+Specifically, ${\boldsymbol{x}}^{\prime}$ achieves a 0.06% reduction in the terminal time compared to ${\boldsymbol{x}}$.
+Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
+VII-C Optimizing a 5th-Order Feasible BBS Trajectory
+Arc-based [36] and discretized optimization [37] methods are two main approaches to solving high-order problems.
+Consider a 5th-order problem for high-order chain-of-integrator systems (48a) where ${\boldsymbol{x}}_{0}={\boldsymbol{0}}$, ${{\boldsymbol{x}}_{\mathrm{f}}}={\boldsymbol{e}}_{5}$, $u_{\mathrm{m}}=1$, and ${\boldsymbol{x}}_{\mathrm{m}}=\left(0.8,0.5,0.5,0.5,1\right)$.
+A feasible BBS trajectory ${\boldsymbol{x}}_{1}$ is planned by [36], as shown in Fig.
+6(a).
+Denote $\mathrm{DOF}\triangleq N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|-n$; hence, Corollary 2 implies that $\mathrm{DOF}\leq 0$ holds for an optimal trajectory.
 As shown in Fig.
-2(a), $t_{\mathrm{f}}^{\prime}$ achieves local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$ shown in Fig.
-2(b).
-It can be observed in Fig.
-2(b4) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory.
-In other words, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint; hence, ${\boldsymbol{t}}^{\prime}$ cannot be further optimized under fixing ${\mathcal{L}}$.
-Among them,
-$\begin{dcases}{\boldsymbol{t}}=\left(1.7500,2.1562,3.1562,4.5342,5.5342,\right.\\ \quad\left.6.0342,6.5221,7.0100,7.2226,8.9165,9.8604\right),\\ {\boldsymbol{t}}^{\prime}=\left(1.7500,2.1562,3.1562,4.5550,5.5550,\right.\\ \quad\left.6.2023,6.6621,7.1218,7.2166,8.9104,9.8543\right).\end{dcases}$ (68)
-Therefore, ${\boldsymbol{x}}^{\prime}$ achieves a $0.6\%$ reduction in the terminal time compared to ${\boldsymbol{x}}$.
-The proposed state-centric necessary condition is verified through the simulation experiment.
-More efficient optimization algorithms based on the proposed state-centric necessary condition can be developed in future works.
-VII Conclusions
-This paper has set out to establish an innovative theoretical framework for the augmented switching law for time-optimal control of controllable linear systems and its applications.
-The proposed augmented switching law represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form.
-Specifically, the equality and inequality constraints induced by the augmented switching law represent a sufficient and necessary condition for the feasibility of the Bang-Singular-Bang trajectory.
-Based on the augmented switching law, a given feasible trajectory can be disturbed with the fixing augmented switching law and feasibility guarantees, resulting in a state-centric necessary condition for optimal control since any disturbed feasible trajectory should have a strictly longer terminal time than the optimal terminal time due to the uniqueness of optimal control.
-The proposed necessary condition states that the Jacobian matrix induced by the augmented switching law must not be full row rank.
-High order conditions can be derived based on the local optimality of the optimal trajectory.
-Different from the traditional necessary condition based on adjoining methods, the developed necessary condition requires only state and control information without dependence on costate information.
-The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrators systems with full box constraints.
-Under the condition that every arc is strictly feasible except for the constraints active during the whole arc, the proposed state-centric necessary condition implies that the number of arcs should be no greater than the sum of the order of constrained arcs plus the order of the problem.
-Considering the chattering phenomenon induced by the 2nd order state constraints, this paper proves a recursive equation for the junction time before the chattering limit point.
-The above two conclusions are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed state-centric necessary conditions.
+6(b), $\mathrm{DOF}_{1}=6>0$ which does not satisfy the proposed state-centric necessary condition.
+Fix the first 5 arcs as well as the ASL, and vary the motion time of the 6th arc to search the minimal terminal time until reaching the boundary of the feasible set.
+If some arcs reach zero time, then delete the system behavior from the ASL.
+Repeat the above process, resulting in trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$.
+As shown in Figs.
+6(a) and 7(b), the terminal time decreases through iteration, and the DOF finally reaches 0 after 5 iterative steps.
+Finally, the optimized trajectory ${\boldsymbol{x}}_{6}$ achieves a 17.7% reduction in the terminal time compared to the original trajectory ${\boldsymbol{x}}_{1}$.
+In fact, it remains challenging to plan near-optimal high-order trajectories using arc-based methods despite advantages in computational accuracy and analytical simplicity.
+The proposed state-centric condition has the potential to optimize BBS trajectories planned by arc-based methods, which will be investigated in future works.
+Figure 6: Iteration of a 5th-order feasible BBS trajectory. The initial trajectory ${\boldsymbol{x}}_{1}$ is planned by [36]. (a) The states and controls. (b) The ASLs.
+A discretized optimization method [37] is applied to plan the trajectory ${\boldsymbol{x}}_{\text{base}}$ for comparison, where the time domain is discretized into 1000 intervals.
+An optimal solution is obtained with a 0.016% relative error of terminal time compared to ${\boldsymbol{x}}_{6}$, as shown in Fig.
+7(b).
+Therefore, [37] can plan optimal or near-optimal trajectories for the high-order problem.
+However, $u_{\text{base}}$ exhibits oscillation, as shown in Fig.
+7(a1), while $u_{i}$ is BBS.
+Defining the total variation of the control as $T_{\mathrm{v}}\triangleq\int_{0}^{{t_{\mathrm{f}}}}\left|u\left(t\right)\right|\mathrm{d}t$ to describe the trajectory stability, ${\boldsymbol{x}}_{5}$ achieves an 8.9% $T_{\mathrm{v}}$ of ${\boldsymbol{x}}_{\text{base}}$.
+For discretized optimization like [37], the open-loop errors of terminal states are difficult to constrain due to numerical error in each interval.
+Define $E_{\mathrm{s}}=\sqrt{\sum_{k=1}^{n}\left(\frac{\hat{x}_{{\mathrm{f}}k}-x_{{\mathrm{f}}k}}{x_{\mathrm{m}k}}\right)^{2}}$ where $\hat{\boldsymbol{x}}_{\mathrm{f}}=\left(\hat{x}_{{\mathrm{f}}k}\right)_{k=1}^{n}$ is directly integrated from the control input.
+As shown in Fig.
+7(d), arc-based trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$, reduce at least 5 orders of magnitude of $E_{\mathrm{s}}$ compared to ${\boldsymbol{x}}_{\text{base}}$.
+Hence, an approach combining arc-based methods and the proposed state-centric necessary condition has the potential to plan near-optimal trajectories with higher trajectory stability and computational accuracy compared to discretized optimization.
+Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [36], the optimized trajectory ${\boldsymbol{x}}_{6}$ through the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [37].
+(a) The states and controls.
+(b) The terminal time.
+(c) The total variation of the control.
+(d) The open-loop error of terminal states.
+VIII Conclusions
+This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input.
+The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form.
+Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility.
+Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one.
+The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be full row rank.
+In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.
+The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrator systems with full box constraints from both theoretical and numerical perspectives.
+An upper bound of the number of arcs is determined in a general sense, and a recursive equation for the junction time in chattering induced by 2nd-order constraints is derived.
+Note that the above two conclusions are challenging to prove by costate-based conditions.
+Numerical experiments showed that the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information, and it has the potential to optimize feasible BBS trajectories with higher trajectory stability and computational accuracy compared to discretized optimization methods.
 Acknowledgment
-This work was supported by the National Key Research and Development Program of China under Grant 2023YFB4302003.
+This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB4302003, and in part by the National Natural Science Foundation of China under Grant 624B2077.
+Appendix A An Example of Application of the Proposed State-Centric Necessary Condition
+Let’s introduce Corollary 4 of order $n$ as follows.
+We will show why it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions, and how to obtain the result based on the proposed state-centric necessary condition.
+Specifically, we will try to derive the result based on PMP-based/state-centric necessary conditions as examples.
+A-A Problem Statement
+Consider the following simplified 4th-order chain-of-integrator system:
+$\min$ ${t_{\mathrm{f}}}$ (62)
+s.t. $\dot{x}_{1}=u,\,\dot{x}_{k}=x_{k-1},\,k\in\left\{2,3,\dots,n\right\},$
+$\left|u\right|\leq 1,x_{2}\leq 1,$
+${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}}.$
+Now, the question is: Is there a $\left({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{\mathrm{f}}\right)$, s.t.
+the optimal control is chattering?
+In other words, can the optimal control switch infinitely many times in a finite interval $\left[0,{t_{\mathrm{f}}}\right]$?
+(For the case where $n=4$, the answer is NO according to Corollary 4 in the manuscript.)
+A-B Preliminaries
+To compare fairly the proposed state-centric necessary condition and PMP-based necessary conditions, we provide some preliminaries for both approaches.
+Assume that chattering occurs in the optimal control. Then, $\exists\left\{t_{i}\right\}_{i\in{\mathbb{N}}}\subset\left[0,{t_{\mathrm{f}}}\right]$ increasing strictly monotonically, s.t.
+  • $\left\{t_{i}\right\}_{i\in{\mathbb{N}}}$ has a finite limit point, i.e.,
+    $t_{\infty}\triangleq\lim_{i\to\infty}t_{i}\in\left[0,{t_{\mathrm{f}}}\right].$ (63)
+    Furthermore, the chattering (unconstrained) arcs converge to a constrained arc at $t_{\infty}$, i.e.,
+    $\lim_{t\to t_{\infty}}x_{1}=0,\,\lim_{t\to t_{\infty}}x_{2}=1.$ (64)
+  • $\forall i\in{\mathbb{N}}$, $x_{2}\left(t_{i}\right)=1$ and $x_{1}\left(t_{i}\right)=0$ hold.
+  • $\forall i\in{\mathbb{N}}$, the control in $\left(t_{i},t_{i+1}\right)$ satisfies
+    $u\left(t\right)=\begin{dcases}-1,&t\in\left(t_{i},\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1}\right),\\ 1,&t\in\left(\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1},\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1}\right),\\ -1,&t\in\left(\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1},t_{i+1}\right).\end{dcases}$ (65)
+The proof of the above conclusions can be found in our related work [32] (Section IV-B).
+The above conclusions are visualized in Figure 8.
+Figure 8: Chattering phenomena in time-optimal control for 4th-order chain-of-integrator systems with constraints on $u$ and $x_{2}$.
+A-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
+Now, we try to derive the existence/non-existence of chattering based on PMP-based necessary conditions, which will be shown challenging.
+The Hamiltonian of problem (62) is
+${\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)=\lambda_{0}+\lambda_{1}u+\sum_{k=2}^{n}\lambda_{k}x_{k-1}+\eta\left(x_{2}-1\right).$ (66)
+The costate vector ${\boldsymbol{\lambda}}$ satisfies the Hamilton’s equations, i.e.,
+$\begin{dcases}\dot{\lambda}_{2}=-\lambda_{3}+\eta,\\ \dot{\lambda}_{n}=0,\\ \dot{\lambda}_{k}=-\lambda_{k+1},\,k\in\left\{1,3,4,5,\dots,n-1\right\}.\\ \end{dcases}$ (67)
+Furthermore, $\forall t\in\left[0,{t_{\mathrm{f}}}\right]$, it holds that
+$\begin{dcases}\lambda_{0}\geq 0,\,\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\not={\boldsymbol{0}},\\ \eta\left(t\right)\geq 0,\,\eta\left(t\right)\left(x_{2}\left(t\right)-1\right)=0,\\ {\mathcal{H}}\left(t\right)\equiv 0.\end{dcases}$ (68)
+PMP implies that it holds a.e. that
+$u\left(t\right)=\begin{dcases}-1,&\lambda_{1}\left(t\right)>0,\\ 0,&\lambda_{1}\left(t\right)\equiv 0\text{ holds for a period},\\ 1,&\lambda_{1}\left(t\right)<0.\\ \end{dcases}$ (69)
+Specifically, in a period, $u\equiv 0$ if and only if a constrained arc occurs where $x_{2}\equiv 1$.
+Furthermore, a junction of the costate vector ${\boldsymbol{\lambda}}$ occurs when the state constraint $x_{2}\leq 1$ switches between active and inactive.
+Specifically, at junction time $t_{i}$, $i\in{\mathbb{N}}$, it holds that
+$\begin{dcases}\lambda_{2}\left(t_{i}^{+}\right)\leq\lambda_{2}\left(t_{i}^{-}\right),\\ \lambda_{k}\left(t_{i}^{+}\right)=\lambda_{k}\left(t_{i}^{-}\right).\end{dcases}$ (70)
+Consider the trajectory in $\left[t_{0},t_{\infty}\right]$. Denote
+$\begin{dcases}\lambda_{0k}\triangleq\lambda_{k}\left(t_{0}^{+}\right),\,\forall k\in\left\{1,2,\dots,n\right\},\\ \mu_{i}=\lambda_{2}\left(t_{i}^{-}\right)-\lambda_{2}\left(t_{i}^{+}\right)\geq 0,\,\forall i\in{\mathbb{N}}^{*}.\end{dcases}$ (71)
+Then costates in $\left(t_{0},t_{\infty}\right)$ are
+$\begin{dcases}\lambda_{k}\left(t\right)=\sum_{j=k}^{n}\frac{\lambda_{0j}}{\left(j-k\right)!}\left(t_{0}-t\right)^{j-k},\,k\in\left\{3,4,\dots,n\right\},\\ \lambda_{2}\left(t\right)=\sum_{j=2}^{n}\frac{\lambda_{0j}}{\left(j-2\right)!}\left(t_{0}-t\right)^{j-2}-\sum_{i=1}^{\infty}\mu_{i}\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right),\\ \lambda_{1}\left(t\right)=\sum_{j=1}^{n}\frac{\lambda_{0j}}{\left(j-1\right)!}\left(t_{0}-t\right)^{j-1}-\sum_{i=1}^{\infty}\mu_{i}\left(t_{i}-t\right)\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right),\end{dcases}$ (72)
+where
+$\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right)=\begin{dcases}1,&t\in\left(t_{i},t_{\infty}\right),\\ 0,&\text{otherwise}.\end{dcases}$ (73)
+In other words, $\forall k\geq 3$, $\lambda_{k}\left(t\right)$ is a polynomial of order $\left(n-k-1\right)$.
+$\forall k\leq 2$, $\lambda_{k}\left(t\right)$ is a piecewise polynomial of order $\left(n-k-1\right)$ due to the junction condition (70).
+For convenience, let
+$\lambda_{1i}\left(t\right)=\sum_{j=1}^{n}\frac{\lambda_{0j}}{\left(j-1\right)!}\left(t_{0}-t\right)^{j-1}-\sum_{p=1}^{i-1}\mu_{p}\left(t_{p}-t\right),\,i\in{\mathbb{N}}^{*}.$ (74)
+Then,
+$\lambda_{1}\left(t\right)=\lambda_{1i}\left(t\right),\,\forall i\in{\mathbb{N}}^{*},\,t\in\left(t_{i-1},t_{i}\right).$ (75)
+According to (65) and (69), it holds that
+$\lambda_{1i}\left(\frac{1}{4}t_{i}+\frac{3}{4}t_{i-1}\right)=\lambda_{1i}\left(\frac{3}{4}t_{i}+\frac{1}{4}t_{i-1}\right)=0,\,\forall i\in{\mathbb{N}}.$ (76)
+In other words, $\forall i\in{\mathbb{N}}$, $\exists\hat{t}_{i}\not\in\left[t_{i-1},t_{i}\right]$, s.t.
+$\lambda_{1i}\left(t\right)=\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),$ (77)
+where $p_{i}$ is a polynomial of order no higher than $n-3$. Furthermore,
+$\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0.$ (78)
+Therefore, a sufficient condition for the non-existence of chattering is $\mathcal{C}=\varnothing$, where
+$\mathcal{C}=\left\{\begin{aligned} &\left(\left(\mu_{i}\right)_{i\in{\mathbb{N}}^{*}},\left(t_{i}\right)_{i\in{\mathbb{N}}},\left(p_{i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}},\left(\lambda_{1i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}}\right):\\ &\qquad t_{i}<t_{i+1},\,\lim_{i\to\infty}t_{i}<\infty,\,\lambda_{0n}\not=0,\,\mu_{i}\geq 0,\\ &\qquad p_{i}\left(\bullet\right)\text{ is a polynomial of order no higher than }n-3,\\ &\qquad\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0,\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1i}\left(t\right)=\lambda_{0n}\frac{\left(-1\right)^{n-1}}{\left(n-1\right)!}\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1\,i+1}\left(t\right)=\lambda_{1i}\left(t\right)+\mu_{i}\left(t-t_{i}\right)\end{aligned}\right\}.$ (79)
+It is significantly challenging to prove that $\mathcal{C}_{\pm}=\varnothing$.
+A key step is deriving a recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$.
+However, the above key step requires eliminating the polynomial $p_{i}$, which involves more complex computations.
+As far as we have tried, it is too complex to obtain the recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$ based on (79).
+In contrast, our proposed Corollary 4 can directly derive the recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$ in a compact and simple form.
+A-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
+Denote
+$\tau_{i}=t_{\infty}-t_{i}.$ (80)
+Then, chattering is allowed only if $\tau_{0}\in{\mathbb{R}}_{++}$ and $\lim_{i\to\infty}\tau_{i}=0$.
+According to Corollary 4, the recursive expression for $\left(\tau_{i}\right)_{i\in{\mathbb{N}}^{*}}$ is
+$\det\left(\left(\tau_{i-k}+3\tau_{i-k-1}\right)^{j+1}-3\left(3\tau_{i-k}+\tau_{i-k-1}\right)^{j+1}+3\left(\tau_{i-k+1}+3\tau_{i-k}\right)^{j+1}-\left(3\tau_{i-k+1}+\tau_{i-k}\right)^{j+1}\right)_{j\in\left[n-2\right],k\in\left[n-2\right]}=0.$ (81)
+Once the initial value $\left(\tau_{i}\right)_{i=1}^{n-2}$ is given, the whole sequence $\left(\tau_{i}\right)_{i\in\infty}$ is determined.
+In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}=\varnothing$, where
+[…]
+Evidently, the condition on (82) is much simpler than the condition on (79).
+Specifically, the condition on (82) requires analyzing the convergence of a series for the recursive expressions (81) in the form of a system of polynomial equations.
+A-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
+Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}=\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (62).
+Note that the above claim is a novel theoretical result.
+Let $n=4$. Then, the recursive expression (81) for $\left(\tau_{i}\right)_{i\in{\mathbb{N}}^{*}}$ is
+$\det\left(\left(\tau_{i-k}+3\tau_{i-k-1}\right)^{j+1}-3\left(3\tau_{i-k}+\tau_{i-k-1}\right)^{j+1}+3\left(\tau_{i-k+1}+3\tau_{i-k}\right)^{j+1}-\left(3\tau_{i-k+1}+\tau_{i-k}\right)^{j+1}\right)_{j\in\left[2\right],k\in\left[2\right]}$ (83)
+$=$ $144\left(\tau_{i-1}-\tau_{i-3}\right)\left(\tau_{i}-\tau_{i-2}\right)f\left(\tau_{i-1}-\tau_{i};\tau_{i-2}-\tau_{i-1},\tau_{i-3}-\tau_{i-2}\right)=0,$
+where
+$f\left(z;z_{1},z_{2}\right)=\left(z_{1}-z_{2}\right)z^{2}+\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right)z-z_{2}\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right).$ (84)
+Note that $\tau_{i}<\tau_{i-1}<\tau_{i-2}<\tau_{i-3}$. Let $z_{i}=\tau_{i-1}-\tau_{i}>0$. Then,
+$f\left(z_{i};z_{i-1},z_{i-2}\right)=0.$ (85)
+$\lim_{i\to\infty}\tau_{i}=0$ implies that $\lim_{i\to\infty}z_{i}=0$. Without loss of generality, let $z_{1}>z_{2}$. Then,
+$\begin{dcases}f\left(z_{2};z_{1},z_{2}\right)=\left(z_{1}-z_{2}\right)^{2}z_{2}>0,\\ f\left(0;z_{1},z_{2}\right)=-z_{2}\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right)<0.\end{dcases}$ (86)
+Note that $f\left(z;z_{1},z_{2}\right)$ is a quadratic function of $z$.
+Therefore, $f\left(z;z_{1},z_{2}\right)=0$ has two roots in $\left(-\infty,0\right)$ and $\left(0,z_{2}\right)$, respectively.
+Since $z_{3}>0$ and $f\left(z_{3};z_{1},z_{2}\right)=0$, it holds that $0<z_{3}<z_{2}<z_{1}$.
+For the same reason recursively, it holds that
+$0<z_{i}<z_{i-1},\,\forall i\geq 2$ (87)
+Let $r_{i}=1-\frac{z_{i+1}}{z_{i}}\in\left(0,1\right)$. Then, (85) implies that
+$g\left(r_{i+1},r_{i}\right)\triangleq r_{i+1}^{2}-\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)r_{i+1}+1=0.$ (88)
+Note that
+$\begin{dcases}g\left(0,r_{i}\right)=1>0,\\ g\left(r_{i},r_{i}\right)=-\frac{r_{i}\left(2-r_{i}\right)^{2}}{1-r_{i}}<0,\\ g\left(1,r_{i}\right)=-1-\frac{1}{r_{i}\left(1-r_{i}\right)}<0.\end{dcases}$ (89)
+(88) has two roots in $\left(0,r_{i}\right)$ and $\left(1,+\infty\right)$, respectively.
+Hence, $10<r_{i+1}<r_{i}<$ 1.
+Denote $r_{\infty}\triangleq\lim_{i\to\infty}r_{i}\in\left[0,1\right]$.
+Then, $g\left(r_{\infty},r_{\infty}\right)$ implies that $r_{\infty}=0$.
+Note that
+$r_{i+1}=\frac{1}{2}\left[-\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)+\sqrt{\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)^{2}-4}\right]=r_{i}-4r_{i}^{2}+\mathcal{O}\left(r_{i}^{3}\right),\,i\to\infty.$ (90)
+By Stolz-Cesàro theorem, it holds that
+$\lim_{i\to\infty}i\left(\frac{z_{i}}{z_{i+1}}-1\right)=\lim_{i\to\infty}\frac{ir_{i}}{1-r_{i}}=\lim_{i\to\infty}\frac{i}{\frac{1}{r_{i}}}=\lim_{i\to\infty}\frac{i+1-i}{\frac{1}{r_{i+1}}-\frac{1}{r_{i}}}=\lim_{i\to\infty}\frac{r_{i}^{2}\left(1+\mathcal{O}\left(r_{i}\right)\right)}{4r_{i}^{2}+\mathcal{O}\left(r_{i}^{3}\right)}=\frac{1}{4}\in\left(0,1\right).$ (91)
+By Raabe-Duhamel’s test, it holds that
+$\infty=\sum_{i=1}^{\infty}z_{i}=\tau_{0}-\tau_{\infty}=t_{\infty}-t_{0}<\infty,$ (92)
+which is a contradiction.
+Therefore, the set $\widehat{\mathcal{C}}$ defined in (82) is empty. Therefore, chattering does not exist in the optimal control of problem (62) when $n=4$.
+A-F Summary
+In this appendix, we provide an example to compare the proposed state-centric necessary condition and PMP-based necessary conditions.
+  1. 1.
+    We show that it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions in the considered optimal control problem.
+The key step is to derive the recursive expression for junction times, which involves complex computations.
+  2. 2.
+    We show that the proposed state-centric necessary condition can directly derive the recursive expression for junction times in a compact and simple form, which is much simpler than the condition based on PMP-based necessary conditions.
+Specifically, we prove that chattering does not exist in the considered optimal control problem of order 4.

### v3

@@
 Given a feasible trajectory, the perturbed trajectory under the constraints of ASL is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information.
-A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures.
+A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not of full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures.
 The proposed necessary condition is applied to high-order chain-of-integrator systems with full box constraints, contributing to some theoretical results challenging to reason by costate-based conditions.
@@
 However, the above necessary conditions require information on both states and costates, while costates are usually unobservable in practice, resulting in a challenge in determining the optimality of a given feasible trajectory by costate-based conditions.
+Figure 1: Graphical roadmap of the proposed theoretical framework. Main contributions regarding representations and optimal conditions are highlighted in red.
 Based on PMP, one can get rid of dependence on costates by deriving the optimal trajectory’s structure.
@@
 Walther et al.
-[17] computated switching surfaces for some systems through Gröebner basis.
+[17] computed switching surfaces for some systems using a Gröebner basis.
 Patil et al.
@@
 Another type of necessary condition leverages Bellman’s principle of optimality [22], asserting that a sub-arc of an optimal trajectory is itself optimal for the induced sub-problem, without explicitly involving costates.
-The Hamilton-Jacobi-Bellman (HJB) equation [23] represents a typical necessary condition in the form of partial differential equations, which is widely used in optimal control [24, 25].
+With wide applications in optimal control [23, 24], the Hamilton-Jacobi-Bellman (HJB) equation [25] serves as a fundamental necessary condition which is also sufficient under suitable regularity conditions.
 Evans and James [26] investigated the HJB equation for time-optimal control.
 Wolenski and Zhuang [27] showed that the minimal time function is the unique proximal solution to the HJB equation.
-However, most numerical HJB-based methods require high computational cost [28] due to the discretized grids.
+However, most numerical HJB-based methods require high computational cost [28].
 From a theoretical perspective, HJB-based methods also fail to preserve arc structures, which hinders the analytical reasoning.
-Compared with the PMP-based and HJB-based necessary conditions, this paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information.
+This paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information.
+A graphical roadmap is shown in Fig.
+1.
 Sections III and IV derive trajectories’ structures on arcs and keypoints for feasibility, respectively.
-Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively.
+Guided by the knowledge of trajectory structures, Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively.
+Proofs of propositions and theorems are provided in Appendix A.
 The contributions of this paper are as follows:
@@
   2. 2.
-    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not full row rank.
+    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not of full row rank.
 The condition is derived from the fact that if a BBS trajectory is optimal, then all perturbed feasible trajectories have a strictly longer terminal time than the original one.
@@
     This paper proposes a potential optimization idea based on the developed necessary condition.
-Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not full row rank.
+Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not of full row rank.
 During iteration, the feasibility is guaranteed based on the developed ASL framework.
@@
   4. 4.
-    The proposed state-centric necessary condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain.
+    The proposed state-centric condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain.
 An upper bound on the number of arcs is determined in a general sense, while case-by-case analysis is necessary in traditional methods.
-A recursive equation of junction time is reasoned in chattering induced by 2nd-order constraints, which is challenging to prove by costate analysis.
+A recursive equation for the junction time when the state is tangent to the constraints’ boundaries is derived in chattering arcs induced by 2nd-order constraints, which is challenging to prove by costate analysis.
 As a result, one could determine the existence of chattering induced by 2nd-order constraints in chain-of-integrator systems in future works.
@@
 This section formulates the time-optimal control problem for controllable linear systems and introduces some necessary assumptions. The optimal control problem is as follows:
-$\min_{u}\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t={t_{\mathrm{f}}},$ (1a)
+$\min_{u}\quad$ ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (1a)
 ${\mathrm{s.t.}}\quad$ $\dot{{\boldsymbol{x}}}\left(t\right)={\boldsymbol{A}}x\left(t\right)+{\boldsymbol{b}}u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (1b)
@@
 (7) is the well-known bang-bang and singular (BBS) control law [11, 29].
-In other words, the optimal control is always ±$\pm{u_{\mathrm{m}}}$, except during singular arcs.
-Note that the objective function $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
+In other words, the optimal control is always $\pm{u_{\mathrm{m}}}$, except during singular arcs.
+Note that the objective function ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
 $\forall t\in\left[0,t_{\mathrm{f}}\right],\,{\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)\equiv 0.$ (8)
@@
 An arc of a trajectory is defined as a continuous segment conditioned by a fixed dynamic equation, i.e., $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ where $\widehat{{\boldsymbol{A}}}$ and $\widehat{{\boldsymbol{b}}}$ are constant.
-Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A and Section III-B, respectively.
-Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C.
-Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
-III-A Unconstrained Arcs in Problem (1a)
+Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A.
+Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-B.
+Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-C.
+III-A Unconstrained and Constrained Arcs in Problem (1a)
 In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds almost everywhere (a.e.), and the control does not switch.
-The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
+Proposition 1 states that $u\equiv\pm{u_{\mathrm{m}}}$ holds in an unconstrained arc.
 Proposition 1.
@@
 that ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not=0$ and $u\equiv-{u_{\mathrm{m}}}{\mathrm{sgn}}\left({\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\right)$.
-Proof.
-Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period.
-By (5), ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ implies that ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ holds a.e.
-in an unconstrained arc.
-Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$.
-Therefore,
-$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}=\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\left({\boldsymbol{\lambda}}^{\top}{\boldsymbol{b}}\right)\equiv 0,$ (9)
-i.e., ${\boldsymbol{\lambda}}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\equiv{\boldsymbol{0}}$.
-However, by (2), ${\boldsymbol{\lambda}}\equiv 0$ holds a.e.
-Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$.
-Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$ holds a.e.
-∎
 Remark.
-By $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$, either ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period or ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 for finite times, while the former case is ruled out in an unconstrained arc by Proposition 1.
-The control switches between $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$ when ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0, where two unconstrained arcs are connected.
-III-B Constrained Arcs in Problem (1a)
+In an unconstrained arc, $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$ holds since ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$.
+Two cases could occur: (a) ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ holds for a period; (b) ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 for finite times.
+Case (a) is ruled out in an unconstrained arc by Proposition 1.
+In Case (b), the control switches between $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$ when ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0, where two unconstrained arcs are connected.
 Compared to unconstrained arcs, constrained arcs exhibit more complex behavior.
@@
 ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$.
-The following proposition characterizes a constrained arc.
+Proposition 2 characterizes a constrained arc.
+Proposition 3 points out that multiple inequality constraints are allowed to be active in a constrained arc.
 Proposition 2.
@@
 ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (10), then under the driving of (9), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in $\left[t_{1},t_{2}\right]$.
-Proof.
-(2) implies ${\boldsymbol{c}}_{p}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\not={\boldsymbol{0}}$ since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$; hence, $r_{p}$ exists and serves as the minimum non-zero column index of a non-vzero matrix.
-For $r\in\left[r_{p}-1\right]$, considering the $r$th-order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$.
-Hence, (10) holds.
-Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
-${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (11).
-Under the driving of (10), note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}=0$; hence, (11) holds in $\left[t_{1},t_{2}\right]$.
-∎
-It is allowed that multiple inequality constraints are active in a constrained arc, as proved in Proposition 3.
 Lemma 1.
@@
 Furthermore, if ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$ satisfies $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on $\mathbb{R}$}.
-Proof.
-The proof of (12) is omitted since it is elementary [31].
-Assume that $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
-Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t.
-$p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$.
-Therefore, $\forall r\geq n$, $\exists p_{r}^{*}$ s.t.
-$p_{r}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{r}^{*}\left({\boldsymbol{A}}\right)$.
-Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$.
-Since $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, Note that $\forall r\geq n+1$,
-$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$ (13)
-It can be proved that ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)$ is analytic. By (13), $\forall t\in{\mathbb{R}}$,
-${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right).$ (14)
-In other words, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$. ∎
 Proposition 3.
-[…]
+Consider a constrained arc for $t\in\left[t_{1},t_{2}\right]$.
+$\mathcal{P}\subset\left[P\right]$.
+Assume that during the arc, $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds.
+Then, $\forall p\in\mathcal{P}$, $t\in\left[t_{1},t_{2}\right]$, it holds that
 $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0,\\ \forall q\in\mathcal{P},\,r\in\left[n\right],\,{\boldsymbol{c}}_{q}^{\top}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right)^{r}{\boldsymbol{x}}=0.\end{dcases}$ (12)
@@
 Then, under the driving of (9) with an arbitrary $p\in\mathcal{P}$, it holds that $\forall q\in\mathcal{P}$, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}+d_{q}\equiv 0$ in $\left[t_{1},t_{2}\right]$.
-Proof.
-Assume that $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in the constrianed arc.
-By Proposition 2, (10) holds $\forall p\in\mathcal{P}$.
-$\forall q\in\mathcal{P}$, since ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv{\mathrm{const}}$, (15) can be derived through considering the $r$th-order derivative of ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}$ for $r\in\left[n\right]$.
-Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
-${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (15).
-$\exists p\in\mathcal{P}$, (15) holds in $\left[t_{1},t_{2}\right]$.
-Then, $\forall q\in\mathcal{P}$, $r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds.
-By Lemma 1, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv-d_{q}$ holds.
-∎
 Example 1.
-Consider a chain-of-integrator system of order $n$, i.e.,
-[…]
+Consider a chain-of-integrator system of order $n$:
+$\begin{dcases}\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right],\,1<k\leq n,\\ \dot{x}_{1}\left(t\right)=u\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right].\end{dcases}$ (13)
 Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$.
-$x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold.
+$x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained arc is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold.
 In this case, $x_{2}\equiv 0$ implies $x_{1}\equiv 0$ and $u\equiv 0$; hence, $x_{1}+x_{3}\leq 1$ implies $x_{3}\equiv 1$.
 Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are inconsistent, i.e., in a constrained arc, if $x_{1}+x_{3}\equiv 1$ holds, then $x_{2}\equiv 1$ cannot hold.
-III-C System Behavior
+III-B System Behavior
 Definition 1.
@@
 ${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right)+u_{0}\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (14)
-For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not=\varnothing$.
+For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not$$\varnothing$.
 (9) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}$, where $\widehat{{\boldsymbol{a}}}_{p}$ is defined in (9).
@@
 $\widehat{{\boldsymbol{a}}}_{p}\not=\widehat{{\boldsymbol{a}}}_{q}$, it is guaranteed by the consistency of $\mathcal{P}$ that $\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{a}}}_{q}^{\top}{\boldsymbol{x}}$ during the arc.
-For convenience, let $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$.
+Denote by $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$.
 Then, $\mathcal{S}$} in Definition 1 is well-defined.
-By Lemma 1, during ${\mathcal{S}}$,
+By Lemma 1, it holds during $\mathcal{S}$} that
 ${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{\widehat{{\boldsymbol{A}}}\left(t-t_{0}\right)}{\boldsymbol{x}}\left(t_{0}\right).$ (15)
 Example 2.
-Consider a chain-of-integrator system (16) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
-The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
+Consider a chain-of-integrator system (13) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
+The $\left(2k-1\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$.
 A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$th-order, where $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$, and $u\equiv 0$.
@@
 The trajectory is tangent to $\left\{x_{3}=0\right\}$ at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, and $\mathcal{T}_{2}^{(2)}$.
-III-D Switching Law
-Based on the analysis in Section III-C, if the optimal trajectory consists of a finite number of arcs, then the optimal trajectory can be represented by a sequence of system behaviors.
+III-C Switching Law
+Based on the analysis in Section III-B, if the optimal trajectory consists of a finite number of arcs, then the optimal trajectory can be represented by a sequence of system behaviors.
 The switching law is defined as follows:
@@
 For the example in Fig.
-1, the switching law is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}$.
+2, the switching law is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}$.
 ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{3}$ are unconstrained arcs where ${\boldsymbol{x}}>{\boldsymbol{0}}$ holds a.e.
@@
 If ${\boldsymbol{x}}\left(t_{1}\right)$ satisfies (21) and (10) for $p\in\mathcal{P}_{2}$, then $\exists\varepsilon>0$, ${\mathcal{S}}_{1}$ occurs for $\left(t_{1}-\varepsilon,t_{1}\right)$ and ${\mathcal{S}}_{2}$ occurs for $\left(t_{1},t_{1}+\varepsilon\right)$.
-Proof.
+Remark.
+Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection.
+Note that the equality constraints at the connection time, i.e., (10), (17), (19), and (21), are allowed to be linearly dependent.
+When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
+IV Keypoints for Trajectory Feasibility
+The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, while the feasibility during an arc remains inadequately investigated.
+To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of keypoints in an arc.
+As formally defined in Definition 5, a keypoint, where the trajectory touches state constraints’ boundaries in isolation, directly influences the feasibility of the perturbed trajectory.
+For the example shown in Fig.
+2, the time points at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, $\mathcal{T}_{2}^{(2)}$, and ends of arcs are keypoints.
+Section IV-A analyzes additional constraints of an arc’s end besides Proposition 4.
+The tangent marker is proposed in Section IV-B to describe the feasibility of a BBS trajectory tangent to constraints’ boundaries.
+Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
+IV-A Additional Constraints at the End of An Arc
+Proposition 4 fully discusses the feasibility of the connection of two adjacent arcs w.r.t.
+constraints induced by $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$.
+However, the constraints on $u$ and $\boldsymbol{x}$}, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ still lacks discussion.
+The following proposition provides a sufficient and necessary condition for the feasibility at the end of an arc.
+Proposition 5.
+Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
+  1. 1.
+    $\forall p\not\in\mathcal{P}$, $\exists\varepsilon>0$, s.t.
+${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\leq 0$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds:
+    1. (a)
+      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}<0$.
+    2. (b)
+      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$.
+Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
+$\forall r\in\left[\hat{r}_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
+    Similar conclusions hold for $\left[t_{1}-\varepsilon,t_{1}\right]$.
+  2. 2.
+    Assume that $\mathcal{S}$} is a constrained arc.
+Then, $\exists\varepsilon>0$, s.t.
+$u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds, where $p\in\mathcal{P}$:
+    1. (a)
+      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<{u_{\mathrm{m}}}$.
+    2. (b)
+      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
+Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
+$\forall r\in\left[\hat{r}_{p}-1\right]$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
+    3. (c)
+      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
+Furthermore, $\forall r\in\left[n\right]$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
+    Similar conclusions hold for $u\geq-{u_{\mathrm{m}}}$ and $\left[t_{1}-\varepsilon,t_{1}\right]$.
+Consider two connected arcs ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
+Proposition 4 guarantees the feasibility of constraints active along ${\mathcal{S}}_{1}$ or ${\mathcal{S}}_{2}$.
+If some constraints are active at the ends of arcs but are inactive in ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ a.e., then additional constraints at the connection, as defined in Definition 3, should be introduced for feasibility based on Proposition 5.
+Definition 3.
+[…]
+Remark.
+Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 5 is directly determined.
+In Definition 3, the additional end-constraint $\mathcal{E}$} does not include the equality constraints induced by $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
+To avoid over-determination, the equality constraints in $\mathcal{E}$} and $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ are made linearly independent through elimination.
+The elimination process is guaranteed by the feasibility of the given trajectory.
+${\mathcal{L}}=\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(1)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(2)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{2}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{(N)}\right\}_{k=1}^{M_{N}}\right)$ (22)
+IV-B Tangent Markers in An Arc
+Proposition 4 and Proposition 5 analyzed the feasibility of an arc near the end points.
+An arc might be tangent to some constrained boundary, where small perturbations may affect the feasibility of the arc.
+The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to constraints’ boundaries.
+Proposition 6.
+Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $\left[t_{0},t_{2}\right]$.
+$\mathcal{S}$} is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists\hat{r}\in\left[n\right]$ is even, s.t.
+$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall r\in\left[\hat{r}-1\right],\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (23)
+Regardless of whether an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained, $\mathcal{S}$} can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$.
+It is noteworthy that $\mathcal{S}$} can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $u\leq{u_{\mathrm{m}}}$.
+For this case, $\mathcal{S}$} should be constrained.
+Arbitrarily consider $p\in\mathcal{P}$.
+According to Proposition 2, $\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$.
+In other words, the constraints on $u$ are equivalent to the constraints on $\boldsymbol{x}$}.
+Hence, Proposition 6 can apply for both state constraints and control constraints.
+For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$.
+The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $u\leq{u_{\mathrm{m}}}$, respectively.
+Another noteworthy fact is that an arc $\mathcal{S}$} can be tangent to multiple constraints $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$ at a time point $t_{1}$.
+Then, the necessary and sufficient condition for feasibility is $\forall p\in\mathcal{P}^{\prime}$, (23) holds.
+The tangent marker is defined as follows:
+Definition 4.
+Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $t\in\left[t_{0},t_{2}\right]$.
+Assume that at $t_{1}\in\left(t_{0},t_{2}\right)$, the arc $\mathcal{S}$} is tangent to some constraints’ boundaries whose indices constitute a nonempty set $\mathcal{P}^{\prime}\subset\left[P+2\right]$.
+By Proposition 6, the tangent condition (23) is equivalent to $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right$] has full row rank.
+Then, the tangent marker is defined as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$.
+Remark.
+In Definition 4, the equality constraints induced by (23) can be linearly dependent.
+However, for a given feasible trajectory, the equalities should have a feasible solution; hence, equality constraints with full row rank, i.e., $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, can be constructed through elimination.
+Therefore, the tangent marker is well-defined.
+IV-C Augmented Switching Law
+Sections III, IV-A, and IV-B analyze the feasibility of the optimal trajectory at the end points of arcs and the tangent points to the constrained boundaries.
+To characterize the optimal trajectory, this section proposes the augmented switching law, which represents the input control and the feasibility of the trajectory in a compact form.
+Definition 5.
+Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$.
+During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$.
+At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$.
+Then, time points at tangent markers and ends of arcs are called keypoints for trajectory feasibility.
+The augmented switching law (ASL) $\mathcal{L}$} is defined in (22).
+Remark.
+If a feasible trajectory with a finite number of arcs follows the BBS control law (7), then Propositions 4, 5, and 6 hold for the trajectory.
+For this case, the ASL can also represent the feasible trajectory.
+Remark.
+If a feasible non-chattering BBS trajectory is perturbed by a sufficiently small perturbation, then the perturbed trajectory can only exceed constraints’ boundaries at the keypoints since the distance between the original trajectory and constraints’ boundaries is 0 at the keypoints, while it is strictly positive at other points.
+Therefore, the ASL can not only characterize the input control but also describe the feasibility of the trajectory.
+Remark.
+The switching law in Definition 2 corresponding to the ASL in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
+It can be observed that the ASL provides more detailed information on the feasibility of the trajectory at keypoints than the switching law.
+Notably, neither the switching law nor the ASL provides the motion time of each arc.
+To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as $\mathcal{S}$} in an ASL.
+If $\mathcal{E}$} is empty, then $\mathcal{E}$} is omitted.
+The additional end-constraints before ${\mathcal{S}}_{1}$ and after ${\mathcal{S}}_{N}$ are omitted since the initial and terminal states are given, i.e., ${\boldsymbol{x}}_{0}$ and ${\boldsymbol{x}}_{\mathrm{f}}$.
+For the example shown in Fig.
+2, the ASL is $\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{1}^{(1)}\right\}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{2},\left\{{\mathcal{T}}_{2}^{(1)},{\mathcal{T}}_{2}^{(2)}\right\}\right){\mathcal{S}}_{3}$.
+For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 31].
+The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively.
+The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively.
+Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$.
+The tangent marker where $\boldsymbol{x}$} is tangent to $x_{{\mathrm{m}}k}$ and $x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively.
+The additional end-constraint ${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $x_{j}\leq x_{{\mathrm{m}}j}$, respectively.
+Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc.
+Based on the above simplified notations, an example is provided to illustrate the ASL.
+Example 3.
+Consider a position-to-position problem for 4th-order chain-of-integrator systems with full box constraints shown in Fig.
+3.
+A feasible suboptimal trajectory with a BBS form, planned by the manifold-intercepted method (MIM) [20], is shown in Fig.
+3(a).
+It can be observed that the ASL is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
+Figure 3: Two examples of ASLs.
+For a 4th-order chain-of-integrator system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$.
+(a) A non-chattering feasible trajectory planned by [20].
+(b) A chattering optimal trajectory planned by [31].
+(b2) is the plot of $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
+Example 4.
+Our previous work [31] pointed out that the chattering phenomenon occurs in the optimal trajectory for the problem in Example 3.
+The chattering optimal trajectory is shown in Fig.
+3(b1).
+Fig.
+3(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied.
+It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$.
+According to Bellman’s principle of optimality [22], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is sufficiently small.
+The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
+V A State-Centric Necessary Condition
+The theoretical framework for the ASL of time-optimal control for linear systems is established in Section IV.
+Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a).
+Section V-A discusses the uniqueness of the time-optimal control.
+Then, given a feasible BBS trajectory, Section V-B analyzes the feasibility of the perturbed BBS trajectory.
+Based on the conclusions in Sections V-A and V-B, Section V-C provides a necessary condition of optimality.
+V-A Uniqueness of Time-Optimal Control
+The uniqueness of the time-optimal control for problem (1a) is proved in Theorem 1 as preliminaries.
+Theorem 1.
+Under Assumption 1, the time-optimal control of problem (1a) is unique in an a.e.
+sense.
+In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal, then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e.; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
+Remark.
+Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an a.e.
+sense when there exists at most one chattering limit point.
+The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
+According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly longer terminal time than the optimal terminal time.
+The above observation will be applied in Section V-B.
+V-B Feasibility of the Perturbed BBS Trajectory
+Consider a feasible BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ of problem (1a) without chattering.
+The ASL is denoted as $\mathcal{L}$}.
+For convenience of discussion, the keypoints of $\mathcal{L}$} occur at $t_{i}$, where $0=t_{0}<t_{1}<\dots<t_{M}=t_{\mathrm{f}}$.
+In other words, $\left\{t_{i}\right\}_{i=0}^{M}$ consists of time points at ends of arcs and tangent markers.
+Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$, and $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$.
+This section discusses the feasibility of a perturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ represented by the ASL $\mathcal{L}$$^{\prime}$, where the time vector ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ is close to $\boldsymbol{t}$}, i.e., $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}\right\|$ is sufficiently small.
+For convenience, the notation $\left\|\bullet\right\|$ refers to $\left\|\bullet\right\|_{\infty}$ in this section if not specified.
+The feasibility of the perturbed trajectory is proved by two steps.
+Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory.
+In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory.
+Secondly, Theorem 2 guarantees the feasibility of the perturbed trajectory based on Propositions 4, 5, 6, and 7.
+In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible.
+Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
+Proposition 7.
+Consider a BBS trajectory ${\boldsymbol{x}}\left(t\right)$ satisfying $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$.
+Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$.
+$\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
+$\forall\varepsilon\in\left(0,\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}\right)$, $\forall{\boldsymbol{x}}^{\prime}\left(t\right)$ is a perturbed trajectory induced by ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ with $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$,
+$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon$ (24)
+where $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$ and $t_{0}^{\prime}=t_{0}$.
+Theorem 2.
+Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is $\mathcal{L}$} in problem (1a).
+Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
+$\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is a trajectory also represented by $\mathcal{L}$$^{\prime}$$\mathcal{L}$}, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
+${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by $\boldsymbol{t}$} and $\boldsymbol{t}$$^{\prime}$ in the same way as that in Proposition 7.
+Theorem 2 points out that if the perturbed trajectory is represented by the same ASL to the original trajectory and the perturbance is small enough, then the perturbed trajectory is feasible.
+In other words, Theorem 2 provides a novel variational approach under the fixed ASL with a feasibility guarantee near constraints’ boundaries.
+V-C State-Centric Necessary Condition of Optimality
+Based on Theorem 2, this section provides an approach to constructing a feasible perturbed trajectory with the same ASL.
+According to Theorem 1, this section shows that the existence of such a perturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
+The Jacobian matrix of equality constraints induced by the ASL can be calculated through Proposition 8.
+Proposition 8.
+${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$, and ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ is fixed.
+Consider the mapping ${\boldsymbol{x}}_{i}:{\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}\mapsto{\boldsymbol{x}}\left(t_{i}\right)$.
+Then, $\forall i,j\in\left[M\right]$,
+$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}{\boldsymbol{P}}_{j,i}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\,j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ {\boldsymbol{0}},\,j>i,\end{dcases}$ (25)
+where ${\boldsymbol{P}}_{j,i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}$.
+Utilizing the notations in Section V-B, the ASL provides a system of equalities on every keypoints.
+Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank.
+For a constrained arc with active constraints $\mathcal{P}$, the equality constraints induced by $\mathcal{P}$ are added in only one end instead of both ends of the arcs since one end satisfies $\mathcal{P}$ implies that the whole arc satisfies $\mathcal{P}$.
+Note that the additional end-constraints at both ends are eliminated by equality constraints induced by $\mathcal{P}$.
+Similar processes are applied if $u\equiv{u_{\mathrm{m}}}$ or $u\equiv-{u_{\mathrm{m}}}$ during a constrained arc.
+${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$.
+Then, Theorem 3 provides a necessary condition of optimality.
+Theorem 3 (State-Centric Necessary Condition).
+Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1.
+Denote the equality constraints induced by the ASL as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$.
+Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
+Remark.
+In contrast to existing costate-based necessary conditions [9], Theorem 3 requires no information on costates, which significantly simplifies the computation.
+Compared to existing state-centric necessary conditions [23], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
+Some examples are provided in Sections VI-B and VI-C.
+Remark.
+[…]
+Remark.
+If a given non-chattering BBS trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical optimization of the trajectory based on Theorem 3.
+Numerical examples are provided in Section VII.
+VI Applications in Chain-of-Integrator Systems
+This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a).
+For convenience, the chain-of-integrator system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
+Time-optimal control for chain-of-integrator systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved.
+The problem can be summarized as follows:
+$\min\quad$ ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (26a)
+${\mathrm{s.t.}}\quad$ $\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),\,1<k\leq n,\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (26b)
+$\dot{x}_{1}\left(t\right)=u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (26c)
+${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}},$ (26d)
+$\left|x_{k}\left(t\right)\right|\leq x_{{\mathrm{m}}k},\,1\leq k\leq n,\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (26e)
+$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (26f)
+This section provides some applications of Theorem 3.
+Section VI-A proves a trivial conclusion from a state-centric perspective.
+Section VI-B and Section VI-C prove two corollaries that are challenging to prove by traditional costate-based necessary conditions.
+VI-A The Case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$
+The case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$ has a well-known conclusion that the optimal control switches no more than $\left(n-1\right)$ times [32].
+This section proves the above conclusion based on the proposed Theorem 3.
+Corollary 1.
+In problem (26a), assume that $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$. Then, the optimal control switches no more than $\left(n-1\right)$ times.
+Corollary 1 implies that the switching law is in the form of $\overline{0}\underline{0}\overline{0}\underline{0}\dots$ or $\underline{0}\overline{0}\underline{0}\overline{0}\dots$ with no more than $n$ arcs.
+The case in Section VI-A is trivial since no state constraints exist and the costate vector follows a fixed equation that $\forall k\in\left[n-1\right]$, $\dot{\lambda}_{i}=-\lambda_{i+1}$ and $\lambda_{n}\equiv{\mathrm{const}}$.
+However, when the state constraints are induced, the behavior of the optimal control can be complex.
+On the one hand, the multiplier $\boldsymbol{\eta}$} in (4) can be non-zero during a constrained arc.
+On the other hand, the junction of $\boldsymbol{\lambda}$} can occur when a state constraint switches between active and inactive [7].
+For this case, the costate analysis can be complex, while the proposed state-centric necessary condition provides a simple and effective approach for analyzing the optimal control.
+VI-B The Case where Only System Behaviors Occur
+The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to constraints’ boundaries, is widely applied in existing works on trajectory planning [19, 33, 20].
+Denote the ASL of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$.
+$\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$.
+The non-existence of tangent markers and additional end-constraints in $\mathcal{L}$} means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility.
+An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
+Corollary 2.
+Assume that the ASL consists of system behaviors without tangent markers and additional end-constraints except at $t_{\mathrm{f}}$}. Assume that:
+  1. 1.
+    $\forall 1\leq j<i$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=j+1}^{i}\left|{\mathcal{S}}_{k}\right|<i-j$.
+  2. 2.
+    $\forall i<j\leq N$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{j-1}\left|{\mathcal{S}}_{k}\right|<j-i$.
+  3. 3.
+    $\forall i\in\left[N\right]$, if $\left|{\mathcal{S}}_{i}\right|>0$ and $\forall i<j\leq N$, $\left|{\mathcal{S}}_{j}\right|<\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{N}\left|{\mathcal{S}}_{k}\right|\leq N-i$.
+Then, it holds that
+$N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|\leq n.$ (27)
+Remark.
+In our previous work [20], a similar conclusion pointed out that when fixing ${\boldsymbol{x}}_{\mathrm{f}}$}, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the ASL ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$ locally forms a submanifold of dimension $r$.
+Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$.
+However, the local optimality is not discussed in [20].
+Corollary 2 proves that the ASL with $r>n$ fails to achieve optimality.
+It is challenging to reason Corollary 2 by existing costate-based necessary conditions [9] due to complex behaviors of the costates.
+In contrast to Corollary 2, PMP-based approaches need to derive the costates for a given trajectory case-by-case.
+For example, in a 4th-order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$.
+$\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can be a candidate for optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$.
+Though Corollary 2 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing optimal control of problem (26a).
+Numerical examples are provided in Section VII.
+VI-C The Chattering Phenomenon induced by $\left|x_{2}\right|\leq x_{{\mathrm{m}}2}$
+Our previous work [31] points out that if the chattering phenomenon occurs in problem (26a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon.
+An infinite number of unconstrained arcs are joined by constraints’ boundaries, i.e., the unconstrained arcs are tangent to the boundary for infinite times.
+In contrast, constrained arcs do not occur.
+This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition.
+[31] provides the following lemma.
+Lemma 2.
+[31] Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (26a) in a left neighborhood of $t_{\infty}$, where $t_{\infty}$ is the chattering limit point.
+Then, $\exists\left\{t_{3k}\right\}_{k=0}^{\infty}\subset\left[t_{0},t_{\infty}\right]$ increasing strictly monotonically and converging to $t_{\infty}$, s.t.
+$x_{2}$ is tangent to $x_{{\mathrm{m}}2}$ at $t_{3k}$.
+$\forall k\in{\mathbb{N}}^{*}$, $u$ switches at most 2 times during $\left(t_{3k-3},t_{3k}\right)$.
+Corollary 3.
+Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (1a) in a left neighborhood of $t_{\infty}$.
+Let $\tau_{k}=t_{\infty}-t_{3k}$, $\forall k\in{\mathbb{N}}$.
+Denote $f_{m}\left(a,b,c\right)$ as
+$\left(b+3a\right)^{m}-3\left(3b+a\right)^{m}+3\left(c+3b\right)^{m}-\left(3c+b\right)^{m}.$ (28)
+Then, $\forall N\in{\mathbb{N}}^{*}$, $\boldsymbol{J}$$^{\prime}$ does not have full row rank, where
+${\boldsymbol{J}}^{\prime}=\left(f_{i+1}\left(\tau_{j-1},\tau_{j},\tau_{j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[N\right]}.$ (29)
+Furthermore, $\forall k\geq n-2$, $\tau_{k}$ can be calculated by the following recursive equation:
+$\det\left(f_{i+1}\left(\tau_{k-j-1},\tau_{k-j},\tau_{k-j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[n-2\right]}=0.$ (30)
+Remark.
+According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ under the recursive equation (30).
+It is evident that chattering does not occur when $n=3$ since $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$ converges to $\infty$.
+Through more refined derivation, it can be rigorously proved that $n=4$ does not hold.
+In future works, one can try to determine the existence of chattering in problem (26a) of order $n\geq 5$ under constraints on $x_{2}$, where Corollary 3 and Theorem 3 would serve as useful mathematical tools.
+However, it is challenging to obtain the recursive equation (30) for problem (26a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering.
+A comparison between state-centric and PMP-based approaches is provided in Appendix B.
+Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 3 provides an example to apply the proposed necessary condition to a chattering trajectory, where a sub-arc without chattering of the optimal trajectory is investigated.
+VII Numerical Experiments
+Numerical experiments for high-order chain-of-integrator systems (26a) are conducted to provide a potential way to apply the proposed state-centric necessary condition to optimizing trajectories.
+Since this paper focuses on the theoretical necessary condition instead of optimization algorithms, metrics like computational time are not considered.
+Note that this paper does not provide a manual optimization algorithm.
+VII-A Representing a Feasible BBS Trajectory
+Consider problem (1a) whose system matrix is non-diagonalizable with multiple Jordan blocks as follows:
+${\boldsymbol{A}}=\left[\begin{array}[]{cccc}-1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right],\,{\boldsymbol{b}}=\left[\begin{array}[]{c}0\\ 0\\ 2\\ -1\end{array}\right].$ (31)
+The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$.
+As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint.
+As shown in Figs.
+4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper.
+The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$.
+Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$.
+${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$.
+${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$.
+${\mathcal{T}}_{5}^{(1)}$ is a tangent marker where $x_{1}$ is tangent to 0.7-0.7.
+${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$.
+In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$.
+In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$.
+One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
+A numerical optimal control toolbox, i.e., Yop [34], is applied to solving problem (31).
+The number of intervals is set to 200.
+The resulting trajectory without a specified initial solution ${\boldsymbol{x}}_{\text{Yop1}}$ is significantly far from optimal, as shown in Fig.
+4.
+If setting ${\boldsymbol{x}}_{\text{ASL}}$ as the initial solution, then a near-optimal solution ${\boldsymbol{x}}_{\text{Yop2}}$ is obtained.
+The terminal times of ${\boldsymbol{x}}_{\text{ASL}}$, ${\boldsymbol{x}}_{\text{Yop1}}$, and ${\boldsymbol{x}}_{\text{Yop2}}$ are 5.541, 6.896, and 5.624, respectively.
+Furthermore, costates can be reconstructed from states ${\boldsymbol{x}}_{\text{ASL}}$ through meticulous derivations, as shown in Fig.
+4(d).
+The information on costates also verifies the proposed state-centric necessary condition in a way, noting that PMP (7) holds in this example.
+In contrast, the costates of ${\boldsymbol{x}}_{\text{Yop1}}$ and ${\boldsymbol{x}}_{\text{Yop2}}$ do not exist since Yop does not solve a problem based on costates.
+Figure 4: A feasible BBS trajectory with system behaviors, tangent markers, and additional end-constraints.
+In (d), $\hat{\lambda}\triangleq\left(2\lambda_{3}-\lambda_{4}\right)\xi\left(t\right)$ where $\xi\left(t\right)>0$ serves as a scaling factor to enhance the visibility of ${\mathrm{sgn}}\left(2\lambda_{3}-\lambda_{4}\right)$.
+VII-B Determining the Optimality of a 4th-Order Trajectory
+A near-optimal trajectory planning method is proposed in our previous work [20].
+Consider a 4th-order trajectory for high-order chain-of-integrator systems (26a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig.
+5(b).
+The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$.
+Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal.
+Let $\boldsymbol{H}$} be the equality constraints induced by $\mathcal{L}$.
+Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank.
+Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (26a).
+As shown in Fig.
+5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory $\boldsymbol{x}$$^{\prime}$.
+It can be observed in Fig.
+5(b2) that $x_{3}^{\prime}$ is tangent to $x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory.
+Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of $t_{\mathrm{f}}$$^{\prime}$ activates a new constraint.
+Specifically, $\boldsymbol{x}$$^{\prime}$ achieves a 0.06% reduction in the terminal time compared to $\boldsymbol{x}$}.
+Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
+Figure 5: Optimizing a 4th-order near-optimal trajectory.
+(a) Plot of the locally minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$.
+(b) The original and the optimized trajectories.
+VII-C Optimizing a 5th-Order Feasible BBS Trajectory
+Arc-based [35] and discretized optimization [36] methods are two main approaches to solving high-order problems.
+Consider a 5th-order problem for high-order chain-of-integrator systems (26a) where ${\boldsymbol{x}}_{0}={\boldsymbol{0}}$, ${{\boldsymbol{x}}_{\mathrm{f}}}={\boldsymbol{e}}_{5}$, $u_{\mathrm{m}}=1$, and ${\boldsymbol{x}}_{\mathrm{m}}=\left(0.8,0.5,0.5,0.5,1\right)$.
+A feasible BBS trajectory ${\boldsymbol{x}}_{1}$ is planned by [35], as shown in Fig.
+6(a).
+Denote $\mathrm{DOF}\triangleq N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|-n$; hence, Corollary 2 implies that $\mathrm{DOF}\leq 0$ holds for an optimal trajectory.
+As shown in Fig.
+6(b), $\mathrm{DOF}_{1}=6>0$ which does not satisfy the proposed state-centric necessary condition.
+Fix the first 5 arcs as well as the ASL, and vary the motion time of the 6th arc to search the minimal terminal time until reaching the boundary of the feasible set.
+If some arcs reach zero time, then delete the system behavior from the ASL.
+Repeat the above process, resulting in trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$.
+As shown in Figs.
+6(a) and 7(b), the terminal time decreases through iteration, and the DOF finally reaches 0 after 5 iterative steps.
+Finally, the optimized trajectory ${\boldsymbol{x}}_{6}$ achieves a 17.7% reduction in the terminal time compared to the original trajectory ${\boldsymbol{x}}_{1}$.
+In fact, it remains challenging to plan near-optimal high-order trajectories using arc-based methods despite advantages in computational accuracy and analytical simplicity.
+The proposed state-centric condition has the potential to optimize BBS trajectories planned by arc-based methods, which will be investigated in future works.
+Figure 6: Iteration of a 5th-order feasible BBS trajectory. The initial trajectory ${\boldsymbol{x}}_{1}$ is planned by [35]. (a) The states and controls. (b) The ASLs.
+A discretized optimization method [36] is applied to plan the trajectory ${\boldsymbol{x}}_{\text{base}}$ for comparison, where the time domain is discretized into 1000 intervals.
+An optimal solution is obtained with a 0.016% relative error of terminal time compared to ${\boldsymbol{x}}_{6}$, as shown in Fig.
+7(b).
+Therefore, [36] can plan optimal or near-optimal trajectories for the high-order problem.
+However, $u_{\text{base}}$ exhibits oscillation, as shown in Fig.
+7(a1), while $u_{i}$ is BBS.
+Defining the total variation of the control as $T_{\mathrm{v}}\triangleq\int_{0}^{{t_{\mathrm{f}}}}\left|u\left(t\right)\right|\mathrm{d}t$ to describe the trajectory stability, ${\boldsymbol{x}}_{5}$ achieves an 8.9% $T_{\mathrm{v}}$ of ${\boldsymbol{x}}_{\text{base}}$.
+For discretized optimization like [36], the open-loop errors of terminal states are difficult to constrain due to numerical error in each interval.
+Define $E_{\mathrm{s}}=\sqrt{\sum_{k=1}^{n}\left(\frac{\hat{x}_{{\mathrm{f}}k}-x_{{\mathrm{f}}k}}{x_{\mathrm{m}k}}\right)^{2}}$ where $\hat{\boldsymbol{x}}_{\mathrm{f}}=\left(\hat{x}_{{\mathrm{f}}k}\right)_{k=1}^{n}$ is directly integrated from the control input.
+As shown in Fig.
+7(d), arc-based trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$, reduce at least 5 orders of magnitude of $E_{\mathrm{s}}$ compared to ${\boldsymbol{x}}_{\text{base}}$.
+Hence, an approach combining arc-based methods and the proposed state-centric necessary condition has the potential to plan near-optimal trajectories with higher trajectory stability and computational accuracy compared to discretized optimization.
+VIII Conclusion
+This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input.
+The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form.
+Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility.
+Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one.
+The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be of full row rank.
+In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.
+The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrator systems with full box constraints from both theoretical and numerical perspectives.
+An upper bound of the number of arcs is determined in a general sense, and a recursive equation for the junction time in chattering induced by 2nd-order constraints is derived.
+Note that the above two conclusions are challenging to prove by costate-based conditions.
+Numerical experiments showed that the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information, and it has the potential to optimize feasible BBS trajectories with higher trajectory stability and computational accuracy compared to discretized optimization methods.
+Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [35], the optimized trajectory ${\boldsymbol{x}}_{6}$ based-on the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [36].
+(a) The states and controls.
+(b) The terminal time.
+(c) The total variation of the control.
+(d) The open-loop error of terminal states.
+Acknowledgment
+The authors would like to thank the reviewers for their valuable comments.
+This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB4302003, and in part by the National Natural Science Foundation of China under Grant 624B2077.
+Appendix A Proofs of Propositions and Theorems
+A-A Proofs in Section III
+Proof of Proposition 1.
+Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period.
+By (5), ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ implies that ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ holds a.e.
+Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$.
+Therefore,
+$\forall k\in{\mathbb{N}},\,{\boldsymbol{\lambda}}^{\top}{\boldsymbol{A}}^{k}{\boldsymbol{b}}=\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\left({\boldsymbol{\lambda}}^{\top}{\boldsymbol{b}}\right)\equiv 0,$ (32)
+i.e., ${\boldsymbol{\lambda}}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\equiv{\boldsymbol{0}}$.
+However, by (2), ${\boldsymbol{\lambda}}\equiv 0$ holds a.e.
+Then, (8) implies that $\lambda_{0}=0$, which contradicts $\left(\lambda_{0},{\boldsymbol{\lambda}}\right)\not=0$.
+Therefore, ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\not\equiv 0$ holds a.e.
+∎
+Proof of Proposition 2.
+Since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$, the controllability condition (2) implies ${\boldsymbol{c}}_{p}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\not={\boldsymbol{0}}$; hence, $r_{p}$ exists and serves as the minimum non-zero column index of a non-zero matrix.
+For $r\in\left[r_{p}-1\right]$, considering the $r$th-order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (10) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$.
+Hence, (9) holds.
+Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (10).
+Under the driving of (9), note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}=0$; hence, (10) holds in $\left[t_{1},t_{2}\right]$.
+∎
+Proof of Lemma 1.
+The proof of (11) is elementary [37].
+Assume that $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
+Considering the minimal polynomial of $\boldsymbol{A}$}, $\exists p^{*}$, s.t.
+$p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$.
+Therefore, $\forall r\geq n$, $\exists p_{r}^{*}$ s.t.
+$p_{r}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{r}^{*}\left({\boldsymbol{A}}\right)$.
+Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$.
+Since $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, Note that $\forall r\geq n+1$,
+$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$ (33)
+It can be proved that ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)$ is analytic. By (33), $\forall t\in{\mathbb{R}}$,
+${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right).$ (34)
+In other words, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on $\mathbb{R}$}. ∎
+Proof of Proposition 3.
+Assume that $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in the constrained arc.
+By Proposition 2, (9) holds $\forall p\in\mathcal{P}$.
+$\forall q\in\mathcal{P}$, since ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv{\mathrm{const}}$, (12) can be derived through considering the $r$th-order derivative of ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}$ for $r\in\left[n\right]$.
+Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t.
+${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (12).
+$\exists p\in\mathcal{P}$, (12) holds in $\left[t_{1},t_{2}\right]$.
+Then, $\forall q\in\mathcal{P}$, $r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds.
+By Lemma 1, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv-d_{q}$ holds.
+∎
+Proof of Proposition 4.
 Consider the connection between two unconstrained arcs.
@@
 Consider the connection between two constrained arcs.
-Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not=\varnothing$.
+Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not$$\varnothing$.
 Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration.
@@
 ∎
-Remark.
-Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection.
-Note that the equality constraints at the connection time, i.e., (11), (20), (22), and (24), are allowed linearly dependent.
-When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
-IV Keypoints for the Trajectory’s Feasibility
-The switching law proposed in Section III can describe a given optimal trajectory and the feasibility near connections of arcs, while the feasibility during an arc remains inadequately investigated.
-To address this limitation, this section investigates the feasibility of the optimal trajectory on each arc, especially of keypoints in an arc.
-As formally defined in Definition 5, a keypoint, where the trajectory touches state constraints’ boundaries in isolation, directly influences the feasibility of the perturbed trajectory.
-For the example shown in Fig.
-1, the time points at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, $\mathcal{T}_{2}^{(2)}$, and ends of arcs are keypoints.
-Section IV-A analyzes additional constraints of an arc’s end besides Proposition 4.
-The tangent marker is proposed in Section IV-B to describe the feasibility of a BBS trajectory tangent to constraints’ boundaries.
-Finally, Section IV-C proposes the augmented switching law and provides a sufficient and necessary condition for the feasibility near the keypoints.
-IV-A Additional Constraints at the End of An Arc
-Proposition 4 fully discusses the feasibility of the connection of two adjacent arcs w.r.t.
-constraints induced by $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$.
-However, the constraints on $u$ and ${\boldsymbol{x}}$, except for those in $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ still lacks discussion.
-The following proposition provides a sufficient and necessary condition for the feasibility at the end of an arc.
-Proposition 5.
-Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is a system behavior for $t\in\left[t_{0},t_{1}\right]$.
-  1. 1.
-    $\forall p\not\in\mathcal{P}$, $\exists\varepsilon>0$, s.t.
-${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\leq 0$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds:
-    1. (a)
-      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}<0$.
-    2. (b)
-      ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0$.
-Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
-$\forall r\in\left[\hat{r}_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)=0$, and ${\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{0}\right)+\widehat{{\boldsymbol{b}}}\right)<0$.
-    Similar conclusions hold for $\left[t_{1}-\varepsilon,t_{1}\right]$.
-  2. 2.
-    Assume that ${\mathcal{S}}$ is a constrained arc.
-Then, $\exists\varepsilon>0$, s.t.
-$u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if one of the following conditions holds, where $p\in\mathcal{P}$:
-    1. (a)
-      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<{u_{\mathrm{m}}}$.
-    2. (b)
-      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
-Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t.
-$\forall r\in\left[\hat{r}_{p}-1\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
-    3. (c)
-      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$.
-Furthermore, $\forall r\in\left[n\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
-    Similar conclusions hold for $u\geq-{u_{\mathrm{m}}}$ and $\left[t_{1}-\varepsilon,t_{1}\right]$.
-Proof.
+A-B Proofs in Section IV
+Proof of Proposition 5.
 Proposition 5.1 holds for the same reason as that of Proposition 4.
@@
 (2) is equivalent to $\forall r\in\left[0,n\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$.
-Hence, Proposition 5.2 holds.
 ∎
-${\mathcal{L}}=\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(1)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{(2)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{2}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{(N)}\right\}_{k=1}^{M_{N}}\right)$ (28)
-Consider two connected arcs ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
-Proposition 4 guarantees the feasibility of constraints active along ${\mathcal{S}}_{1}$ or ${\mathcal{S}}_{2}$.
-If some constraints are active at the ends of arcs but are inactive in ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ a.e., then additional constraints at the connection, as defined in Definition 3, should be introduced for feasibility based on Proposition 5.
-Definition 3.
-Consider two arcs ${\mathcal{S}}_{i}$ with active constraints $\mathcal{P}_{i}$ for $t\in\left[t_{i-1},t_{i}\right]$, $i\in\left\{1,2\right\}$.
-$\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{1}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank.
-By Proposition 5, state constraints except $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ and control constraints induce $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
-Then, the additional end-constraint of ${\boldsymbol{x}}\left(t_{1}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$.
-Furthermore, if $\widehat{{\boldsymbol{C}}}$ and $\widehat{{\boldsymbol{F}}}$ are empty matrices, then ${\mathcal{E}}$ is called empty.
-Remark.
-Given a feasible arc, the order $\hat{r}_{p}$ in Proposition 5 is directly determined.
-In Definition 3, the additional end-constraint ${\mathcal{E}}$ does not include the equality constraints induced by $\mathcal{P}_{1}\cup\mathcal{P}_{2}$.
-To avoid over-determination, the equality constraints in ${\mathcal{E}}$ and $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ are made linearly independent through elimination.
-The elimination process is guaranteed by the feasibility of the given trajectory.
-IV-B Tangent Markers in An Arc
-Proposition 4 and Proposition 5 analyzed the feasibility of an arc near the end points.
-An arc might be tangent to some constrained boundary, where small perturbations may affect the feasibility of the arc.
-The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to constraints’ boundaries.
-Proposition 6.
-Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $\left[t_{0},t_{2}\right]$.
-${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists\hat{r}\in\left[n\right]$ is even, s.t.
-$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall r\in\left[\hat{r}-1\right],\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
-Proof.
-${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0$.
-$\exists\varepsilon>0$, s.t.
-$B_{\varepsilon}\left(t_{1}\right)\subset\left(t_{0},t_{2}\right)$ and $\forall t\in{\left.B_{\varepsilon}\left(t_{1}\right)\middle\backslash\left\{t_{1}\right\}\right.}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d<0$.
-Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)={\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}\right)$.
+Proof of Proposition 6.
+[…]
 Assume that $\forall r\in\left[n\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$.
@@
 $\forall r\in\left[\hat{r}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
-If $\hat{r}$ is odd, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ strictly crosses 00 at $t_{1}$, which contradicts the tangent condition.
+If $\hat{r}$ is odd, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ strictly crosses 0 at $t_{1}$, which contradicts the tangent condition.
 Hence, $\hat{r}$ is even.
@@
 ∎
-No matter an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained or not, ${\mathcal{S}}$ can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$.
-It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$.
-For this case, ${\mathcal{S}}$ should be constrained.
-Arbitrarily consider $p\in\mathcal{P}$.
-According to Proposition 2, ±$\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$.
-In other words, the constraints on $u$ are equivalent to the constraints on ${\boldsymbol{x}}$.
-Hence, Proposition 6 can apply for both state constraints and control constraints.
-For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$.
-The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$, respectively.
-Another noteworthy fact is that an arc ${\mathcal{S}}$ can be tangent to multiple constraints $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$ at a time point $t_{1}$.
-Then, the necessary and sufficient condition for feasibility is $\forall p\in\mathcal{P}^{\prime}$, (29) holds.
-The tangent marker is defined as follows:
-Definition 4.
-Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $t\in\left[t_{0},t_{2}\right]$.
-Assume that at $t_{1}\in\left(t_{0},t_{2}\right)$, the arc ${\mathcal{S}}$ is tangent to some constraints’ boundaries whose indices constitute a nonempty set $\mathcal{P}^{\prime}\subset\left[P+2\right]$.
-By Proposition 6, the tangent condition (29) is equivalent to $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank.
-Then, the tangent marker is defined as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$.
-Remark.
-In Definition 4, the equality constraints induced by (29) can be linearly dependent.
-However, for a given feasible trajectory, the equalities should have a feasible solution; hence, equality constraints with full row rank, i.e., $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, can be constructed through elimination.
-Therefore, the tangent marker is well-defined.
-IV-C Augmented Switching Law
-Section III, Section IV-A, and Section IV-B analyze the feasibility of the optimal trajectory at the end points of arcs and the tangent points to the constrained boundaries.
-To characterize the optimal trajectory, this section proposes the augmented switching law, which represents the input control and the feasibility of the trajectory in a compact form.
-Definition 5.
-Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$.
-During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$.
-At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$.
-Then, time points at tangent markers and ends of arcs are called keypoints for the trajectory’s feasibility.
-The augmented switching law (ASL) ${\mathcal{L}}$ is defined in (28).
-Remark.
-If a feasible trajectory with a finite number of arcs follows the BBS control law (7), then Propositions 4, 5, and 6 hold for the trajectory.
-For this case, the ASL can also represent the feasible trajectory.
-Remark.
-If a feasible non-chattering BBS trajectory is perturbed by a sufficiently small perturbation, then the perturbed trajectory can only exceed constraints’ boundaries at the keypoints since the distance between the original trajectory and constraints’ boundaries is 0 at the keypoints, while it is strictly positive at other points.
-Therefore, the ASL can not only characterize the input control but also describe the feasibility of the trajectory.
-Remark.
-The switching law in Definition 2 corresponding to the ASL in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
-It can be observed that the ASL provides more detailed information on the feasibility of the trajectory at keypoints than the switching law.
-Notably, neither the switching law nor the ASL provides the motion time of each arc.
-To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an ASL.
-If ${\mathcal{E}}$ is empty, then ${\mathcal{E}}$ is omitted.
-The additional end-constraints before ${\mathcal{S}}_{1}$ and after ${\mathcal{S}}_{N}$ are omitted since the initial and terminal states are given, i.e., ${\boldsymbol{x}}_{0}$ and ${\boldsymbol{x}}_{\mathrm{f}}$.
-For the example shown in Fig.
-1, the ASL is $\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{1}^{(1)}\right\}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{2},\left\{{\mathcal{T}}_{2}^{(1)},{\mathcal{T}}_{2}^{(2)}\right\}\right){\mathcal{S}}_{3}$.
-For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 32].
-The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively.
-The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively.
-Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$.
-The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively.
-The additional end-constraint =${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively.
-Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc.
-Based on the above simplified notations, an example is provided to illustrate the ASL.
-Example 3.
-Consider a position-to-position problem for 4th-order chain-of-integrator systems with full box constraints shown in Fig.
-2.
-A feasible suboptimal trajectory with a BBS form, planned by the manifold-intercepted method (MIM) [20], is shown in Fig.
-2(a).
-It can be observed that the ASL is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
-Figure 2: Two examples of ASLs.
-For a 4th-order chain-of-integrator system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$.
-(a) A non-chattering feasible trajectory planned by [20].
-(b) A chattering optimal trajectory planned by [32].
-(b2) is the plot of $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
-Example 4.
-Our previous work [32] pointed out that the chattering phenomenon occurs in the optimal trajectory for the problem in Example 3.
-The chattering optimal trajectory is shown in Fig.
-2(b1).
-Fig.
-2(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied.
-It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$.
-According to Bellman’s principle of optimality [22], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is sufficiently small.
-The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
-V A State-Centric Necessary Condition
-The theoretical framework for the ASL of time-optimal control for linear systems is established in Section IV.
-Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a).
-Section V-A discusses the uniqueness of the time-optimal control.
-Then, given a feasible BBS trajectory, Section V-B analyzes the feasibility of the perturbed BBS trajectory.
-Based on the conclusions in Sections V-A and V-B, Section V-C provides a necessary condition of optimality.
-V-A Uniqueness of Time-Optimal Control
-The uniqueness of the time-optimal control for problem (1a) is proved in Theorem 1 as preliminaries.
-Theorem 1.
-Under Assumption 1, the time-optimal control of problem (1a) is unique in an a.e.
-sense.
-In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal, then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e.; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
-Proof.
+A-C Proofs in Section V
+Proof of Theorem 1.
 If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then it holds a.e.
@@
 Therefore, both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are constrained arcs.
-Assume that $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}\not=\varnothing$.
+Assume that $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}\not$$\varnothing$.
 Arbitrarily consider $p\in\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}$.
@@
 ∎
-Remark.
-Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an a.e.
-sense when there exists at most one chattering limit point.
-The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
-According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time.
-The above observation will be applied in Section V-B.
-V-B Feasibility of the Perturbed BBS Trajectory
-[…]
-The feasibility of the perturbed trajectory is proved by two steps.
-Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory.
-In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory.
-Secondly, Theorem 2 proves the perturbed trajectory’s feasibility based on Propositions 4, 5, 6, and 7.
-In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible.
-Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
-Proposition 7.
-Consider a BBS trajectory ${\boldsymbol{x}}\left(t\right)$ satisfying $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$.
-Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$.
-$\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
-$\forall\varepsilon\in\left(0,\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}\right)$, $\forall{\boldsymbol{x}}^{\prime}\left(t\right)$ is a perturbed trajectory induced by ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ with $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$,
-$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon$ (30)
-where $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$ and $t_{0}^{\prime}=t_{0}$.
-Proof.
+Proof of Proposition 7.
 Since $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$, it holds that $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$.
@@
 $\begin{dcases}{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau,\\ {\boldsymbol{x}}^{\prime}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)+\int_{\hat{t}_{2i}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\tau\right)}{\boldsymbol{b}}_{i}\mathrm{d}\tau.\end{dcases}$ (38)
-Therefore, $\forall t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$, it holds that
-$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$ (32)
-$\leq$ $C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|.$
-$\forall i\in\left[M\right]$, let
-$\begin{dcases}j=i+1,\,j^{\prime}=i,&\text{if }t_{i}\leq t_{i}^{\prime},\\ j=i,\,j^{\prime}=i+1,&\text{if }t_{i}>t_{i}^{\prime}.\end{dcases}$ (33)
+Therefore, $\forall t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$, $\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\leq C_{3}$ implies that
+$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|.$ (39)
+$\forall i\in\left[M\right]$, let (a) $j=i+1$, $j^{\prime}=i$ if $t_{i}\leq t_{i}^{\prime}$; (b) $j=i$, $j^{\prime}=i+1$ if $t_{i}>t_{i}^{\prime}$.
 Then, $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{j}{\boldsymbol{x}}+{\boldsymbol{b}}_{j}$ and $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{j^{\prime}}{\boldsymbol{x}}+{\boldsymbol{b}}_{j^{\prime}}$.
@@
 $\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\leq 2\left(C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$ (42)
-Therefore, (30) holds.
-Theorem 2.
-Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is ${\mathcal{L}}$ in problem (1a).
-Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t.
-$\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is trajectory also represented by ${\mathcal{L}}^{\prime}={\mathcal{L}}$, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
-${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by ${\boldsymbol{t}}$ and ${\boldsymbol{t}}^{\prime}$ in the same way as that in Proposition 7.
-Proof.
+Therefore, (24) holds. ∎
+Proof of Theorem 2.
 Fix ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$. Firstly, let
@@
 Furthermore, $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds.
-Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, the above property on $\left({\boldsymbol{x}}\left(t\right),{\boldsymbol{t}}\right)$ holds for $\left({\boldsymbol{x}}^{\prime}\left(t\right),{\boldsymbol{t}}^{\prime}\right)$ as well.
+Since $\mathcal{L}$$\mathcal{L}$$^{\prime}$, the above property on $\left({\boldsymbol{x}}\left(t\right),{\boldsymbol{t}}\right)$ holds for $\left({\boldsymbol{x}}^{\prime}\left(t\right),{\boldsymbol{t}}^{\prime}\right)$ as well.
 Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig.
-3(a).
-${\mathcal{L}}={\mathcal{L}}^{\prime}$ implies that ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\equiv 0$ holds.
+8(a).
+$\mathcal{L}$$\mathcal{L}$$^{\prime}$ implies that ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\equiv 0$ holds.
 Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right)$.
 Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig.
-3(b).
+8(b).
 The feasibility of ${\boldsymbol{x}}^{\prime}\left(t\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right$] is discussed in two parts, i.e., far from and near the keypoints.
@@
 Then, $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ is continuous in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$.
-If $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ holds, then let $r_{j}^{(p)}=0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$; otherwise, according to Propositions 4, 5, and 6, $\exists r_{j}^{(p)}\in\left[n\right]$, s.t.
-$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ and $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$.
+If $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ holds, then let $r_{j}^{(p)}=0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$; otherwise, by Propositions 4, 5, and 6, $\exists r_{j}^{(p)}\in\left[n\right]$, s.t.
+$f_{p,j}\triangleq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ and $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$.
 Let $t_{1,j}^{(p)}\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$ is sufficiently small, s.t.
-$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ holds $\forall t\in\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$.
+$\forall t\in\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$, $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)<\frac{1}{2}f_{p,j}<0$ holds.
 $t_{2,j}^{(p)}\leq t_{i^{(p)}_{j}}$ is constructed similarly.
 Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{(p)},t_{i^{(p)}_{j}}\right)$.
+Figure 8: Feasibility of the perturbed trajectory near the keypoints. (a) Constrained arc. (b) Unconstrained arc.
 Due to the continuity of $f_{p}$, $\sup_{t\in\left[t_{1,j}^{(p)},t_{2,j}^{(p)}\right]}f_{p}\left(t\right)<0$ holds.
@@
 $f_{p}\left(t\right)\leq\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right$].
-By (30) and (38), $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$, $f_{p}^{\prime}\left(t\right)\leq 0$ is feasible since
-$f_{p}^{\prime}\left(t\right)\leq$ $f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (40)
-$\leq$ $\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)-\frac{1}{4}f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0.$
-Figure 3: Feasibility of the perturbed trajectory near the keypoints. (a) Constrained arc. (b) Unconstrained arc.
-For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)=0$ and $t_{i_{j-1}^{(p)}}<t_{1,j}^{(p)}$, let
-$\varepsilon<\hat{\varepsilon}\triangleq-\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}+1\right)^{-1}.$ (41)
+By (24) and (44), $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right$], $f_{p}^{\prime}\left(t\right)\leq 0$ is feasible since $f_{p}^{\prime}\left(t\right)\leq f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\leq\frac{1}{4}f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$.
+For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)=0$ and $t_{i_{j-1}^{(p)}}<t_{1,j}^{(p)}$, let $\varepsilon<\hat{\varepsilon}\triangleq-f_{p,j}\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}+1\right)^{-1}$.
 Hence, $\forall t\in\left[t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right]$, it holds that
-$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
-$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$
+$\leq$ $\frac{1}{2}f_{p,j}-\frac{1}{4}f_{p,j}<0.$ (46)
 If $t_{i^{(p)}_{j-1}}^{\prime}<t_{i^{(p)}_{j-1}}$, let $\varepsilon$ be sufficiently small, s.t.
@@
 Then, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]\subset\left[t_{i^{(p)}_{j-1}}-\varepsilon,t_{i^{(p)}_{j-1}}\right$], it holds that
-$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r_{j}^{(p)}-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ (44)
+$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r_{j}^{(p)}-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$
 $+$ $\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left(\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)\right\|\right)$
-$\leq$ $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
+$\leq$ $f_{p,j}-\frac{1}{2}f_{p,j}<0.$ (48)
 Therefore, once $\varepsilon$ is sufficiently small, then $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}<0$ holds in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$.
-${\mathcal{L}}^{\prime}={\mathcal{L}}$ implies that $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right)$.
+$\mathcal{L}$$^{\prime}$$\mathcal{L}$} implies that $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right)$.
 Therefore, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible.
@@
 Thirdly, the control constraints $\left|u^{\prime}\right|\leq{u_{\mathrm{m}}}$ can be proved by applying a similar analysis when $\varepsilon$ is sufficiently small. ∎
-Theorem 2 points out that if the perturbed trajectory is represented by the same ASL to the original trajectory and the perturbance is small enough, then the perturbed trajectory is feasible.
-In other words, Theorem 2 provides a novel variational approach under the fixed ASL with a feasibility guarantee near constraints’ boundaries.
-V-C State-Centric Necessary Condition of Optimality
-Based on Theorem 2, this section provides an approach to constructing a feasible perturbed trajectory with the same ASL.
-According to Theorem 1, this section shows that the existence of such a perturbed trajectory implies that the original trajectory is not optimal, resulting in a state-centric necessary for non-chattering optimal control of problem (1a).
-The Jacobian matrix of equality constraints induced by the ASL can be calculated through Proposition 8.
-Proposition 8.
-${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$, and ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ is fixed.
-Consider the mapping ${\boldsymbol{x}}_{i}:{\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}\mapsto{\boldsymbol{x}}\left(t_{i}\right)$.
-Then, $\forall i,j\in\left[M\right]$,
-$\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\begin{dcases}{\boldsymbol{P}}_{j,i}\left(\left({\boldsymbol{A}}_{j}-{\boldsymbol{A}}_{j+1}\right){\boldsymbol{x}}_{j}+{\boldsymbol{b}}_{j}-{\boldsymbol{b}}_{j+1}\right),\,j<i,\\ {\boldsymbol{A}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{b}}_{i},\,j=i,\\ {\boldsymbol{0}},\,j>i,\end{dcases}$ (45)
-where ${\boldsymbol{P}}_{j,i}={\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}{\mathrm{e}}^{{\boldsymbol{A}}_{i-1}\left(t_{i-1}-t_{i-2}\right)}\dots\mathrm{e}^{{\boldsymbol{A}}_{j+1}\left(t_{j+1}-t_{j}\right)}$.
-Proof.
+Proof of Proposition 8.
 $\forall j>i$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}={\boldsymbol{0}}$ holds evidently.
@@
 $=$ ${\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t_{i}-t_{i-1}\right)}\left(\left({\boldsymbol{A}}_{i-1}-{\boldsymbol{A}}_{i}\right){\boldsymbol{x}}_{i-1}+{\boldsymbol{b}}_{i-1}-{\boldsymbol{b}}_{i}\right).$
-$\forall i>1$, $j<i-1$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ holds, which implies (45).
+So $\forall i>1$, j<i1j<i-1, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$.
 ∎
-Utilizing the notations in Section V-B, the ASL provides a system of equalities on every keypoints.
-Denote that $\forall i\in\left[M\right]$, ${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{g}}_{i}={\boldsymbol{0}}$ and ${\boldsymbol{C}}_{i}{\boldsymbol{x}}_{i}+{\boldsymbol{d}}_{i}<{\boldsymbol{0}}$, where ${\boldsymbol{x}}_{i}={\boldsymbol{x}}\left(t_{i}\right)$ and ${\boldsymbol{F}}_{i}$ has full row rank.
-For a constrained arc with active constraints $\mathcal{P}$, the equality constraints induced by $\mathcal{P}$ are added in only one end instead of both ends of the arcs since one end satisfies $\mathcal{P}$ implies that the whole arc satisfies $\mathcal{P}$.
-Note that the additional end-constraints at both ends are eliminated by equality constraints induced by $\mathcal{P}$.
-Similar processes are applied if $u\equiv{u_{\mathrm{m}}}$ or $u\equiv-{u_{\mathrm{m}}}$ during a constrained arc.
-${\boldsymbol{F}}_{i}{\boldsymbol{x}}_{i}\left({\boldsymbol{t}}\right)+{\boldsymbol{g}}_{i}$ for $i\in\left[M\right]$ induce a function ${\boldsymbol{H}}\left({\boldsymbol{t}}\right)$.
-Then, Theorem 3 provides a necessary condition of optimality.
-Theorem 3 (State-Centric Necessary Condition).
-Assume that the optimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{*}\left(t\right)$ of problem (1a) satisfies Assumption 1.
-Denote the equality constraints induced by the ASL as ${\boldsymbol{H}}\left({\boldsymbol{t}}^{*}\right)={\boldsymbol{0}}$, where ${\boldsymbol{t}}^{*}=\left(t_{i}^{*}\right)_{i=1}^{M}$ is the arriving time of each keypoints and $t_{0}=0$.
-Then, $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}_{1:\left(M-1\right)}}\left({\boldsymbol{t}}^{*}\right)$ does not have full row rank.
-Proof.
+$\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
+Proof of Theorem 3.
 Assume that ${\boldsymbol{H}}:{\mathbb{R}}^{M}\to{\mathbb{R}}^{M^{\prime}}$.
@@
 ∎
-Remark.
-In contrast to existing costate-based necessary conditions [9], Theorem 3 requires no information on costates, which significantly simplifies the computation.
-Compared to existing state-centric necessary conditions [24], Theorem 3 is effective in theoretical reasoning and has a lower computational complexity in numerical computation.
-Some examples are provided in Sections VI-B and VI-C.
-Remark.
-[…]
-Remark.
-If a given non-chattering BBS trajectory does not satisfy the necessary condition in Theorem 3, then one can perform numerical optimization of the trajectory based on Theorem 3.
-Numerical examples are provided in Section VII.
-VI Applications in Chain-of-Integrator Systems
-This section provides some applications of the state-centric necessary condition, i.e., Theorem 3, for the optimal control of problem (1a).
-For convenience, the chain-of-integrator system with box constraints is considered in this section, where the notations have been introduced in Section IV-C.
-Time-optimal control for chain-of-integrator systems with box constraints is an open and challenging problem in the optimal control domain, yet to be resolved.
-The problem can be summarized as follows:
-$\min\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (48a)
-[…]
-$\dot{x}_{1}\left(t\right)=u\left(t\right),\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (48c)
-${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}},$ (48d)
-[…]
-$\left|u\left(t\right)\right|\leq{u_{\mathrm{m}}},\,\forall t\in\left[0,t_{\mathrm{f}}\right].$ (48f)
-This section provides some applications of Theorem 3.
-Section VI-A proves a trivial conclusion from a state-centric perspective.
-Section VI-B and Section VI-C prove two corollaries that are challenging to prove by traditional costate-based necessary conditions.
-VI-A The Case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$
-The case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$ has a well-known conclusion that the optimal control switches no more than $\left(n-1\right)$ times [33].
-This section proves the above conclusion based on the proposed Theorem 3.
-Corollary 1.
-In problem (48a), assume that $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$. Then, the optimal control switches no more than $\left(n-1\right)$ times.
-Proof.
+A-D Proofs in Section VI
+Proof of Corollary 1.
 Chattering does not occur since no state inequality constraints exist.
@@
 In other words, $\forall i\in\left[N\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $u\left(t\right)\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$.
-$u_{i}\not=u_{i-1}$, implies that $\Delta u_{i}\not=0$.
-Then,
-${\boldsymbol{J}}=\left[-\Delta u_{i+1}{\boldsymbol{\phi}}\left(t_{N}-t_{i}\right)\right]_{i=1}^{N-1}\in{\mathbb{R}}^{n\times\left(N-1\right)}.$ (49)
-The Vandermonde form of ${\boldsymbol{J}}$ implies that ${\mathrm{rank}}{\boldsymbol{J}}=\min\left\{n,N-1\right\}$. Theorem 3 implies that ${\mathrm{rank}}{\boldsymbol{J}}<n$; hence, $N-1<n$. ∎
-Corollary 1 implies that the switching law is in the form of $\overline{0}\underline{0}\overline{0}\underline{0}\dots$ or $\underline{0}\overline{0}\underline{0}\overline{0}\dots$ with no more than $n$ arcs.
-The case in Section VI-A is trivial since no state constraints exist and the costate vector follows a fixed equation that $\forall k\in\left[n-1\right]$, $\dot{\lambda}_{i}=-\lambda_{i+1}$ and $\lambda_{n}\equiv{\mathrm{const}}$.
-However, when the state constraints are induced, the behavior of the optimal control can be complex.
-On the one hand, the multiplier ${\boldsymbol{\eta}}$ in (4) can be non-zero during a constrained arc.
-On the other hand, the junction of ${\boldsymbol{\lambda}}$ can occur when a state constraint switches between active and inactive [7].
-For this case, the costate analysis can be complex, while the proposed state-centric necessary condition provides a simple and effective approach for analyzing the optimal control.
-VI-B The Case where Only System Behaviors Occur
-The case where only system behaviors occur, i.e., the optimal trajectory consists of a finite number of unconstrained arcs and constrained arcs that are not tangent to constraints’ boundaries, is widely applied in existing works on trajectory planning [19, 34, 20].
-Denote the ASL of the trajectory as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}{\mathcal{E}}_{N}$, where ${\mathcal{E}}_{N}$ is induced by ${\boldsymbol{x}}\left(t_{N}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$.
-$\forall k\in{\mathbb{N}}$, denote $\left|\overline{k}\right|=\left|\underline{k}\right|=k$, ${\mathrm{sgn}}\left(\overline{k}\right)=+1$, and ${\mathrm{sgn}}\left(\underline{k}\right)=-1$.
-The non-existence of tangent markers and additional end-constraints in ${\mathcal{L}}$ means that $\forall i\in\left[N\right]$, the arc ${\mathcal{S}}_{i}$ achieves strictly feasibility.
-An arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ in $\left[t_{0},t_{1}\right]$ is strictly feasible if $\forall p\not\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ holds in $\left[t_{0},t_{1}\right]$.
-Corollary 2.
-Assume that the ASL consists of system behaviors without tangent markers and additional end-constraints except at ${t_{\mathrm{f}}}$. Assume that:
-  1. 1.
-    $\forall 1\leq j<i$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=j+1}^{i}\left|{\mathcal{S}}_{k}\right|<i-j$.
-  2. 2.
-    $\forall i<j\leq N$, if $\left|{\mathcal{S}}_{j}\right|\geq\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{j-1}\left|{\mathcal{S}}_{k}\right|<j-i$.
-  3. 3.
-    $\forall i\in\left[N\right]$, if $\left|{\mathcal{S}}_{i}\right|>0$ and $\forall i<j\leq N$, $\left|{\mathcal{S}}_{j}\right|<\left|{\mathcal{S}}_{i}\right|$, then $\sum_{k=i}^{N}\left|{\mathcal{S}}_{k}\right|\leq N-i$.
-Then, it holds that
-$N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|\leq n.$ (50)
-$\left[\begin{array}[]{ccccccccc}-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{3}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3}-t_{2}\right)&0&0&0&0&\dots&0&0\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&-{u_{\mathrm{m}}}&\dots&0&0\\ -2{u_{\mathrm{m}}}\left(t_{6}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{6}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{6}-t_{5}\right)&0&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ -2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}&0&\dots&-2{u_{\mathrm{m}}}&2{u_{\mathrm{m}}}\\ -2{u_{\mathrm{m}}}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{2}\right)&0&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{5}\right)&0&\dots&-2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}\left(t_{3N}-t_{3N-1}\right)\\ -2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{1}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{2}\right)&{\boldsymbol{0}}&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{4}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{5}\right)&{\boldsymbol{0}}&\dots&-2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-2}\right)&2{u_{\mathrm{m}}}{\boldsymbol{\phi}}_{3:n}\left(t_{3N}-t_{3N-1}\right)\\ \end{array}\right]$
-Proof.
+Note that ${\boldsymbol{J}}=\left[-\Delta u_{i+1}{\boldsymbol{\phi}}\left(t_{N}-t_{i}\right)\right]_{i=1}^{N-1}\in{\mathbb{R}}^{n\times\left(N-1\right)}$.
+The Vandermonde form of $\boldsymbol{J}$} implies that ${\mathrm{rank}}{\boldsymbol{J}}=\min\left\{n,N-1\right\}$ since $\Delta u_{i}\not=0$.
+By Theorem 3, ${\mathrm{rank}}{\boldsymbol{J}}<n$; hence, N1<nN-1<n.
+∎
+Proof of Corollary 2.
 For a constrained arc ${\mathcal{S}}_{i}$, add the equality constraint at $t_{i}$, i.e., ${\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|}$.
 Note that if ${\mathcal{S}}_{i}$ is constrained, i.e., $\left|{\mathcal{S}}_{i}\right|\not=0$, then $u\equiv 0$ in ${\mathcal{S}}_{i}$.
-Therefore, the constraints induced by ${\mathcal{L}}$ are
-$\begin{dcases}{\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|},\,\forall i\in\left[N\right],\,\left|{\mathcal{S}}_{i}\right|\not=0,\\ {\boldsymbol{x}}_{N}={{\boldsymbol{x}}_{\mathrm{f}}}.\end{dcases}$ (52)
+Therefore, the constraints induced by $\mathcal{L}$} are ${\boldsymbol{x}}_{N}={{\boldsymbol{x}}_{\mathrm{f}}}$ and ${\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|}$, $\forall i\in\left[N\right]$, $\left|{\mathcal{S}}_{i}\right|\not=0$.
 By Condition 3, $\left|{\mathcal{S}}_{N}\right|=0$.
-Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$.
-According to Proposition 8, ${\boldsymbol{J}}$ contains the following rows:
-${\boldsymbol{J}}=\left(-\Delta u_{j+1}{\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]},$ (53)
-where $\delta_{A}$ is the indicator function of condition $A$, i.e.,
-$\delta_{A}=\begin{dcases}1,&\text{condition }A\text{ holds},\\ 0,&\text{condition }A\text{ does not hold}.\end{dcases}$ (54)
+Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and $\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$.
+According to Proposition 8, ${\boldsymbol{J}}=\left(-\Delta u_{j+1}{\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}$, where $\delta_{A}$ is the indicator function of condition $A$.
 Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that $\boldsymbol{A}$} has full row rank if and only if $\boldsymbol{B}$} has full row rank.
-By $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it holds that
-${\boldsymbol{J}}\leftrightarrow\left({\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}.$ (55)
-By iteratively taking differences between adjacent rows of (55), dividing by the time difference, and eliminating, it can be proved that ${\boldsymbol{J}}$ has full rank based on Conditions 1, 2, and 3.
+By $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it holds that ${\boldsymbol{J}}\leftrightarrow\left({\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}$.
+By iteratively taking differences between adjacent rows of the right-hand side, dividing by the time difference, and eliminating, it can be proved that $\boldsymbol{J}$} has full rank based on Conditions 1, 2, and 3.
 The proof is similar to our previous work [20] and is omitted due to space limitation.
-According to Theorem 3, the number of rows should be greater than the number of columns, i.e.,
-$\sum_{i\in\mathcal{I}}\left|{\mathcal{S}}_{i}\right|=\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|+n>N-1.$ (56)
-Therefore, (50) holds. ∎
-Remark.
-In our previous work [20], a similar conclusion pointed out that when fixing ${{\boldsymbol{x}}_{\mathrm{f}}}$, then the set of $\left({\boldsymbol{x}}_{0},{\boldsymbol{t}}\right)$ that satisfies the ASL ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$ locally forms a submanifold of dimension $r$.
-Among them, $r=N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|$.
-However, the local optimality is not discussed in [20].
-Corollary 2 proves that the ASL with $r>n$ fails to achieve optimality.
-It is challenging to reason Corollary 2 by existing costate-based necessary conditions [9] due to complex behaviors of the costates.
-In contrast to Corollary 2, PMP-based approaches need to derive the costates for a given trajectory case-by-case.
-For example, in a 4th-order problem, $\mathcal{L}_{1}=\underline{01}\overline{01}\underline{0}\overline{2}\underline{01}\overline{01}\underline{0}$ can be feasible, but it cannot be optimal since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=5>4$.
-$\mathcal{L}_{2}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$ can be a candidate for optimal control since $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=4$.
-Though Corollary 2 is not a sufficient condition for optimal control, it provides a simple and effective approach to analyzing optimal control of problem (48a).
-Numerical examples are provided in Section VII.
-VI-C The Chattering Phenomenon induced by $\left|x_{2}\right|\leq x_{{\mathrm{m}}2}$
-Our previous work [32] points out that if the chattering phenomenon occurs in problem (48a), then there exists one and only one inequality state constraint that switches between active and inactive during the chattering phenomenon.
-An infinite number of unconstrained arcs are joined by constraints’ boundaries, i.e., the unconstrained arcs are tangent to the boundary for infinite times.
-In contrast, constrained arcs do not occur.
-This section discusses the chattering phenomenon induced by $x_{2}\leq x_{{\mathrm{m}}2}$ based on the proposed state-centric necessary condition.
-[32] provides the following lemma.
-Lemma 2.
-[32] Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (48a) in a left neighborhood of $t_{\infty}$, where $t_{\infty}$ is the chattering limit point.
-Then, $\exists\left\{t_{3k}\right\}_{k=0}^{\infty}\subset\left[t_{0},t_{\infty}\right]$ increasing strictly monotonically and converging to $t_{\infty}$, s.t.
-$x_{2}$ is tangent to $x_{{\mathrm{m}}2}$ at $t_{3k}$.
-$\forall k\in{\mathbb{N}}^{*}$, $u$ switches at most 2 times during $\left(t_{3k-3},t_{3k}\right)$.
-Corollary 3.
-Assume that the chattering phenomenon is induced by $x_{2}\leq x_{{\mathrm{m}}2}$ in problem (1a) in a left neighborhood of $t_{\infty}$.
-Let $\tau_{k}=t_{\infty}-t_{3k}$, $\forall k\in{\mathbb{N}}$.
-Denote $f_{m}\left(a,b,c\right)$ as
-$\left(b+3a\right)^{m}-3\left(3b+a\right)^{m}+3\left(c+3b\right)^{m}-\left(3c+b\right)^{m}.$ (57)
-Then, $\forall N\in{\mathbb{N}}^{*}$, ${\boldsymbol{J}}^{\prime}$ does not have full row rank, where
-${\boldsymbol{J}}^{\prime}=\left(f_{i+1}\left(\tau_{j-1},\tau_{j},\tau_{j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[N\right]}.$ (58)
-Furthermore, $\forall k\geq n-2$, $\tau_{k}$ can be calculated by the following recursive equation:
-$\det\left(f_{i+1}\left(\tau_{k-j-1},\tau_{k-j},\tau_{k-j+1}\right)\right)_{i\in\left[n-2\right],j\in\left[n-2\right]}=0.$ (59)
-Proof.
-Since $\forall k\in{\mathbb{N}}^{*}$, ${\boldsymbol{x}}_{1:2}\left(t_{3k-3}\right)={\boldsymbol{x}}_{1:2}\left(t_{3k}\right)=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, it can be proved that the control $u$ in $\left(t_{3k-3},t_{3k}\right)$ is
-$\begin{dcases}u\left(t\right)\equiv u_{3k-2}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-3},t_{3k-2}\right),\\ u\left(t\right)\equiv u_{3k-1}={u_{\mathrm{m}}},&t\in\left(t_{3k-2},t_{3k-1}\right),\\ u\left(t\right)\equiv u_{3k}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-1},t_{3k}\right).\end{dcases}$ (60)
-Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$.
+According to Theorem 3, the number of rows should be greater than the number of columns, i.e., $\sum_{i\in\mathcal{I}}\left|{\mathcal{S}}_{i}\right|=\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|+n>N-1$.
+∎
+Proof of Corollary 3.
+Since $\forall k\in{\mathbb{N}}^{*}$, ${\boldsymbol{x}}_{1:2}\left(t_{3k-3}\right)={\boldsymbol{x}}_{1:2}\left(t_{3k}\right)=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, it can be proved that
+$\begin{dcases}u\left(t\right)\equiv u_{3k-2}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-3},t_{3k-2}\right),\\ u\left(t\right)\equiv u_{3k-1}={u_{\mathrm{m}}},&t\in\left(t_{3k-2},t_{3k-1}\right),\\ u\left(t\right)\equiv u_{3k}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-1},t_{3k}\right),\end{dcases}$ (51)
+where $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$.
 $\forall N\in{\mathbb{N}}^{*}$, the ASL between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$.
@@
 Considering the rank of columns $\left(k-n+3\right)$ through $k$ of $\boldsymbol{J}$$^{\prime}$, (30) holds.
-Remark.
-According to Corollary 3, the chattering phenomenon can be induced by $x_{2}\leq x_{{\mathrm{m}}2}$ only if $\lim_{k\to\infty}\tau_{k}=0$ and $\tau_{k}>0$ under the recursive equation (59).
-It is evident that chattering does not occur when $n=3$ since $\tau_{k}=2\tau_{k-1}-\tau_{k-2}$ converges to $\infty$.
-Through more refined derivation, it can be rigorously proved that $n=4$ does not hold.
-In future works, one can try to determine the existence of chattering in problem (48a) of order $n\geq 5$ under constraints on $x_{2}$, where Corollary 3 and Theorem 3 would serve as useful mathematical tools.
-However, it is challenging to obtain the recursive equation (59) for problem (48a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering.
-A comparison between state-centric and PMP-based approaches is provided in Appendix A.
-Although Theorem 3 holds under the assumption of the non-existence of the chattering phenomenon, Corollary 3 provides an example to apply the proposed necessary condition to a chattering trajectory, where a sub-arc without chattering of the optimal trajectory is investigated.
-VII Numerical Experiments
-Numerical experiments for high-order chain-of-integrator systems (48a) are conducted to provide a potential way to apply the proposed state-centric necessary condition to optimizing trajectories.
-Since this paper focuses on the theoretical necessary condition instead of optimization algorithms, metrics like computational time are not considered.
-Note that this paper does not provide a manual optimization algorithm.
-VII-A Representating a Feasible BBS Trajectory
-Consider problem (1a) whose system matrix is non-diagonalizable with multiple Jordan blocks as follows:
-${\boldsymbol{A}}=\left[\begin{array}[]{cccc}-1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right],\,{\boldsymbol{b}}=\left[\begin{array}[]{c}0\\ 0\\ 2\\ -1\end{array}\right].$ (61)
-The constraints are $\left|u\right|\leq 1$, $x_{1}\geq-0.7$, and $x_{3}+x_{4}\leq 0.5$.
-As defined in Proposition 2, $x_{1}\geq-0.7$ is a 3rd-order state constraint.
-As shown in Figs.
-4(a-c), a feasible BBS trajectory ${\boldsymbol{x}}_{\text{ASL}}$ consists of full features defined in this paper.
-The ASL is ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}{\mathcal{S}}_{4}\left({\mathcal{S}}_{5},\left\{{\mathcal{T}}_{5}^{(1)}\right\}\right){\mathcal{S}}_{6}{\mathcal{E}}_{6}{\mathcal{S}}_{7}{\mathcal{S}}_{8}$.
-Specifically, ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{4}$, and ${\mathcal{S}}_{8}$ are unconstrained arcs where $u\equiv-1$.
-${\mathcal{S}}_{2}$, ${\mathcal{S}}_{5}$, and ${\mathcal{S}}_{7}$ are unconstrained arcs where $u\equiv+1$.
-${\mathcal{S}}_{3}$ and ${\mathcal{S}}_{6}$ are constrained arcs where $x_{3}+x_{4}\equiv 0.5$.
-${\mathcal{T}}_{5}^{(1)}$ is a tangent markers where $x_{1}$ is tangent to $-0.7$.
-${\mathcal{E}}_{6}$ is an additional end-constraint requiring $u\left(t_{6}^{-}\right)=1$.
-In other words, the control $u$ is continuous at the connection of the constrained arc ${\mathcal{S}}_{6}$ and the unconstrained arc ${\mathcal{S}}_{7}$.
-In contrast, $u$ is not continuous at the two ends of another constrained arc ${\mathcal{S}}_{3}$.
-One can check that Theorem 3 holds for ${\boldsymbol{x}}_{\text{ASL}}$.
-A numerical optimal control toolbox, i.e., Yop [35], is applied to solving problem (61).
-The number of intervals is set to 200.
-The resulting trajectory without a specified initial solution ${\boldsymbol{x}}_{\text{Yop1}}$ is significantly far from optimal, as shown in Fig.
-4.
-If setting ${\boldsymbol{x}}_{\text{ASL}}$ as the initial solution, then a near-optimal solution ${\boldsymbol{x}}_{\text{Yop2}}$ is obtained.
-The terminal times of ${\boldsymbol{x}}_{\text{ASL}}$, ${\boldsymbol{x}}_{\text{Yop1}}$, and ${\boldsymbol{x}}_{\text{Yop2}}$ are 5.541, 6.896, and 5.624, respectively.
-Furthermore, costates can be reconstructed from states ${\boldsymbol{x}}_{\text{ASL}}$ through meticulous derivations, as shown in Fig.
-4(d).
-The information on costates also verifies the proposed state-centric necessary condition in a way, noting that PMP (7) holds in this example.
-In contrast, the costates of ${\boldsymbol{x}}_{\text{Yop1}}$ and ${\boldsymbol{x}}_{\text{Yop2}}$ do not exist since Yop does not solve a problem based on costates.
-Figure 4: A feasible BBS trajectory with system behaviors, tangent markers, and additional end-constraints.
-In (d), $\hat{\lambda}\triangleq\left(2\lambda_{3}-\lambda_{4}\right)\xi\left(t\right)$ where $\xi\left(t\right)>0$ serves as a scaling factor to enhance the visibility of ${\mathrm{sgn}}\left(2\lambda_{3}-\lambda_{4}\right)$.
-Figure 5: Optimizing a 4th-order near-optimal trajectory.
-(a) Plot of the locally minimal $t_{\mathrm{f}}^{\prime}-t_{\mathrm{f}}$ and $t_{4}^{\prime}-t_{4}$.
-(b) The original and the optimized trajectories.
-VII-B Determining the Optimality of a 4th-Order Trajectory
-A near-optimal trajectory planning method is proposed in our previous work [20].
-Consider a 4th-order trajectory for high-order chain-of-integrator systems (48a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig.
-5(b).
-The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$.
-Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal.
-Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$.
-Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank.
-Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (48a).
-As shown in Fig.
-5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$.
-It can be observed in Fig.
-5(b2) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory.
-Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint.
-Specifically, ${\boldsymbol{x}}^{\prime}$ achieves a 0.06% reduction in the terminal time compared to ${\boldsymbol{x}}$.
-Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
-VII-C Optimizing a 5th-Order Feasible BBS Trajectory
-Arc-based [36] and discretized optimization [37] methods are two main approaches to solving high-order problems.
-Consider a 5th-order problem for high-order chain-of-integrator systems (48a) where ${\boldsymbol{x}}_{0}={\boldsymbol{0}}$, ${{\boldsymbol{x}}_{\mathrm{f}}}={\boldsymbol{e}}_{5}$, $u_{\mathrm{m}}=1$, and ${\boldsymbol{x}}_{\mathrm{m}}=\left(0.8,0.5,0.5,0.5,1\right)$.
-A feasible BBS trajectory ${\boldsymbol{x}}_{1}$ is planned by [36], as shown in Fig.
-6(a).
-Denote $\mathrm{DOF}\triangleq N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|-n$; hence, Corollary 2 implies that $\mathrm{DOF}\leq 0$ holds for an optimal trajectory.
-As shown in Fig.
-6(b), $\mathrm{DOF}_{1}=6>0$ which does not satisfy the proposed state-centric necessary condition.
-Fix the first 5 arcs as well as the ASL, and vary the motion time of the 6th arc to search the minimal terminal time until reaching the boundary of the feasible set.
-If some arcs reach zero time, then delete the system behavior from the ASL.
-Repeat the above process, resulting in trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$.
-As shown in Figs.
-6(a) and 7(b), the terminal time decreases through iteration, and the DOF finally reaches 0 after 5 iterative steps.
-Finally, the optimized trajectory ${\boldsymbol{x}}_{6}$ achieves a 17.7% reduction in the terminal time compared to the original trajectory ${\boldsymbol{x}}_{1}$.
-In fact, it remains challenging to plan near-optimal high-order trajectories using arc-based methods despite advantages in computational accuracy and analytical simplicity.
-The proposed state-centric condition has the potential to optimize BBS trajectories planned by arc-based methods, which will be investigated in future works.
-Figure 6: Iteration of a 5th-order feasible BBS trajectory. The initial trajectory ${\boldsymbol{x}}_{1}$ is planned by [36]. (a) The states and controls. (b) The ASLs.
-A discretized optimization method [37] is applied to plan the trajectory ${\boldsymbol{x}}_{\text{base}}$ for comparison, where the time domain is discretized into 1000 intervals.
-An optimal solution is obtained with a 0.016% relative error of terminal time compared to ${\boldsymbol{x}}_{6}$, as shown in Fig.
-7(b).
-Therefore, [37] can plan optimal or near-optimal trajectories for the high-order problem.
-However, $u_{\text{base}}$ exhibits oscillation, as shown in Fig.
-7(a1), while $u_{i}$ is BBS.
-Defining the total variation of the control as $T_{\mathrm{v}}\triangleq\int_{0}^{{t_{\mathrm{f}}}}\left|u\left(t\right)\right|\mathrm{d}t$ to describe the trajectory stability, ${\boldsymbol{x}}_{5}$ achieves an 8.9% $T_{\mathrm{v}}$ of ${\boldsymbol{x}}_{\text{base}}$.
-For discretized optimization like [37], the open-loop errors of terminal states are difficult to constrain due to numerical error in each interval.
-Define $E_{\mathrm{s}}=\sqrt{\sum_{k=1}^{n}\left(\frac{\hat{x}_{{\mathrm{f}}k}-x_{{\mathrm{f}}k}}{x_{\mathrm{m}k}}\right)^{2}}$ where $\hat{\boldsymbol{x}}_{\mathrm{f}}=\left(\hat{x}_{{\mathrm{f}}k}\right)_{k=1}^{n}$ is directly integrated from the control input.
-As shown in Fig.
-7(d), arc-based trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$, reduce at least 5 orders of magnitude of $E_{\mathrm{s}}$ compared to ${\boldsymbol{x}}_{\text{base}}$.
-Hence, an approach combining arc-based methods and the proposed state-centric necessary condition has the potential to plan near-optimal trajectories with higher trajectory stability and computational accuracy compared to discretized optimization.
-Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [36], the optimized trajectory ${\boldsymbol{x}}_{6}$ through the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [37].
-(a) The states and controls.
-(b) The terminal time.
-(c) The total variation of the control.
-(d) The open-loop error of terminal states.
-VIII Conclusions
-This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input.
-The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form.
-Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility.
-Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one.
-The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be full row rank.
-In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.
-The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrator systems with full box constraints from both theoretical and numerical perspectives.
-An upper bound of the number of arcs is determined in a general sense, and a recursive equation for the junction time in chattering induced by 2nd-order constraints is derived.
-Note that the above two conclusions are challenging to prove by costate-based conditions.
-Numerical experiments showed that the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information, and it has the potential to optimize feasible BBS trajectories with higher trajectory stability and computational accuracy compared to discretized optimization methods.
-Acknowledgment
-This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB4302003, and in part by the National Natural Science Foundation of China under Grant 624B2077.
-Appendix A An Example of Application of the Proposed State-Centric Necessary Condition
+∎
+Appendix B An Example of Application of the Proposed State-Centric Necessary Condition
 Let’s introduce Corollary 4 of order $n$ as follows.
@@
 Specifically, we will try to derive the result based on PMP-based/state-centric necessary conditions as examples.
-A-A Problem Statement
+Note that the fundamental idea of the state-centric approach has already been applied in our previous work [31, Section VI.B].
+B-A Problem Statement
 Consider the following simplified 4th-order chain-of-integrator system:
@@
 (For the case where $n=4$, the answer is NO according to Corollary 4 in the manuscript.)
-A-B Preliminaries
+B-B Preliminaries
 To compare fairly the proposed state-centric necessary condition and PMP-based necessary conditions, we provide some preliminaries for both approaches.
@@
     $u\left(t\right)=\begin{dcases}-1,&t\in\left(t_{i},\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1}\right),\\ 1,&t\in\left(\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1},\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1}\right),\\ -1,&t\in\left(\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1},t_{i+1}\right).\end{dcases}$ (55)
-The proof of the above conclusions can be found in our related work [32] (Section IV-B).
+The proof of the above conclusions can be found in our related work [31, Section VI.B].
 The above conclusions are visualized in Figure 9.
 Figure 9: Chattering phenomena in time-optimal control for 4th-order chain-of-integrator systems with constraints on $u$ and $x_{2}$.
-A-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
+B-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
 Now, we try to derive the existence/non-existence of chattering based on PMP-based necessary conditions, which will be shown challenging.
@@
 The costate vector $\boldsymbol{\lambda}$} satisfies the Hamilton’s equations, i.e.,
-$\begin{dcases}\dot{\lambda}_{2}=-\lambda_{3}+\eta,\\ \dot{\lambda}_{n}=0,\\ \dot{\lambda}_{k}=-\lambda_{k+1},\,k\in\left\{1,3,4,5,\dots,n-1\right\}.\\ \end{dcases}$ (67)
+[…]
 Furthermore, $\forall t\in\left[0,{t_{\mathrm{f}}}\right]$, it holds that
@@
 PMP implies that it holds a.e. that
-$u\left(t\right)=\begin{dcases}-1,&\lambda_{1}\left(t\right)>0,\\ 0,&\lambda_{1}\left(t\right)\equiv 0\text{ holds for a period},\\ 1,&\lambda_{1}\left(t\right)<0.\\ \end{dcases}$ (69)
+[…]
 Specifically, in a period, $u\equiv 0$ if and only if a constrained arc occurs where $x_{2}\equiv 1$.
@@
 $\lambda_{1i}\left(t\right)=\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),$ (67)
-where $p_{i}$ is a polynomial of order no higher than $n-3$. Furthermore,
+where $p_{i}$ is a polynomial of order no higher than n3n-3. Furthermore,
 $\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0.$ (68)
 Therefore, a sufficient condition for the non-existence of chattering is $\mathcal{C}=\varnothing$, where
-$\mathcal{C}=\left\{\begin{aligned} &\left(\left(\mu_{i}\right)_{i\in{\mathbb{N}}^{*}},\left(t_{i}\right)_{i\in{\mathbb{N}}},\left(p_{i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}},\left(\lambda_{1i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}}\right):\\ &\qquad t_{i}<t_{i+1},\,\lim_{i\to\infty}t_{i}<\infty,\,\lambda_{0n}\not=0,\,\mu_{i}\geq 0,\\ &\qquad p_{i}\left(\bullet\right)\text{ is a polynomial of order no higher than }n-3,\\ &\qquad\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0,\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1i}\left(t\right)=\lambda_{0n}\frac{\left(-1\right)^{n-1}}{\left(n-1\right)!}\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1\,i+1}\left(t\right)=\lambda_{1i}\left(t\right)+\mu_{i}\left(t-t_{i}\right)\end{aligned}\right\}.$ (79)
+[…]
 It is significantly challenging to prove that $\mathcal{C}_{\pm}=\varnothing$.
@@
 In contrast, our proposed Corollary 4 can directly derive the recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$ in a compact and simple form.
-A-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
+B-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
 Denote
@@
 Once the initial value $\left(\tau_{i}\right)_{i=1}^{n-2}$ is given, the whole sequence $\left(\tau_{i}\right)_{i\in\infty}$ is determined.
-In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}=\varnothing$, where
+In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}$$\varnothing$, where
 […]
@@
 Specifically, the condition on (72) requires analyzing the convergence of a series for the recursive expressions (71) in the form of a system of polynomial equations.
-A-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
-Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}=\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (62).
+B-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
+Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}$$\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (52).
 Note that the above claim is a novel theoretical result.
@@
 $\det\left(\left(\tau_{i-k}+3\tau_{i-k-1}\right)^{j+1}-3\left(3\tau_{i-k}+\tau_{i-k-1}\right)^{j+1}+3\left(\tau_{i-k+1}+3\tau_{i-k}\right)^{j+1}-\left(3\tau_{i-k+1}+\tau_{i-k}\right)^{j+1}\right)_{j\in\left[2\right],k\in\left[2\right]}$ (73)
-$=$ $144\left(\tau_{i-1}-\tau_{i-3}\right)\left(\tau_{i}-\tau_{i-2}\right)f\left(\tau_{i-1}-\tau_{i};\tau_{i-2}-\tau_{i-1},\tau_{i-3}-\tau_{i-2}\right)=0,$
+[…]
 where
@@
 For the same reason recursively, it holds that
-$0<z_{i}<z_{i-1},\,\forall i\geq 2$ (87)
+$0<z_{i}<z_{i-1},\,\forall i\geq 2.$ (77)
 Let $r_{i}=1-\frac{z_{i+1}}{z_{i}}\in\left(0,1\right)$. Then, (75) implies that
@@
 (78) has two roots in $\left(0,r_{i}\right)$ and $\left(1,+\infty\right)$, respectively.
-Hence, $10<r_{i+1}<r_{i}<$ 1.
+Hence, $0<r_{i+1}<r_{i}<1$.
 Denote $r_{\infty}\triangleq\lim_{i\to\infty}r_{i}\in\left[0,1\right]$.
@@
 Therefore, the set $\widehat{\mathcal{C}}$ defined in (72) is empty. Therefore, chattering does not exist in the optimal control of problem (52) when $n=4$.
-A-F Summary
+B-F Summary
 In this appendix, we provide an example to compare the proposed state-centric necessary condition and PMP-based necessary conditions.